### v2

# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shic

Abstract
LASSO introduces shrinkage bias into estimated coefficients, which can adversely affect the desirable asymptotic normality and invalidate the standard inferential procedure based on the t-statistic. The desparsified LASSO has emerged as a well-known remedy for this issue. In the context of high dimensional predictive regression, the desparsified LASSO faces an additional challenge: the Stambaugh bias arising from nonstationary regressors. To restore the standard inferential procedure, we propose a novel estimator called IVX-desparsified LASSO (XDlasso). XDlasso eliminates the shrinkage bias and the Stambaugh bias simultaneously and does not require prior knowledge about the identities of nonstationary and stationary regressors. We establish the asymptotic properties of XDlasso for hypothesis testing, and our theoretical findings are supported by Monte Carlo simulations. Applying our method to real-world applications from the FRED-MD database — which includes a rich set of control variables — we investigate two important empirical questions: (i) the predictability of the U.S. stock returns based on the earnings-price ratio, and (ii) the predictability of the U.S. inflation using the unemployment rate.
Key words: Forecast, Data-rich Environment, Hypothesis testing, Shrinkage, Unit root
JEL code: C22, C53, C55
If modern techniques would have been useless in predicting the last depression, is there any reason to believe they would do better predicting the next? The authors say that the paucity of data available… made it impractical for them to construct a more complicated and more insightful model of the economy. A richer model, Mr. [Ray] Fair concedes, might have yielded better forecasts.”
“Economic Scene; Forecasters’ Art In 1929 and Now” by Peter Passell. New York Times, August 31, 1988,
1 Introduction
Prediction is one of the fundamental tasks in statistical analysis of economic and financial data. About one century ago, quantitative analysis and forecasting services were made available to businesses and the general public (Dominguez et al.,, 1988). Irving Fisher (1867-1947) spearheaded this initiative, pioneering data-driven forecasting practices grounded in theoretical foundations. His renowned theories, namely the monetary quantity theory, inflation-deflation theory, and index theory, were supported by statistical evidence. For example, Fisher, (1925) was one of the earliest attempts to understand the source of business cycles in association with the price level (inflation), and Fisher, (1926) was the first to report “a statistical relation between unemployment and price changes”, today better known as the Phillips curve. Inference is crucial for integrating economic theory and empirical data, which has become a tradition upheld by successive generations of researchers. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in inference have uncovered new empirical patterns in prediction practices using datasets with temporal features.
This paper aims at a plain quest: in a high dimensional linear predictive regression model when the number of potential regressors is larger than the sample size, how to conduct asymptotically valid inference for a regressor of particular interest. No paper has solved this question before. The challenges are twofold. Firstly, conventional predictive regressions were studied in the low-dimensional context. The notable feature of the predictive regression includes the persistence of the regressors (Stambaugh,, 1999), which upends the standard inferential theory based on the standard $t$-statistics. Secondly, to cope with the high dimensionality, one must estimate the coefficient via some regularization method, for example, when the underlying true regression model is sparse, LASSO (Tibshirani,, 1996) is the off-the-shelf method. It is well-known that the LASSO estimator is biased toward zero due to shrinkage, and follows a non-standard asymptotic distribution that is distinct from the asymptotic normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference based on the $t$-statistic, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors, and the shrinkage bias due to LASSO’s absolute-value penalty.
The above diagnosis hints at a potential solution. In low-dimensional predictive regressions, Phillips and Magdalinos, (2009)’s IVX method leverages a self-generated instrument to alleviate the persistence of the regressor and thereby overcome the Stambaugh bias. In high dimensional cross sectional regressions, Zhang and Zhang, (2014)’s desparsified LASSO (Dlasso) constructs a score vector for the parameter of interest and removes the shrinkage bias via an auxiliary regression. Moreover, both methods provide estimators with asymptotic normal distributions in their respective environments.
Can we alloy these two methods into one procedure to solve the inference of high dimensional predictive regression? We find the answer is no and yes. “No” is in the sense that a naive combination of them does not lead to the desirable results. The affirmative “Yes”, on the other hand, is established upon a deep understanding of the mechanism of both components and adapting them into the specific environment. This research culminates in a new IVX-Desparsified LASSO (XDlasso) estimator that is free from both biases and has an asymptotic normal distribution.
With a predictor of interest in mind, the construction of XDlasso, detailed in Algorithm 1 of Section 2.3, is summarized as follows. First, a workhorse estimator is needed to level the ground for high dimensional predictive regression. When both nonstationary and stationary predictors are present in the regression, recently Mei and Shi, (2024) established the consistency of the standardized LASSO (Slasso) so this becomes a natural candidate for the workhorse. Second, the common practice of generating the instrument in IVX is insufficient — the IV must be scale-standardized to have the stochastic order aligned with all other predictors in Slasso. Finally, the standardized IV (SIV) serves as the target variable for the auxiliary Slasso regression in Dlasso to estimate the shrinkage bias. The XDlasso estimator of the parameter of interest is defined by the bias correction term added to the initial Slasso estimator, and the companion $t$-statistic is employed for statistical inference by comparing it with critical values drawn from the quantiles of the standard normal distribution.
To tackle persistent regressors, the self-generated instrument of IVX in the second step is the key ingredient. This generated IV is less persistent than the unit root. This important feature enables us to decorrelate the IV and other covariates in the auxiliary Slasso regression in the third step, so that the XDlasso estimator boasts these two properties: (i) It is free from the Stambaugh bias and thus enjoys asymptotic normality; (ii) It reduces the order of shrinkage bias to make it correctable. In contrast, the ordinary Dlasso encounters a spurious auxiliary Slasso regression, failing to correct the bias arising from persistent regressors (See Section 2.4 for details). More importantly, XDlasso inference does not require a priori knowledge about the persistence of the regressor of interest and is thus free from pretesting bias. To the best of our knowledge, this is the first methodology to handle the inferential problem in high dimensional predictive regressions ($p\gg n$) with nonstationary predictors. This is also the first paper refining IVX technique for the high dimensional framework.
This paper establishes asymptotic normality for our proposed XDlasso estimator and the convergence rate of its standard error. Specifically, the asymptotically normally distributed XDlasso estimator is $\sqrt{n}$-consistent for a stationary regressor while its convergence accelerates to a faster rate for a unit root regressor. Monte Carlo simulations show that XDlasso successfully removes the bias for inference on the coefficient of a nonstationary regressor, but the ordinary Dlasso does not. Our procedure is applied to the high dimensional macroeconomic FRED-MD dataset (McCracken and Ng,, 2016) with both stationary and persistent variables, studying two important macro-finance problems: financial market return predictability and the Phillips curve in macroeconomics, where certain predictors of interest in both cases are highly persistent.
Literature review. This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (Bickel et al.,, 2009). It is well received and used for economic applications; Belloni et al., (2012), Shi, (2016), Caner and Kock, (2018), and Babii et al., (2022), to name a few. In recent years, the properties of LASSO estimators are studied in various topics in high dimensional time series, including nonstationary time series models (Koo et al.,, 2020; Lee et al.,, 2022; Mei and Shi,, 2024) and inference based on the heteroskedasticity and autocorrelation consistent (HAC) estimation (Babii et al.,, 2020, 2024).
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, Zhang and Zhang, (2014), van de Geer et al., (2014), and Javanmard and Montanari, (2014) have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. Adamek et al., (2023) generalize the Dlasso inference to high dimensional stationary time series. We follow this line of desparsified LASSO literature considering its convenience, typically only requiring a baseline regression and an auxiliary regression. An alternative method for statistical inference is Chernozhukov et al., (2018)’s double machine learning (DML), which has been widely used in cross sectional data (Chernozhukov et al., 2022a, ; Chernozhukov et al., 2022b, ; Fan et al.,, 2024). DML adopts sample splitting or cross fitting to correct the bias in the initial estimator. In our environment with nonstationary time series, however, it is not clear how to split the sample over the time dimension, due to the time varying nature of the data generating processes.
The other strand is the vast literature on predictive regressions. As highlighted by Campbell and Yogo, (2006) and Jansson and Moreira, (2006), non-standard distortion in the asymptotic distribution arises from persistent regressors. The peculiar asymptotic distributions disable the standard inferential procedures. There have been multiple proposals for valid inference, for example, the Bonferroni’s method (Campbell and Yogo,, 2006), the conditional likelihood method (Jansson and Moreira,, 2006), the linear projection method (Cai and Wang,, 2014), the weighted empirical likelihood approach (Zhu et al.,, 2014; Liu et al.,, 2019; Yang et al.,, 2021) and the implication-based inference (Xu,, 2020). Some of these methods are designed for univariate predictive regressions; it would be difficult to carry them over into the high dimensional case where the regularization method is required to cope with the issue of many parameters. On the other hand, Phillips and Magdalinos, (2009)’s IVX estimator gained its popularity by recovering the asymptotic normal distribution, enabling valid inference for mean regressions (Kostakis et al.,, 2015, 2018; Phillips and Lee,, 2013, 2016; Yang et al.,, 2020; Demetrescu et al.,, 2023), and quantile regressions (Lee,, 2016; Fan and Lee,, 2019; Cai et al.,, 2023; Liu et al.,, 2023) with multiple regressors. IVX restores asymptotic normality in low dimension by projecting the persistent regressor onto a self-generated IV, which is a more promising route to solve the problem in high dimension.
With the advent of big data, machine learning methods have spread to time series topics such as nonstationarity (Phillips and Shi,, 2021; Smeekes and Wijler,, 2021; Mei et al.,, 2024), cointegration testing (Onatski and Wang,, 2018; Zhang et al.,, 2019; Bykhovskaya and Gorin,, 2022), and structural breaks (Deshpande et al.,, 2023; Tu and Xie,, 2023).
Layout. The rest of the paper is organized as follows. Section 2 introduces the high dimensional predictive regression model with a mixture of stationary and nonstationary regressors and proposes XDlasso. Section 3 establishes the theoretical results justifying the size and power of the XDlasso inference procedure. Section 4 carries out simulation studies that corroborate the theory. Section 5 applies XDlasso inference to two macro-finance empirical examples. Section 6 concludes the paper. Technical proofs are relegated to the Online Appendices.
Notations. We set up the notations before formal discussions. We define $\boldsymbol{1}\left\{\cdot\right\}$ as the indicator function, and $\Delta$ as the difference operator so that $\Delta x_{t}=x_{t}-x_{t-1}$. The set of natural numbers, integers, and real numbers are denoted as $\mathbb{N}$, $\mathbb{Z}$, and $\mathbb{R}$, respectively. For some $n\in\mathbb{N}$, the integer set $\{1,2,\cdots,n\}$ is denoted as $[n]$, and the space of $n$-dimensional vectors is denoted as $\mathbb{R}^{n}.$ For $x=(x_{t})_{t\in[n]}\in\mathbb{R}^{n}$, the $L_{1}$-norm is $\left\|x\right\|_{1}=\sum_{t=1}^{n}\left|x_{t}\right|$, and the sup-norm is $\|x\|_{\infty}=\sup_{t\in[n]}|x_{t}|$. Let $0_{n}$ be an $n\times 1$ zero vector, $1_{n}$ be an $n\times 1$ vector of ones, and $I_{n}$ be the $n\times n$ identity matrix. For a generic matrix $B,$ let $B_{ij}$ be the $(i,j)$-th element, and $B^{\top}$ be its transpose. Let $\|B\|_{\infty}=\max_{i,j}|B_{ij}|$, and $\lambda_{\min}(B)$ and $\lambda_{\max}(B)$ be the minimum and maximum eigenvalues, respectively. Define $a\wedge b:=\min\left\{a,b\right\}$, and $a\vee b:=\max\left\{a,b\right\}$. An absolute constant is a positive, finite constant that is invariant with the sample size. The abbreviation “w.p.a.1” is short for “with probability approaching one”. We use $\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}$ and $\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ to denote convergence in probability and in distribution, respectively. For any time series $\{a_{t}\}_{t=1}^{n}$, we use $\bar{a}$ to denote its sample mean $n^{-1}\sum_{t=1}^{n}a_{t}$. For any time series $\{a_{t}\}$ and $\{b_{t}\}$, we say they are asymptotically uncorrelated if their sample correlation coefficient $\frac{\sum_{t=1}^{n}(a_{t}-\bar{a})(b_{t}-\bar{b})}{\sqrt{\sum_{t=1}^{n}(a_{t}-\bar{a})^{2}\sum_{t=1}^{n}(b_{t}-\bar{b})^{2}}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$ as $n\to\infty.$
2 Model and Procedure
2.1 High Dimensional Predictive Regression
Suppose that a time series of outcome $y_{t}$ is generated by the following linear predictive regression: [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality, as explained by Mei and Shi, (2024) that in LASSO the intercept can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model.
$y_{t}$ $=W_{t-1}^{\top}\theta^{*}+u_{t}=X_{t-1}^{\top}\beta^{*}+Z_{t-1}^{\top}\gamma^{*}+u_{t},$ (2.1)
where the error term $u_{t}$ is a stationary and martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. We consider two types of regressors with different stochastic properties: $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ is a $p_{x}\times 1$ vector of unit roots satisfying $X_{t}=X_{t-1}+e_{t}$ with $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ being a $p$-dimensional vector of stationary time series, and $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$ is a $p_{z}\times 1$ vector of stationary regressors. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($=p_{x}+p_{z}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$. Define the sparsity index $s=\sum_{j=1}^{p}\boldsymbol{1}\left\{\theta_{j}^{*}\neq 0\right\}$ as the number of nonzero components in the vector of coefficients $\theta^{*}$.
As in the default R program option glmnet::glmnet(x,y), it is a common practice in LASSO to scale-standardize each regressor $w_{j,t}$ by its sample standard deviation (s.d.)  $\widehat{\sigma}_{j}=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}},$ where $\bar{w}_{j}=n^{-1}\sum_{t=1}^{n}w_{j,t-1}$ is the sample mean. The diagonal matrix $D={\rm diag}\left(\widehat{\sigma}_{1},\widehat{\sigma}_{2},\cdots,\widehat{\sigma}_{p}\right)$ stores the sample standard deviations. The standardized LASSO (Slasso) estimator is
$\widehat{\theta}^{{\rm S}}:=\arg\min_{\theta\in\mathbb{R}^{p}}\dfrac{1}{n}\sum_{t=1}^{n}(y_{t}-W_{t-1}^{\top}\theta)^{2}+\lambda\|D\theta\|_{1}.$ (2.2)
Compared to the plain LASSO estimator that replaces $D$ by the identity matrix, the Slasso estimator is scale-invariant: if the regressor $w_{j,t-1}$ is multiplied by a nonzero constant $m$, then the $j$-th coefficient estimator changes proportionally into $\widehat{\theta}_{j}^{{\rm S}}/m.$ The standardization renders the magnitudes of unit roots into the same order as those of stationary regressors, so that the same LASSO tuning parameter $\lambda$ in (2.2) is valid for both stationary and persistent regressors.
2.2 Two Types of Biases
When data are i.i.d. or stationary, LASSO is subject to shrinkage bias as well as nonstandard asymptotic distribution which cannot be used for standard inference (Fu and Knight,, 2000). This motivates Zhang and Zhang, (2014) to bring forth the desparsified LASSO to correct the bias and recover asymptotic normality. With high dimensional unit roots, $\widehat{\theta}^{{\rm S}}$ is subject to the shrinkage bias, while the highly persistent unit root regressors incur the Stambaugh bias, further distorting the standard t-statistics inference. In this section, we examine both the shrinkage bias and the Stambaugh bias, and propose XDlasso for correcting both biases.
We are interested in inference for a null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ for some $j\in[p]$, a prevalent practice in empirical studies. [^2] A joint inference for the fixed-dimensional subvector $\theta_{\mathcal{J}}^{*}=(\theta_{j}^{*})_{j\in\mathcal{J}}$ is feasible for a subset $\mathcal{J}\subset[p]$. However, a subvector will incur additional complexity when persistent and stationary time series are mixed in the high dimensional regressors. For simplicity of exposition, we focus on inference for a scalar coefficient. In a low dimensional linear regression where $p$ is fixed, the Frisch-Waugh-Lovell theorem yields the following formulation of the ordinary least squares (OLS) estimator:
$\widehat{\theta}_{j}^{{\rm OLS}}=\frac{\sum_{t=1}^{n}w_{j,t-1}^{\perp}y_{t}}{\sum_{t=1}^{n}w_{j,t-1}^{\perp}w_{j,t-1}},$
where $w_{j}^{\perp}=(w_{j,0}^{\perp},\dots,w_{j,n-1}^{\perp})^{\top}$ is the OLS residual from regressing $w_{j,t}$ on all other regressors $W_{-j,t}=(w_{k,t})_{k\neq j}$. OLS induces a large variance as $p$ gets large, and becomes infeasible when $p>n$. Now, consider replacing the OLS residual $w_{j}^{\perp}$ by a generic score vector $r_{j}=(r_{j,0},\dots,r_{j,n-1})^{\top}$ to construct an estimator of $\theta_{j}^{*}$ that is linear in $y_{t}$ in the form of $\widehat{\theta}_{j}^{({\rm lin})}=\frac{\sum_{t=1}^{n}r_{j,t-1}y_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.$ Since $y_{t}=w_{j,t-1}\theta_{j}^{*}+W_{-j,t-1}^{\top}\theta_{-j}^{*}+u_{t}$ where $\theta_{-j}^{*}=(\theta_{k}^{*})_{k\neq j}$ is the vector of coefficients excluding the $j$-th entry, the generic estimator $\widehat{\theta}_{j}^{({\rm lin})}$ can be decomposed into
$\widehat{\theta}_{j}^{({\rm lin})}=\theta_{j}^{*}+\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\theta_{-j}^{*}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.=:\theta_{j}^{*}+{\rm N}_{j}+{\rm B}_{j},$
where ${\rm N}_{j}$ is the noise component that determines the asymptotic distribution of $\widehat{\theta}_{j}^{({\rm lin})}$, and ${\rm B}_{j}$ is the potential bias due to the choice of $r_{j}$.
For OLS, the score vector $r_{j}=w_{j}^{\perp}$ is orthogonal to the column space of $W_{-j,\cdot}:=(W_{-j,0},\dots,W_{-j,n-1})^{\top}$, under which ${\rm B}_{j}=0$ and no bias is present. The bias term ${\rm B}_{j}$ pops up whenever $r_{j}$ is not orthogonal to $W_{-j,\cdot}$, which happens if we add a penalty to the OLS objective function. With LASSO penalty we call ${\rm B}_{j}$ shrinkage bias.
Following Zhang and Zhang, (2014), we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}=(\widehat{\theta}_{k}^{{\rm S}})_{k\neq j}$:
$\widehat{\theta}_{j}=\widehat{\theta}_{j}^{({\rm lin})}-\widehat{{\rm B}}_{j}\text{ where }\widehat{{\rm B}}_{j}=\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\widehat{\theta}_{-j}^{{\rm S}}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}$ (2.3)
to compensate for ${\rm B}_{j}$. Equivalently, $\widehat{\theta}_{j}$ can be written as
$\widehat{\theta}_{j}=\widehat{\theta}_{j}^{{\rm S}}+\frac{\sum_{t=1}^{n}r_{j,t-1}\widehat{u}_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}},$ (2.4)
where $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ is the Slasso residual. Though LASSO may shrink $\widehat{\theta}_{j}^{{\rm S}}$ all the way to exactly zero, the second term in (2.4) is continuously distributed and therefore $\widehat{\theta}_{j}$ is a desparsified version of $\widehat{\theta}_{j}^{{\rm S}}$. Elementary calculation yields
$\widehat{\theta}_{j}-\theta_{j}^{*}=\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j}^{{\rm S}})}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}=:{\rm N}_{j}-(\widehat{{\rm B}}_{j}-{\rm B}_{j}),$ (2.5)
where $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ is the approximation error of the shrinkage bias.
Let $\omega_{j}$ denote the standard deviation of ${\rm N}_{j}$. To secure asymptotic normality for ${\rm N}_{j}$, we need an appropriate score vector $r_{j}$ such that as $n\to\infty$:
(R1) ${\rm N}_{j}/\omega_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1),$ and (R2) $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$.
These two results will furnish $\widehat{\theta}_{j}$ with asymptotic normality, validating inference based on the standard $t$-statistic.
The result (R1) is well established by Zhang and Zhang, (2014) under i.i.d. data when the score $r_{j,t}$ is taken as the LASSO residual of regressing $w_{j,t}$ on all other regressors. However, when $w_{j,t}$ is a unit root, this practice leads to a highly persistent $r_{j,t}$ and ruins the asymptotic normality of ${\rm N}_{j}$. The issue that highly persistent regressors produce an asymptotically non-normally distributed OLS estimator that is skewed away from zero is called the Stambaugh bias (Stambaugh,, 1999). To safeguard (R1), we must seek a score that is less persistent than a unit root.
To retain (R2), we again examine the estimation error of the shrinkage bias
${\rm\widehat{B}}_{j}-{\rm B}_{j}$ $=-g_{j}^{\top}(\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}),\ \text{where}\ g_{j}=\dfrac{\sum_{t=1}^{n}W_{-j,t-1}r_{j,t-1}}{\sum_{t=1}^{n}w_{j,t-1}r_{j,t-1}}\cdot$ (2.6)
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\cdot\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.6) suggests that a weak correlation in sup-norm between $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the unit root regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to remove the shrinkage bias.
2.3 IVX-Desparsified LASSO
For inference of $\theta_{j}^{*}$ in low dimensions, Phillips and Magdalinos, (2009)’s IVX method adopts a generated instrument by $w_{j,t}$. When $w_{j,t}$ has a unit root, this self-generated instrument is mildly integrated. The mitigation of persistence will remove the Stambaugh bias when the sample size passes to infinity, and thus recovers asymptotic normality of the IVX estimator. Due to the coexistence of Stambaugh bias and shrinkage bias, the wisdom of IVX cannot be directly transplanted into the high dimensional case. Instead, we integrate the IVX method with the idea of desparsified LASSO after comprehending the mechanism of the biases illustrated in Section 2.2.
Specifically, IVX adopts the following instrumental variable:
$\zeta{}_{j,t}=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}\Delta w_{j,s}$ (2.7)
where $\rho_{\zeta}\in(0,1)$ is a user-determined tuning parameter. Define the sample s.d. of the instrument as $\widetilde{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=0}^{n-1}\left(\zeta_{j,t}-\bar{\zeta}_{j}\right)^{2}.}$ We unify the scale for unit roots and stationary regressors by standardizing the instrument with its s.d.
$\widetilde{\zeta}_{j,t}=\zeta_{j,t}/\widetilde{\varsigma}_{j}.$ (2.8)
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.5) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see Phillips and Magdalinos, (2009). When $w_{j,t}$ is a unit root, the IV $\zeta_{j,t}$ is mildly integrated and less persistent than a unit root. Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either unit roots or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ has a unit root, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small-order $\|g_{j}\|_{\infty}$ where $g_{j}$ is defined in (2.6). If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both unit roots and stationary regressors. To this end, we construct a “residual” score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ (2.9)
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu\|D_{-j}\varphi\|_{1}$ (2.10)
where $\mu$ is the LASSO tuning parameter, and $D_{-j}={\rm diag}(\{\widehat{\sigma}_{k}\}_{k\neq j})$. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.7), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.10).
The score vector $\widehat{r}_{j}$ by (2.9) accommodates stationary and unit root regressors. When $w_{j,t}$ is stationary, the magnitude of the instrument $\widetilde{\zeta}{}_{j,t}$ shares the same order as that of $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is exactly the score in Zhang and Zhang, (2014) for cross-sectional data, which is also used in Adamek et al., (2023) for stationary time series. When $w_{j,t}$ has a unit root, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion below (2.8) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for a unit root regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for both stationary and nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.
Remark 1.
The dependent variable of the auxiliary LASSO regression (2.10) is the standardized instrumental variable (2.8). It is possible to use its original form (2.7). However, the sample s.d. of $\zeta_{j,t}$ passes to infinity if $w_{j,t}$ has a unit root. Thus, without standardizing the instrument, the theoretical order of $\mu$ for a unit root $w_{j,t}$ would be far larger than the order for a stationary regressor. This difference complicates the theoretical justifications. The standardization in (2.8) makes it possible to unify the theoretical convergence rate of $\mu$ regardless of $w_{j,t}$ being a unit root or stationary regressor.
Following (2.4), XDlasso is constructed as
$\widehat{\theta}_{j}^{{\rm XD}}=\widehat{\theta}_{j}^{{\rm S}}+\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\widehat{u}_{t}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ (2.11)
with the standard error
$\widehat{\omega}_{j}^{{\rm XD}}=\dfrac{\sqrt{\widehat{\sigma}_{u}^{2}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}$ (2.12)
where $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}.$ With this point estimator and the associated standard error, we perform the $t$-test for the null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$. Below, we summarize the testing procedure using the XDlasso estimator.
Algorithm 1 (XDlasso Inference for $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$).
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.2). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.7), and standardize it by (2.8). Step3 Run the auxiliary LASSO regression (2.10), and save the residual $\widehat{r}_{j,t}$ in (2.9). Step4 Compute the XDlasso estimator (2.11) and the standard error (2.12). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big{/}\widehat{\omega}_{j}^{{\rm XD}}.$ (2.13) Reject $\mathbb{H}_{0}$ under the significance level $\alpha$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{a}$ is the $100a$-th percentile of the standard normal distribution.
The testing procedure in Algorithm 1 refines the conventional predictive regression inference using modern high dimensional inference techniques. In the following, we elaborate on the necessity of the IVX transformation by explaining the drawback of Dlasso under high dimensional unit roots.
2.4 Necessity of IVX Transformation
Given i.i.d. data, Dlasso proceeds with the following residual as the score vector:
[…]
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu\|D_{-j}\psi\|_{1}.$ (2.15)
where the dependent variable in (2.14) is also scaled here to keep in line with (2.9) and (2.10).
When $w_{j,t}$ is a unit root regressor, this standard procedure fails to correct the bias, hindering valid inference. First, (2.15) is a spurious regression as in Granger and Newbold, (1974). In a low dimensional regression where all regressors have unit roots, the standardized least squares counterpart of the regression (2.15) follows a functional central limit theorem (FCLT)
$\widetilde{\psi}^{(j,{\rm OLS})}$ $=$ $D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{w}_{j,t}$ (2.16)
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ $\widetilde{\psi}^{*}:={\rm diag}(\{\sigma_{k}^{*}/\sigma_{j}^{*}\}_{k\neq j})\left(\int_{0}^{1}\mathcal{B}_{-j}\mathcal{B}_{-j}^{\top}\right)^{-1}\int_{0}^{1}\mathcal{B}_{-j}\mathcal{B}_{j},$
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the least square estimator converges in distribution to a nondegenerate random variable. This is the well-known spurious regression phenomenon in unit root regressions. While (2.16) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.15) in high dimensions depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, as a consequence of the spurious regression, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$, which under large $n$ is expected to be approximated by $\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{*}$ in view of (2.16), remains highly persistent. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector in low dimensions would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. The same issue is also anticipated in the score vector $r_{j,t}$ in (2.14) by LASSO in high dimensions.
Figure 1: Histograms of $t$-statistics from XDlasso and Dlasso
In contrast, with the help of the IVX transformation, the score vector in (2.9) achieves (R1) and (R2). Figure 1 provides an illustrative simulation to compare Dlasso with XDlasso. We set $\beta_{1}^{*}$ (the coefficient of the first unit root) as zero, $n=300$, $(p_{x},p_{z})=(150,300)$, and use i.i.d. innovations (See Eq. (4.1)). We generate the data following the DGP (2.1), with a mixture of unit root and stationary regressors as in Section 4.1. Figure 1 displays the histograms of the $t$-statistics over 2000 replications. The density of XDlasso $t$-statistic is well approximated by $\mathcal{N}(0,1)$, whereas the Dlasso $t$-statistic suffers from a substantial bias.
3 Asymptotic Theory
We now develop the limit theory to shed light on the asymptotic behavior of the XDlasso estimator. The bias correction of XDlasso is mounted on the workhorse estimator Slasso. Slasso’s consistency is a prerequisite for the ensuing maneuver. Unsurprisingly, this paper’s assumptions share similarities with those in Mei and Shi, (2024). We first state our theoretical assumptions and then highlight and explain the differences between the two papers.
Regarding the asymptotic framework, we define the number of regressors $p=p\left(n\right)$ and the sparsity index $s=s\left(n\right)$ as deterministic functions of the sample size $n$. In asymptotic theory, we explicitly send $n\to\infty$, and it is understood that $p(n)\to\infty$ as $n\to\infty$, while $s\left(n\right)$ is allowed to be either fixed or divergent. Recall that the stationary $e_{t}=X_{t}-X_{t-1}\in\mathbb{R}^{p_{x}}$ is the innovation of the unit root regressor. We assume that the stationary high dimensional vector $v_{t}=(e_{t}^{\top},Z_{t}^{\top})^{\top}$ is generated by the linear transformation of innovations $\varepsilon_{t}=(\varepsilon_{k,t})_{k\in[p]}$:
$v_{t}=\Phi\varepsilon_{t}$ (3.1)
where $\Phi$ is a $p\times p$ deterministic matrix. Let $\mathcal{F}_{t}$ denote the $\sigma$-field generated by $\{u_{s},\varepsilon_{s}\}_{s\leq t}$.
Assumption 1.
Suppose that $u_{t}$ and $\varepsilon_{t}$ are strictly stationary. Moreover, $u_{t}$ is a martingale difference sequence (m.d.s.) such that $\mathbb{E}\left(u_{t}|\mathcal{F}_{t-1}\right)=0$ and $\mathbb{E}(u_{t}^{2}|\mathcal{F}_{t-1})=\sigma_{u}^{2}>0$. There exist absolute constants $C_{u}$, $b_{u}$, $C_{\varepsilon}$, and $b_{\varepsilon}$ such that for all $t\in\mathbb{Z}$ and $a>0$,
$\Pr\left(|u_{t}|>a\right)$ $\leq C_{u}\exp(-a/b_{u}),$ (3.2)
$\Pr\left(|\varepsilon_{k,t}|>a\right)$ $\leq C_{\varepsilon}\exp(-a/b_{\varepsilon}),$ (3.3)
for all $k\in[p]$. Furthermore, $\varepsilon_{k,t}$ and $\varepsilon_{\ell,t}$ are independent for all $k\neq\ell.$
The following Assumption 2 imposes restrictions on the $\alpha$-mixing coefficients that characterize the time dependence of the innovations $u_{t}$ and $\varepsilon_{t}$. For any two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, define $\alpha(\mathcal{A},\mathcal{B})=\sup_{A\in\mathcal{A},B\in\mathcal{B}}|\Pr\left(A\cap B\right)-\Pr(A)\Pr(B)|$ and $\alpha(d)=\sup_{s\in\mathbb{Z}}\alpha(\sigma(\{u_{t},\varepsilon_{t}\}_{t\leq s}),\sigma(\{u_{t},\varepsilon_{t}\}_{t\geq s+d}))$.
Assumption 2.
There exist some absolute constants $C_{\alpha}$, $c_{\alpha}$, $r$, $c_{\varepsilon}$ such that
$\alpha(d)\leq C_{\alpha}\exp\left(-c_{\alpha}d^{r}\right),\ \ \forall d\in\mathbb{Z},$ (3.4)
and $\mathbb{\mathbb{E}}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]\geq c_{\varepsilon}$ for any $k\in[p].$
The following Assumption 3 depicts the contemporary correlation of $v_{t}$ defined as (3.1). Define $\Omega=\Phi\Phi^{\top}$ where $\Phi$ has appeared in (3.1).
Assumption 3.
There are absolute constants $c_{\Omega}$, $C_{\Omega}$ and $C_{L}$ such that: (a) $c_{\Omega}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq C_{\Omega}$; (b) $\max_{j\in[p]}\sum_{\ell=1}^{p}\left|\Phi_{j\ell}\right|\leq C_{L}$.
We specify the user-determined parameter in (2.7) as
$\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ (3.5)
with absolute constants $C_{\zeta}>0$ and $\tau\in(0,1)$. The choice of $\tau$ determines the persistence of the IV $\zeta_{j,t}$, which is elaborated in Remark 5. The following Assumption 4 characterizes the rate of variable dimension $p$ and sparsity index $s$ with respect to the sample size $n$.
Assumption 4.
(a) $p=O(n^{\nu})$ for an arbitrary $\nu>0$ and (b) $s=O(n^{\frac{1}{4}(\tau\wedge(1-\tau))-\xi}\wedge p^{1-\xi})$ for an arbitrary small $\xi>0$.
In the following, we expound the differences between the assumptions above and those in Mei and Shi, (2024). First, Assumption 1 imposes the m.d.s. and conditional homoskedasticity conditions for the error term $u_{t}$. Three remarks are in order to justify these two conditions.
Remark 2.
Though the m.d.s. assumption is not required in Mei and Shi, (2024), in this paper it is essential for the asymptotic normality of XDlasso. Without the m.d.s. assumption, we would call for long-run covariances that not only complicate the procedures but also rule out stationary regressors in the theory. See Phillips and Lee, (2016, Remark 2.3) for detailed discussions.
Remark 3.
In empirical finance, the m.d.s assumption is commonly imposed on the error term, especially for testing asset return predictability (Zhu et al.,, 2014; Kostakis et al.,, 2015). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the efficient market hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s. assumption plausible. [^3] Mild model misspecification with approximate sparsity can be accommodated by our framework; see Belloni et al., (2012) and Mei and Shi, (2024, Remark 1).
Remark 4.
It is possible to extend our methodology and theoretical results to conditional heteroskedastic errors. Under low dimensional predictive regression, Kostakis et al., (2015, Theorem 1) show that the homoskedastic-only standard error of the IVX estimator is robust to conditional heteroskedastic error terms if the regressor of interest is persistent. We conjecture that this result applies to XDlasso, and the expression of our standard error (2.12) is robust to conditional heteroskedasticity in our two empirical applications, where each predictor of interest is persistent. A complete theory for conditional heteroskedasticity in high dimensional predictive regression will deserve a standalone paper for future research.
Second, Mei and Shi, (2024) assumed the innovations follow linear processes, and imposed the mixing condition and the lower bounded long run variances through the coefficients in the linear process. In contrast, our Assumption 2 does not assume any specific form of the linear process for $u_{t}$ and $\varepsilon_{t}$, but directly imposes the exponentially decaying rate for the mixing coefficient and the lower bound of the long run variances.
Third, Assumption 4(a) follows Mei and Shi, (2024) that allows the variable dimension $p$ to diverge at a polynomial rate of the sample size $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Assumption 4(b) imposes a more restrictive condition for the sparsity index $s$, compared to $s=o(n^{1/4})$ in Mei and Shi, (2024), which is reasonable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.13) below.
Remark 5.
In practice the sparsity index $s$ is unknown. We therefore recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (Phillips and Lee,, 2016; Kostakis et al.,, 2015) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). The context is different in our setting of high dimensional predictive regression, under which An excessively large $\tau$ will require a very small sparsity index $s$ relative to $n$ and thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1.
By Assumptions 14, we first display a lemma for the consistency of the initial workhorse Slasso estimator. For ease of exposition, we present the convergence rate of Slasso for the coefficients standardized by the sample standard deviations.
Lemma 1.
Under Assumptions 14, there exists some absolute constant $C_{{\rm m}}$ such that when the LASSO tuning parameter (2.2) for the main regression satisfies $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
$\|D(\widehat{\theta}^{{\rm S}}-\theta^{*})\|_{1}$ $=O_{p}\left(\dfrac{s^{2}}{\sqrt{n}}(\log p)^{6+\frac{1}{2r}}\right).$ (3.6)
Lemma 1 provides the consistency of the standardized coefficients, which is necessary to establish the asymptotic normality of XDlasso. The rates of tuning parameters are technical devices for proofs. In practice, we recommend cross-validation for the tuning parameter selection.
3.1 The Auxiliary Regression
While the convergence of Slasso in the main regression is well expected, new theory must be developed for the auxiliary regression (2.10), again by the workhorse Slasso when $w_{j,t}$ is either a unit root or stationary regressor. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be simultaneously applicable to the two very different types of time series in a unified manner. Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of unit root and stationary regressors, respectively.
We first examine the case when $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in Phillips and Lee, (2016) yields the following rate of convergence in OLS:
$D_{-j}\widehat{\varphi}^{(j){\rm OLS}}=D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{\zeta}_{j,t}=O_{p}\left(1/\sqrt{n^{\tau\wedge(1-\tau)}}\right).$ (3.7)
In high dimensional models, the sample Gram matrix is rank-deficient and the FCLT no longer works. Thanks to the $L_{1}$ penalization, Slasso has a comparable local-to-zero order as (3.7), which is shown in Proposition 1 below.
We then turn to $j\in\mathcal{M}_{z}$, and slightly abuse the notation $Z_{-j,t}$ to denote the vector of stationary regressors excluding $w_{j,t}$. Define
$\varphi_{0z}^{(j)*}:=\mathbb{E}\left(Z_{-j,t}Z_{-j,t}^{\top}\right)^{-1}\mathbb{E}\left(Z_{-j,t}w_{j,t}\right)$ (3.8)
as the linear projection of $w_{j,t}$ on $Z_{-j,t}$, and normalize it by the standard deviation of the IV as
$\varphi_{z}^{(j)*}=\widehat{\varsigma}_{j}^{-1}\varphi_{0z}^{(j)*}.$ (3.9)
The “pseudo-true” model for the LASSO regression (2.10) is
$\widetilde{\zeta}_{j,t}=X_{t}^{\top}0_{p_{x}}+Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t}=Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t},$ (3.10)
[…]
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$ and thus asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficients $\varphi_{0z}^{(j)*}$ satisfy $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the LASSO estimator $\widehat{\varphi}^{(j)}$ in (2.10).
We impose the following assumption on the coefficient $\varphi_{0z}^{(j)*}$.
Assumption 5.
$\|\varphi_{0z}^{(j)*}\|_{0}\leq s$ with $\varphi_{0z}^{(j)*}$ defined as (3.8) and $s$ specified in Assumption 4. Moreover, $\|\varphi_{0z}^{(j)*}\|_{1}\leq C_{1}$ for some absolute constant $C_{1}.$
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.10). This is similar to the high dimensional sparse instrumental variable regression; see Zhu, (2018) and Gold et al., (2020). In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of Zhang and Cheng, (2017, pp. 759). Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.10). Parallel conditions naturally hold for $j\in\mathcal{M}_{x}$ as the pseudo-true coefficients are zero in view of the local-to-zero OLS estimate displayed in (3.7).
The following proposition formally lays out the convergence rate of the auxiliary LASSO estimator (2.10).
Proposition 1.
Suppose that $j\in\mathcal{M}_{z}$ and Assumptions 15 hold. Then, there exists some absolute constant $C_{{\rm a}}>0$ such that when $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right)$ (3.11)
where $\varphi^{(j)*}={\bf 1}\{j\in\mathcal{M}_{x}\}\cdot 0_{p-1}+{\bf 1}\{j\in\mathcal{M}_{z}\}\cdot(0_{p_{x}}^{\top},\varphi_{z}^{(j)*\top})^{\top}$.
Remark 6.
The rate of tuning parameter $\mu$ specified in Proposition 1 induces the following Karush–Kuhn–Tucker (KKT) condition
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu}{2}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.12)
which is sufficient to bound the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ in (2.5). Nevertheless, it is still necessary to establish the consistency of $\widehat{\varphi}^{(j)}$. This consistency result not only helps us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.13), which guarantees the asymptotic size of the test, but is also critical for the convergence rate of the standard error that governs the asymptotic power.
These key theoretical results about asymptotic size and power are displayed in the next section.
3.2 Asymptotic Normality
The consistency of the auxiliary regression built in Proposition 1 guarantees the following decomposition of the statistic:
$\frac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}={\rm sgn}_{j}\cdot\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}+O_{p}\left(\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right),$ (3.13)
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or $-1$ with probability one as $\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.5). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term of the right-hand side of (3.13) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
Theorem 1.
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{{\rm m}}$ and $C_{{\rm a}}$ such that when $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, respectively, as $n\to\infty$, we have
$(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*})\big{/}\widehat{\omega}_{j}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (3.14)
The asymptotic normality in Theorem 1 is our main theoretical result that delivers the valid asymptotic size of the hypothesis testing for $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ using the $t$-statistic $t_{j}^{{\rm XD}}$ in (2.13). The following Theorem 2 provides the convergence rate of the estimated standard error, which characterizes the asymptotic power of the test.
Theorem 2.
Conditions in Theorem 1 yield
$\widehat{\omega}_{j}^{{\rm XD}}=\boldsymbol{1}\left\{j\in\mathcal{M}_{x}\right\}\cdot O_{p}\left(1/\sqrt{n^{1+\tau}}\right)+\boldsymbol{1}\left\{j\in\mathcal{M}_{z}\right\}\cdot O_{p}\left(1/\sqrt{n}\right).$
The convergence rates displayed in Theorem 2 are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, for the null hypothesis $\mathbb{H}_{0}:\beta_{j}^{*}=0$, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for a unit root $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of unit root regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (Phillips,, 2015; Lee et al.,, 2022), and the test by XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.
Remark 7.
Unlike the convergence rates of Slasso in Lemma 1, the convergence rate of XDlasso in Theorem 2 is independent of the variable dimension $p$ and the sparsity index $s$. In view of (2.5), in XDlasso the order of the approximation error of the shrinkage bias, $\widehat{{\rm B}}_{j}-{\rm B}_{j}$, depends on $p$ and $s$, but it is dominated by the order of the noise component ${\rm N}_{j}$. The convergence rate of XDlasso therefore follows the order of ${\rm N}_{j}$, which relates only to the sample size $n$.
Remark 8.
Our theoretical results do not cover cointegrated data. Mei and Shi, (2024, Theorem 4) show that in the presence of cointegration, Slasso over-penalizes the coefficients of cointegrated regressors and shrinks them all the way to zero, regardless of their true values. We are unaware of any regularization method that achieves consistent estimation in the $p\gg n$ scheme with cointegrated variables mixed with unit roots and stationary ones. Without consistency, inference about the cointegrated variable’s true parameter is beyond reach at this moment.
Let us summarize the insights gained from the theoretical development. The fundamental principle behind the Dlasso method, as discussed in Zhang and Zhang, (2014), is based on the Frisch-Waugh-Lovell theorem. This theorem purges the influence of other control variables to overcome shrinkage bias. In the context of persistent predictors, the two-stage least squares approach, as developed in Magdalinos and Phillips, (2009) is used to eliminate Stambaugh bias. Each piece is a modern version of a classical idea. To adapt these existing procedures to high dimensional predictive regressions, we must rely on the consistency of Slasso, as proved in Mei and Shi, (2024). This consistency is crucial for both the main regression and the auxiliary regression. Moreover, as mentioned in Remark 5, the construction of the IV and the choice of $\tau$ must be tailored to balance the two types of predictors. This approach is original in that such a restriction is unique to high dimensional models and has not been studied before even in conventional low-dimensional settings.
4 Simulation
In this section, we evaluate the performance of the proposed XDlasso inference procedure by comparing its test size and power with those of Dlasso. Although Dlasso, as well as other existing LASSO-bias-correction procedures, is not designed to handle persistent regressors, we include this comparison to highlight the value added by the IVX transformation. Furthermore, to demonstrate the robustness of the XDlasso approach, we benchmark it using tests based on infeasible estimators where an oracle reveals the locations of nonzero coefficients.
4.1 Setup
We consider the linear predictive regression model in (2.1). The vector of stationary components, denoted as $v_{t}=\left(u_{t},\Delta X_{t}^{\top},Z_{t}^{\top}\right)^{\top}$, is generated by the two different processes:
Case I (IID Innovations): $v_{t}\sim\text{i.i.d. }\mathcal{N}\left(0,\Sigma\right),$ (4.1)
Case II (AR(1) Innovations): $v_{t}=R_{n}v_{t-1}+\xi_{t},\ \ \xi_{t}\sim\text{i.i.d. }\mathcal{N}\left(0,\Sigma\right),$ (4.2)
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d. satisfying the m.d.s. condition in Assumption 1, while the unit root generators $\Delta X_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
$\Sigma_{ij}=\begin{cases}0,&\text{if }(i,j)\text{ is associated with }Z_{t}\text{ and }u_{t};\\ 0.5^{|i-j|},&\text{otherwise}.\end{cases}$
Recall $0_{p}$ is a $p$-dimensional zero vector and $1_{p}$ is a $p$-dimensional vector of ones. The true coefficient vectors are:
$\beta^{*}=\left(\beta_{1}^{*},\dfrac{0.5}{\sqrt{n}}\times 1_{4}^{\top},0_{p_{x}-5}^{\top}\right)^{\top},\ \ \gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},0_{p_{z}-5}^{\top})^{\top}.$
The specification involves four active unit root regressors and four active stationary regressors. The $1/\sqrt{n}$ scaling balances the regression by normalizing the coefficients of the unit root regressors. We test the hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ with sample sizes $n\in\{200,300,400,500,600\}$ and the dimensionality pairs $(p_{x},p_{z})\in\{(50,100),(100,150),(150,300)\}.$ We conduct 2000 replications for each setting.
We compare the finite sample performance of the testing procedure XDlasso, as described in Algorithm 1, and Dlasso with the score vector in (2.14) and (2.15). In addition, we consider two infeasible testing procedures as benchmarks. Using the known active set of regressors, we conduct IVX inference (IVX oracle) and the standard $t$-test based on the OLS estimator (OLS oracle), employing only the regressor of interest and the active regressors, which form a low-dimensional predictive regression model. We set $C_{\zeta}=5$ and $\tau=0.5$ for the parameter $\rho_{\zeta}$ specified as (3.5). As discussed in Remark 5, the choice $\tau=$ 0.5 admits the weakest sparsity condition, and thus effectively improves the finite sample performance.
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rates of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following Lee et al., (2022) to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
$C_{{\rm m}}^{\left(q\right)}=\lambda^{(q)}n_{0}^{1/2}/(\log p_{0})^{\frac{3}{2}+\frac{1}{2r}},\ \ C_{{\rm a}}^{\left(q\right)}=\mu^{(q)}n_{0}^{[(1-\tau)\wedge\tau]/2}/(\log p_{0})^{2+\frac{1}{2r}},$
where $r=1$ and $\tau=0.5$ are chosen in the simulation. We then fix $\widehat{C}_{\star}=\mathrm{median}(C_{\star}^{\left(1\right)},\ldots,C_{\star}^{\left(500\right)})$ for $\star\in\left\{{\rm m},{\rm a}\right\}$ in the full-scale experiments. The tuning parameters are then set as $\widehat{\lambda}=\widehat{C}_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\widehat{\mu}=\widehat{C}_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ as in Theorem 1.
4.2 Results
We first investigate the empirical size of different testing methods at a 5% nominal significance level when the true coefficients $\beta_{1}^{*}=0$ and $\gamma_{1}^{*}=0$. Tables 1 and 2 report the empirical sizes under the IID and AR(1) innovations, respectively. Foremost among the findings is that XDlasso effectively controls the empirical size for both $\beta_{1}^{*}$, associated with a unit root regressor, and $\gamma_{1}^{*}$, associated with a stationary regressor. This performance stands in sharp contrast to that of Dlasso and OLS oracle, which exhibit severe size distortions for $\beta_{1}^{*}$. Such distortions can be attributed to the failure to account for the Stambaugh bias arising from nonstationarity. Furthermore, the results yield noteworthy insights regarding the tuning parameter selection. The empirical size of XDlasso with both cross-validated and calibrated tuning parameters is aligned with the nominal level. This result not only validates the asymptotic rates of tuning parameters specified in Theorem 1 but also supports CV as a feasible data-driven tuning parameter selection method in practice. Lastly, we investigate the efficiency and robustness of XDlasso in comparison with alternatives. When compared to the unbiased but infeasible “IVX oracle” estimator, the confidence intervals produced by XDlasso are only slightly wider. In contrast to the estimators solely for low-dimensional data, XDlasso demonstrates robustness by accommodating high dimensional covariates without compromising much in efficiency.
Table 1: Empirical size and length of confidence interval: IID innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.043 0.194 0.105 0.075 0.051 0.223 0.340 0.106 0.061 0.229 0.420 0.159
300 0.042 0.137 0.104 0.050 0.054 0.165 0.418 0.079 0.060 0.172 0.499 0.119
400 0.042 0.109 0.107 0.038 0.048 0.133 0.466 0.064 0.063 0.136 0.557 0.093
500 0.046 0.091 0.101 0.031 0.055 0.112 0.499 0.054 0.065 0.116 0.586 0.077
600 0.048 0.078 0.099 0.025 0.059 0.098 0.522 0.047 0.063 0.101 0.595 0.063
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.051 0.192 0.099 0.076 0.046 0.222 0.313 0.102 0.062 0.230 0.482 0.158
300 0.045 0.138 0.098 0.051 0.044 0.164 0.416 0.077 0.063 0.170 0.577 0.129
400 0.048 0.110 0.100 0.038 0.055 0.131 0.483 0.063 0.065 0.136 0.649 0.106
500 0.048 0.091 0.095 0.030 0.054 0.111 0.553 0.053 0.071 0.117 0.694 0.089
600 0.046 0.080 0.105 0.025 0.049 0.095 0.611 0.046 0.066 0.101 0.745 0.077
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.043 0.189 0.106 0.074 0.045 0.217 0.261 0.098 0.047 0.227 0.436 0.141
300 0.042 0.138 0.098 0.050 0.053 0.157 0.365 0.073 0.057 0.169 0.566 0.113
400 0.049 0.108 0.106 0.037 0.057 0.126 0.450 0.059 0.061 0.136 0.633 0.095
500 0.045 0.092 0.095 0.030 0.052 0.106 0.518 0.050 0.056 0.115 0.690 0.087
600 0.045 0.077 0.098 0.025 0.055 0.093 0.558 0.044 0.065 0.098 0.741 0.074
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.048 0.365 0.057 0.320 0.066 0.329 0.071 0.291 0.069 0.325 0.070 0.289
300 0.045 0.289 0.051 0.262 0.058 0.266 0.058 0.241 0.061 0.266 0.061 0.240
400 0.052 0.247 0.055 0.226 0.063 0.230 0.060 0.211 0.065 0.230 0.060 0.211
500 0.046 0.219 0.052 0.202 0.066 0.205 0.057 0.190 0.065 0.206 0.060 0.190
600 0.047 0.198 0.049 0.185 0.058 0.187 0.058 0.174 0.060 0.188 0.058 0.175
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.043 0.364 0.043 0.319 0.061 0.328 0.057 0.291 0.066 0.322 0.059 0.287
300 0.040 0.290 0.047 0.261 0.050 0.266 0.053 0.241 0.055 0.264 0.056 0.239
400 0.044 0.248 0.052 0.226 0.054 0.229 0.052 0.211 0.054 0.229 0.050 0.210
500 0.051 0.219 0.052 0.202 0.060 0.204 0.054 0.189 0.060 0.205 0.058 0.189
600 0.049 0.198 0.050 0.184 0.056 0.186 0.053 0.174 0.058 0.187 0.054 0.174
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.051 0.365 0.061 0.320 0.073 0.331 0.077 0.294 0.078 0.319 0.080 0.284
300 0.041 0.290 0.057 0.262 0.072 0.266 0.067 0.241 0.073 0.261 0.070 0.237
400 0.046 0.247 0.061 0.226 0.072 0.229 0.064 0.210 0.067 0.227 0.065 0.209
500 0.055 0.219 0.060 0.203 0.065 0.204 0.066 0.189 0.064 0.203 0.067 0.188
600 0.047 0.198 0.058 0.185 0.062 0.185 0.065 0.173 0.063 0.185 0.063 0.173
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table 2: Empirical size and length of confidence: AR(1) innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.046 0.139 0.106 0.054 0.059 0.174 0.407 0.080 0.084 0.177 0.469 0.136
300 0.044 0.097 0.115 0.035 0.048 0.123 0.478 0.060 0.076 0.127 0.545 0.096
400 0.049 0.079 0.100 0.027 0.044 0.099 0.522 0.049 0.066 0.101 0.612 0.070
500 0.049 0.065 0.097 0.021 0.059 0.083 0.556 0.041 0.074 0.084 0.610 0.057
600 0.048 0.056 0.096 0.017 0.055 0.071 0.563 0.036 0.071 0.073 0.620 0.047
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.039 0.143 0.092 0.054 0.050 0.169 0.371 0.077 0.082 0.177 0.516 0.133
300 0.055 0.101 0.103 0.036 0.046 0.123 0.480 0.058 0.081 0.129 0.624 0.108
400 0.046 0.080 0.101 0.027 0.047 0.097 0.550 0.047 0.073 0.102 0.682 0.082
500 0.046 0.067 0.099 0.022 0.050 0.081 0.614 0.040 0.071 0.085 0.739 0.068
600 0.047 0.058 0.101 0.018 0.052 0.069 0.654 0.035 0.073 0.073 0.761 0.056
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.042 0.140 0.111 0.053 0.059 0.164 0.310 0.073 0.081 0.172 0.498 0.115
300 0.040 0.101 0.098 0.036 0.054 0.118 0.436 0.055 0.068 0.126 0.589 0.089
400 0.048 0.079 0.100 0.027 0.046 0.093 0.522 0.045 0.068 0.100 0.670 0.073
500 0.048 0.066 0.106 0.022 0.055 0.079 0.586 0.038 0.067 0.084 0.728 0.067
600 0.047 0.057 0.107 0.018 0.049 0.067 0.624 0.033 0.071 0.073 0.750 0.056
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.046 0.371 0.060 0.308 0.067 0.339 0.054 0.357 0.069 0.336 0.074 0.276
300 0.049 0.293 0.056 0.251 0.072 0.272 0.050 0.286 0.074 0.270 0.070 0.229
400 0.050 0.248 0.051 0.217 0.062 0.233 0.053 0.244 0.065 0.232 0.060 0.200
500 0.050 0.219 0.056 0.194 0.064 0.207 0.050 0.216 0.064 0.206 0.062 0.181
600 0.046 0.198 0.052 0.177 0.055 0.188 0.044 0.196 0.057 0.188 0.059 0.166
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.050 0.373 0.054 0.308 0.064 0.340 0.043 0.356 0.069 0.331 0.066 0.273
300 0.045 0.293 0.052 0.250 0.056 0.271 0.042 0.284 0.057 0.269 0.059 0.228
400 0.045 0.249 0.050 0.216 0.055 0.232 0.046 0.243 0.058 0.231 0.057 0.199
500 0.052 0.219 0.054 0.194 0.055 0.206 0.043 0.215 0.052 0.206 0.062 0.180
600 0.047 0.198 0.047 0.177 0.058 0.186 0.051 0.195 0.057 0.187 0.053 0.165
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.050 0.372 0.055 0.308 0.085 0.340 0.051 0.355 0.084 0.329 0.081 0.271
300 0.051 0.293 0.062 0.251 0.077 0.271 0.054 0.285 0.075 0.266 0.073 0.226
400 0.049 0.248 0.059 0.217 0.074 0.231 0.045 0.243 0.071 0.229 0.073 0.198
500 0.046 0.219 0.056 0.194 0.064 0.205 0.051 0.216 0.061 0.204 0.072 0.179
600 0.051 0.197 0.061 0.177 0.059 0.185 0.053 0.195 0.064 0.185 0.065 0.164
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
We now turn to the empirical power of the XDlasso inference. Figure 2 plots power curves for the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ under varying true coefficient values. In our analysis, we vary the coefficient $\beta_{1}^{*}$ or $\gamma_{1}^{*}$ from 00 to $0.5$ while maintaining the other equal to 00. Across various configurations, XDlasso exhibits increasingly high power against the null hypothesis as the true coefficient moves away from 0 and the sample size $n$ increases. Furthermore, Figure 3 compares the power for testing $\beta_{1}^{*}$ and $\gamma_{1}^{*}$, which reveals that the power associated with a unit root regressor surpasses that of a stationary regressor. This observation provides finite sample evidence to support the theoretical results in Theorem 2. Specifically, it corroborates the faster convergence of standard error for unit root regressors compared with stationary regressors, thereby inducing higher power in the hypothesis testing.
Figure 2: Power curves of XDlasso inference
(a) IID Innovations
(b) AR(1) Innovations
Notes: The left and right panels correspond to DGPs (4.1) and (4.2), respectively. In each subplot, the first row depicts to the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, across various $\left(p_{x},p_{z}\right)$ configurations, while the second row pertains to $\gamma_{1}^{*}$, associated with a stationary regressor. The empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution.
Figure 3: Power of XDlasso inference for nonstationary and stationary regressors
Notes: This figure plots the power curves under $n=600.$ The first and second rows correspond to DGPs (4.2) and (4.1), respectively. In each subplot, blue lines represent the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, while red lines represent that for $\gamma_{1}^{*}$, associated with a stationary regressor. Empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.13) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ is the 97.5-th percentile of the standard normal distribution.
5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see Smeekes and Wijler, (2018), Medeiros et al., (2021), and Giannone et al., (2021), for example. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2024, sourced from the FRED-MD dataset by McCracken and Ng, (2016).
5.1 Predictability of Stock Return Using Earnings-Price Ratio
In spite of a prominent economist and statistician with a national reputation, Irving Fisher suffered devastating personal finance losses during the 1929 Great Depression (Dimand and Geanakoplos,, 2005). Were he alive today, he would have mixed feelings about the fact that return predictability remains inconclusive after a century of extensively studies. Among the various predictors examined, the predictive power of valuation ratios, particularly the earnings-price ratio, has been subject to widespread discussions. In this section, we investigate the predictability of stock returns using the log earnings-price ratio while incorporating high dimensional mixed-root control covariates.
5.1.1 Data
Our analysis focuses on predicting the monthly return of the S&P 500 index, calculated as $\text{Return}_{t}=\log(P_{t})-\log(P_{t-1}),$ where $P_{t}$ refers to S&P 500 (S&P’s Common Stock Price Index: Composite). The primary predictor of interest is the log earnings-price ratio. We obtain it from the variable S&P PE ratio (S&P’s Composite Common Stock: Price-to-Earnings Ratio), denoted as $\text{PE}_{t}$. The log earnings-price ratio is calculated by inverting the original price-earning ratio as $\text{logEP}_{t}=\log(1/\text{PE}_{t}).$
Figure B.1 displays the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2024. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table B.1. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.081$, nonstationarity is not rejected under the 5% significance level. Note that the nonstationary log earnings-price ratio is likely to have predictive power for the stationary monthly return of S&P 500, since we allow for a local-to-zero coefficient to balance the different scales between a stationary outcome and a nonstationary regressor (Phillips,, 2015). Theorem 2 and the paragraph that follows illuminate that inference by XDlasso has the power to detect a wide range of local-to-zero alternatives.
Return predictability can be time-varying and sporadic (Tu and Xie,, 2023). Campbell and Yogo, (2006) discover that the return predictability test results can vary depending on the inclusion of the samples after 1994, and the break is revisited by Zhu et al., (2014). Following this empirical finding, we divide the sample into two periods: pre-1994 (January 1960 - December 1993) and post-1994 (January 1994 - April 2024). The S&P 500 return demonstrates stationarity across both subperiods with ADF $p$-values below 1%. In contrast, the persistence of the log earnings-price ratio is ambiguous. Nonstationarity is evident in the Pre-1994 period with a large ADF test $p$-value $0.776$, while the null hypothesis of nonstationarity is rejected at the 1% significance level in the Post-1994 period. The ambiguity in stationarity of the log earnings-price ratio motivates the use of the XDlasso procedure, a unified approach for both stationary and nonstationary regressors without prior knowledge of their persistence.
Our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables are mixed with stationary and nonstationary time series. To facilitate the stationarization, FRED assigns a transformation code (TCODE) to each variable denoting the recommended transformation of the potentially nonstationary time series into a stationary one. Given the stationary pattern of the dependent variable, in our analysis, we follow the TCODE to transform the 110 time series and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in getting rid of nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at 5% significance level for both Pre-1994 and Post 1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control covariates, to demonstrate the necessity of incorporating high dimensional control covariates in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
Table 3: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ : Full sample and subperiods
Sample Periods IVX Dlasso XDlasso
Est. S.E. Est. S.E. Est. S.E.
Full Sample (Jan. 1960 - Apr. 2024) -0.023** 0.010 0.008* 0.006 0.002 0.016
Pre-1994 (Jan. 1960 - Dec. 1993) -0.058** 0.026 0.025** 0.010 0.057* 0.032
Post-1994 (Jan 1994 - Apr. 2024) -0.022 0.015 0.000 0.007 -0.001 0.019
Notes: ”Est.” and ”S.E.” represent the estimates and standard error, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. The tuning parameter for LASSO estimation is selected through 10-fold cross-validation. In XDlasso, instruments are generated based on (2.7) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table 3 presents the point estimates and testing results for $\theta_{1}^{*}$. IVX delivers negative point estimates across all sample periods and detects significant effects in the full sample period and Pre-1994 subperiod at the 5% significance level. However, the resulting negative relation between stock return and earning performance contradicts the economic mechanism between the two variables. Obviously, a large number of potential confounding variables are present in the dataset, which may be the culprit in producing the counterintuitive result from the simple regression.
Dlasso, which incorporates the high dimensional controls, reverses the signs of estimates. It returns a positively significant the coefficient for the full sample at the 10% significance level, and detects stronger evidence of predictability, at 5% significance level, in the Pre-1994 period. Further accounting for the potential nonstationarity in the predictors, XDlasso provides different estimates and standard errors from Dlasso. Notably, the standard errors of Dlasso are 60% to 68% smaller than those of XDlasso across sample periods, suggesting Dlasso may underestimate the uncertainty. Consequently, XDlasso reveals that there is no significant predictive power of log earnings-price ratio for stock return in the full sample. Moreover, the evidence of predictor power in Pre-1994 is weakened to 10% level compared with that of Dlasso at 5% level. Our empirical findings are consistent with the general recognition that there is little predictability in the financial market, thereby supporting the efficient market hypothesis.
5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As Engemann, (2020) pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded by Fisher, (1926, 1973) though, Phillips, (1958) popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation among not only modern economic studies, but also policymakers. [^4] Mary Daly, San Francisco Fed President, delivered at Daly, (2019) a negative view on the Phillips curve that “As for the Phillips curve… most arguments today center around whether it’s dead or just gravely ill. Either way, the relationship between unemployment and inflation has become very difficult to spot.” John Williams, New York Fed President, expressed a different opinion that “The Phillips curve is the connective tissue between the Federal Reserve’s dual mandate goals of maximum employment and price stability. Despite regular declarations of its demise, the Phillips curve has endured. It is useful, both as an empirical basis for forecasting and for monetary policy analysis.” See Engemann, (2020) for more details. Empirical findings find that inflation rate can be either positively or negatively correlated to unemployment, depending on the shocks to the economies, the policies, and the lag orders (Niskanen,, 2002; Gordon,, 2011, 2013). Given the ongoing debate, We revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
5.2.1 Data
The inflation rate, as the outcome variable in the predictive regression, is calculated by $\text{Inflation}_{t}=\left(\log({\rm CPI}_{t})-\log({\rm CPI}_{t-1})\right)\times 100,$ where ${\rm CPI}_{t}$ denotes the Consumer Price Index for All Urban Consumers: All Items (CPI). The unemployment rate, as the predictor of interest, denoted as $\text{Unrate}_{t}$, is retrieved in its original form under the name UNRATE.
Similar to our first empirical application in Section 5.1, we follow Benati, (2015) to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960–July 1979), Volcker and Greenspan (August 1979–January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2024). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. [^5] Pre-Volcker (before August 1979): This period was characterized by high and volatile inflation, with the Federal Reserve lacking a clear nominal anchor. Volcker and Greenspan (August 1979–January 2006): Volcker, who served as the Federal Reserve Chairman from August 1979 to August 1987, implemented aggressive anti-inflation measures, notably raising interest rates to historically high levels. Greenspan succeeded Volcker and continued to focus on maintaining price stability during his tenure, which contributed to a period of low and stable inflation known as the “Great Moderation”. Bernanke, Yellen, and Powell (February 2006–December 2019): Bernanke’s tenure as Fed Chairman was marked by the Great Recession and the implementation of unconventional monetary policies, such as quantitative easing, aimed at stimulating the economy and preventing deflation. Yellen and Powell carried on these policies. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
Figure B.2 plots the inflation and unemployment rates over our sample period. Visual inspection suggests that the unemployment rate is more persistent than the inflation rate. Table B.2 further reports the AR(1) coefficient estimates and ADF test $p$-values of the inflation and the unemployment rate under different sample periods. The inflation rate demonstrates stationary patterns in most periods, with AR(1) coefficient estimates ranging from 0.54 to 0.64. However, the inflation rate during the pre-Volcker period shows a slight upward trend, and nonstationarity is indicated by the ADF test with a $p$-value of 0.22. In contrast, the inflation rate is found to be stationary in the other two subperiods. The unemployment rate, on the other hand, appears nonstationary with AR(1) coefficient estimates close to 1. The ADF test rejects the null hypothesis of nonstationarity at a 10% significance level for both the full sample and the Volcker-Greenspan period. However, during the Pre-Volcker and Bernanke-Yellen-Powell periods, there is strong evidence of nonstationarity in the unemployment rate. The stationarity of both inflation and the unemployment rate in different periods is again ambiguous, which prompts the use of XDlasso from an agnostic perspective.
In addition to the unemployment rate, we incorporate all other 110 macroeconomic variables from the FRED-MD dataset as controls after the TCODE transformation as in Section 5.1, in view of the pattern of the dependent variable. Still, a significant proportion, from 11% to 17% across subperiods, of the transformed variables exhibit nonstationarity according to the ADF test at 5% significance level.
5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes all transformed covariates. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
Table 4: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ : Full sample and subperiods
Sample Periods IVX Dlasso XDlasso
Est. S.E. Est. S.E. Est. S.E.
Full Sample (Jan. 1960 - Apr. 2024) -0.023** 0.010 0.019*** 0.006 -0.021 0.065
Pre-Volcker (Jan. 1960 - July 1979) 0.080 0.069 0.074*** 0.017 0.013 0.224
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.035 0.063 -0.021 0.019 0.155 0.120
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2024) -0.044* 0.026 0.000 0.008 0.020 0.027
Notes: ”Est.” and ”S.E.” represent the estimates and standard error, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. The tuning parameter for LASSO estimation is selected through 10-fold cross-validation. In XDlasso, instruments are generated based on (2.7) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table 4 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso and XDlasso across sample periods. The results by XDlasso in different subperiods are unanimous that the unemployment rate does not have significant predictive power in forecasting the inflation rate across all periods. These results add new insight to the recent debates on the Phillips curve: the predictive power of unemployment for inflation appears to be diminished when we control for other high dimensional time series as covariates. In contrast, IVX and Dlasso yield contradictory results. Dlasso delivers significantly positive estimates for the full sample and the Pre-Volcker period at the 1% significance level. Conversely, IVX detects significant negative effects for the full sample and Bernanke-Yellen-Powell period at the 5% and 10% significance levels, respectively. These discrepancies may be attributed to the failure of IVX to control for confounders and the neglect of the Stambaugh bias in Dlasso arising from nonstationarity. The empirical findings echo Mankiw, (2024)’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.”
6 Conclusion
This paper proposes XDlasso to overcome the difficulties in hypothesis testing for high dimensional predictive regressions with stationary and nonstationary regressors. XDlasso fuses the IVX technique from time series econometrics and the debiasing technique from the high dimensional statistics, thereby reducing the order of biases to make them readily correctable. We establish the asymptotic normality and convergence rate of XDlasso. The validity of our methods is further evidenced by simulation studies and empirical applications.
One limitation of this paper is that it considers only the unit roots for persistent regressors. In addition to unit roots, there are many other models that portray the complex nonstationarity in real-world data, including local unit roots, fractional integration, cointegration, structural breaks, to name a few. LASSO’s consistency and inference under these time series models are yet to be studied, and one may attempt to generalize inference of the null to high dimensional parameters. These extension are currently being investigated.
Appendix A Technical Proofs
[…]
Section A.1 proves Lemma 1 that establishes consistency of Slasso under general mixing conditions without an explicit linear process. Section 3.1 includes the proofs for the results in Section 3.1 that constructs consistency of the auxiliary LASSO regression adapted for bias correction. Section A.3 includes the proofs for the theorems in Section 3.2 about the asymptotic distribution of the XDlasso estimator and the order of its standard error, which determine the size and power of the XDlasso inference.
A.1 Proof of Lemma 1
The assumptions imposed in Section 3 slightly differ from those in Mei and Shi, (2024) (MS24, hereafter). Specifically, the linear process assumption in MS24 is extended to more general $\alpha$-mixing and sub-exponential conditions. The main modification of the proof is the Gaussian approximation error deduced in MS24’s Lemma B.4. The following lemma re-establishes Gaussian approximation errors for the stationary components $v_{t}$ defined as (3.1).
Lemma A.1.
Under Assumptions 14, there exist standard Brownian motions $\{\mathcal{B}_{k}(t)\}$ with independent increment $\mathcal{B}_{k}(t)-\mathcal{B}_{k}(s)\sim\mathcal{N}(t-s)$ for $t\geq s\geq 0$ such that
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
where $V_{k}^{*}=\mathbb{E}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]$ is the long-run variance of $\{\varepsilon_{k,t}\}.$
Remark A.1.
The convergence rate in Lemma A.1 is less sharp than that in MS24’s Lemma B.4, since we work with $\alpha$-mixing sequences.
The proof of Lemma 1 follows exactly the same roadmap of the proof of MS24’s Theorem 2. Main modifications lie in the convergence rates of the Gaussian approximation in the proof of MS24’s Proposition B.4. Specifically, we follow the notations in MS24 for this proof only. The convergence rate below MS24’s (B.48) becomes
$\sup_{j\in[p+1]}\sup_{0\leq t\leq n-1}n^{-1/2}|\xi_{jt}-\zeta_{jt}|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right).$
The last equality at Page 20 in MS24’s supplement becomes
$\|\widehat{\Sigma}-\widehat{\Upsilon}\|_{\infty}=O_{p}\left(n^{3/4}(\log p)^{3/2}\right).$
Then Equation (B.49) becomes
$\dfrac{\left|\delta^{\top}(\widehat{\Sigma}-\widehat{\Upsilon})\delta\right|}{n\|\delta\|_{2}^{2}}\leq 4L^{2}sO_{p}\left(n^{-1/4}(\log p)^{3/2}\right)=o_{p}\left(\dfrac{L^{-2}}{s\log p}\right)$
under Assumption 4 in the current paper, which implies $s^{2}=o(n^{1/4-\xi})$ for an arbitrarily small $\xi$.
We prove Lemma A.1 to complete assist the proof of Lemma 1.
Proof of Lemma A.1.
Define $V_{k,t}=\mathbb{E}\left[\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}\right)^{2}\right].$ The proof includes the following two steps:
  1. 1.
    The variance of the partial sum $V_{k,t}$ is well approximates by the long run variance $V_{k}^{*}$ scale by $t$, in the sense that
    $\sup_{k\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|$ $=O\left(\dfrac{1}{n}\right).$ (A.1)
  2. 2.
    Strong Gaussian approximation: There exist standard Brownian motions $\{\mathcal{B}_{k}(t)\}$ with independent increment $\mathcal{B}_{k}(t)-\mathcal{B}_{k}(s)\sim\mathcal{N}(t-s)$ for $t\geq s\geq 0$ such that
    $\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|=O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right).$ (A.2)
Given the two steps above, Lemma A.1 follows by the triangular inequality
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|$
$\leq$ $\sup_{j\in[p],t\in[n]}\left|\dfrac{1}{\sqrt{n}}\left(\mathcal{B}_{k}(V_{k,t})-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|+\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|$
$=$ $O_{p}\left(\sqrt{\log(np)\cdot\sup_{k\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|}\right)+\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|$
$=$ $O_{p}\left(\sqrt{\dfrac{\log(np)}{n}}\right)+O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right)$
$=$ $O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
where the second row applies the fact that $n^{-1/2}\left(\mathcal{B}_{k}(V_{k,t})-\mathcal{B}_{k}(t\cdot V_{k}^{*})\right)$ for each $(j,t)$ follows a normal distribution of mean zero and variance $\left|\left(t\cdot V_{k}^{*}-V_{k,t}\right)/n\right|$, the third row applies (A.1), (A.2), and the last row applies the assumption $p\geq n^{\nu_{1}}$.
Step 1. Verifying (A.1). Define ${\rm cov}_{k}(d):={\rm cov}(\varepsilon_{k,t},\varepsilon_{k,t-d})$ as the autocovariance function of $\{\varepsilon_{k,t}\}$. Then by some fundamental calculations,
$V_{k,t}=t\cdot{\rm cov}_{k}(0)+2\sum_{d=1}^{t}(t-d)\cdot{\rm cov}_{k}(d).$ (A.3)
In addition,
$t\cdot V_{k}^{*}=t\cdot{\rm cov}_{k}(0)+2t\cdot\sum_{d=1}^{\infty}{\rm cov}_{k}(d).$
Then
$t\cdot V_{k}^{*}-V_{k,t}$ $=2t\sum_{d=1}^{\infty}{\rm cov}_{k}(d)-2\sum_{d=1}^{t}(t-d)\cdot{\rm cov}_{k}(d)$
$=2t\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)+2\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d).$ (A.4)
Recall that Assumption 2 imposes an upper bound for the $\alpha$-mixing coefficient. With $p=q=3$ in Equation (2.2) in the Corollary of Davydov, (1968, pp. 692), we have
$\sup_{j\in[p]}|{\rm cov}_{k}(d)|\leq 12\left(\mathbb{E}|\varepsilon_{k,t}|^{3}\right)^{2/3}\sqrt{\alpha(d)}\leq C\cdot\exp\left(-c_{\alpha}d^{r}/2\right).$ (A.5)
where $C=12\left(\mathbb{E}|\varepsilon_{k,t}|^{3}\right)^{2/3}\sqrt{C_{\alpha}}$ and $C_{\alpha}$ is in Assumption 2. By Equation (B.78) in MS24’s supplement,
$\sup_{t\in[n]}\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\leq\dfrac{2}{c_{\alpha}}\exp\left(-c_{\alpha}d^{r}/4\right),$
and thus uniformly for all $t,$ there exists some absolute constant $C_{1}$ such that
$\sup_{j\in[p],t\in[n]}t\cdot\sum_{d=t+1}^{\infty}|{\rm cov}_{k}(d)|\leq C\cdot\sup_{t\in[n]}t\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\leq\sup_{t\in[n]}\dfrac{4Ct}{c_{\alpha}}\exp\left(-c_{\alpha}t^{r}/4\right)<C_{1}.$ (A.6)
In addition,
$\sup_{j\in[p],t\in[n]}\sum_{d=1}^{t}d\cdot|{\rm cov}_{k}(d)|<C\sum_{d=1}^{\infty}d\cdot\exp\left(-c_{\alpha}d^{r}/2\right)<C_{2}$ (A.7)
for some absolute constant $C_{2}$. Then
$\sup_{j\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|$ $\leq\dfrac{2}{n}\cdot\left(\sup_{j\in[p],t\in[n]}t\cdot\sum_{d=t+1}^{\infty}|{\rm cov}_{k}(d)|+\sup_{j\in[p],t\in[n]}\sum_{d=1}^{t}d\cdot|{\rm cov}_{k}(d)|\right)$
$\leq\dfrac{2(C_{1}+C_{2})}{n},$
which implies (A.1).
Step 2. Verifying (A.2). We use the strong Gaussian approximation from Lin and Lu, (1997)’s Theorem 9.3.1. Specifically, define $g(x)=\exp(x)$. By the sub-exponential tail imposed by Assumption 1, the sub-exponential norm of $\varepsilon_{k,t}$, denoted as $\|\varepsilon_{k,t}\|_{g}$ in Lin and Lu, (1997), is uniformly bounded by an absolute constant. It then suffices to verify the following two conditions required in the aforementioned theorem:
  • (i)
    $V_{k,t}\geq ct$ for some absolute constant $c$.
  • (ii)
    $\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))<\infty$, where the parameter $\delta$ in Lin and Lu, (1997, Theorem 9.3.1) is taken as 2.
Then by Theorem 9.3.1 of Lin and Lu, (1997), for any $j\in[p]$
$\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})=O_{a.s.}\left(V_{k,t}^{1/4}(\log V_{k,t})^{3/2}\right).$ (A.8)
By (A.3) and (A.5),
$V_{k,t}$ $=t\left[{\rm cov}_{k}(0)+2\sum_{d=1}^{t}(1-d/t)\cdot{\rm cov}_{k}(d)\right]$
$\leq t\left[{\rm cov}_{k}(0)+2C\cdot\sum_{d=1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\right]=O(t)$
uniformly for all $(k,t)$. Then by (A.8),
$\sup_{j\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|=O_{a.s.}\left(\sup_{t\in[n]}\dfrac{t^{1/4}(\log t)^{3/2}}{\sqrt{n}}\right)=O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right),$
which leads to (A.2). It then suffices to show the Conditions (i) and (ii) above.
Proof of (i). By (A.4),
$\dfrac{V_{k,t}}{t}=V_{k}^{*}-2\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)-\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\to V_{k}^{*}$
as $t\to\infty$, where the limit applies (A.6) and (A.7). By Assumption 2, the long-run variance $V_{k}^{*}$ is bounded from below by some absolute constant. This result implies that $V_{k,t}/t$ is lower bounded by some absolute constant uniformly for all $(k,t)$.
Proof of (ii). This is a direct corollary of the exponential decaying mixing coefficient imposed by Assumption 2, in the sense that
$\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))\leq\sum_{d=1}^{\infty}C_{\alpha}^{1/4}\exp\left(-c_{\alpha}d^{r}/4\right)\cdot c_{\alpha}d^{r}<\infty.$
This completes to proof of Lemma A.1. ∎
A.2 Proofs for Section 3.1
A.2.1 Unit Root Regressors
We first introduce and prove several technical lemmas.
Lemma A.2.
Under Assumptions 14,
$\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t-1}|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p},$ (A.9)
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=1}^{n}x_{j,t-1}u_{t}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.10)
Furthermore, recall that $\widehat{\sigma}_{j}=\sqrt{n^{-1}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}}.$ Then
$n(\log p)^{-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\leq\sup_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p),$ (A.11)
$\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}\sup_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1,$ (A.12)
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=1}^{n}\widehat{\sigma}_{j}^{-1}w_{j,t-1}u_{t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}.$ (A.13)
Proof of Lemma A.2.
Eq. (A.9) comes from MS24’s (B.12). (A.10) can be justified by MS24’s Propositions 1 and 3(b), and the latter also implies (A.13). (A.11) and (A.12) are deduced in the proof of MS24’s Theorem 3. ∎
Lemma A.3.
Under Assumptions 14, for any fixed $j\in\mathcal{M}_{x}$
$\sup_{t\in[n]}|\zeta_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.14)
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=0}^{n-1}e_{k,t}\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}},$ (A.15)
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}X_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$ (A.16)
Furthermore, recall that $\widehat{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=1}^{n}(\zeta_{j,t-1}-\bar{\zeta}_{j})^{2}}.$ Then for any $j\in\mathcal{M}_{x}$,
$\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}.$ (A.17)
Proof of Lemma A.3.
We work on these inequalities one by one.
Proof of (A.14). Recall from (3.5) that $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$. When $j\in\mathcal{M}_{x}$, we have $\zeta{}_{j,t}=\sum_{s=1}^{t}\rho_{\zeta}^{s}e_{j,t-s}$ for $t\geq 1$, which is a partial sum of $\alpha$-mixing sup-exponential components $e_{j,t-s}$ weighted by $\rho_{\zeta}^{s}.$ Define
$a_{n}:=\lfloor n^{\tau}(\log p)^{2}\rfloor.$
Note that $\rho_{\zeta}^{s}e_{j,t-s}$ is sub-exponential with an exponentially decaying $\alpha$-mixing coefficient, and thus $\zeta{}_{j,t}$ is the partial sum of $t$ observations from a sub-exponential and $\alpha$-mixing time series. MS24’s Lemma B.2 yields
$\sup_{t\leq a_{n}}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{a_{n}\cdot\log p}=O\left[n^{\tau/2}(\log p)^{3/2}\right].$ (A.18)
In addition, when $t>a_{n}$,
$|\zeta{}_{j,t}|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\left|\sum_{a_{n}<s\leq t}\rho_{\zeta}^{s}e_{j,t-s}\right|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|.$ (A.19)
By the same arguments for (A.18), we bound the two sums on the right-hand side of (A.19) by
$\sup_{a_{n}<t\leq n}\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.20)
and
$\sup_{a_{n}<t\leq n}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{(n-a_{n})\cdot\log p}\text{.}$
Besides, under the assumption $p\geq n^{\nu_{1}},$ the sequence
$\rho_{\zeta}^{a_{n}}=\left(1-C_{\zeta}/n^{\tau}\right)^{\lfloor n^{\tau}(\log p)^{2}\rfloor}=O\left(\exp\left(-C_{\zeta}(\log p)^{2}\right)\right)=O\left(p^{-C_{\zeta}\log p}\right)$
converges to zero faster than the reciprocal of any polynomial function of $n.$ Thus,
$\sup_{a_{n}<t\leq n}\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\rho_{\zeta}^{a_{n}}\sqrt{(n-a_{n})\cdot\log p}=o\left(n^{\tau/2}(\log p)^{3/2}\right).$ (A.21)
By (A.19), (A.20), and (A.21), it follows
$\sup_{a_{n}<t\leq n}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}$ (A.22)
and then (A.14) follows by (A.18) and (A.22).
Proof of (A.15). Using the result (A.14), the proof of (A.15) follows exactly the same procedures in the proof of MS24’s Proposition B.2 about deviation bound (DB) for unit root. Essential modifications include:
  1. 1.
    Change the bound $\sup_{t}|x_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}$ in MS24’s Equation (B.12) into
    $\sup_{t}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$
    for a fixed $j$, which has been established (A.14). With this result, we can deduce the same upper bound of $T_{1}$ in MS24’s Equation (B.13).
  2. 2.
    Change the expression of $T_{2}$ above into MS24’s Equation (B.12)
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|.$
    The upper bound of the new expression of $T_{2}$ can be analyzed in the same way as MS24.
  3. 3.
    Change the definition of the event $\mathcal{X}_{t}$ below MS24’s Equation (B.16) into
    $\mathcal{X}_{t}=\{|\zeta{}_{j,t}|\leq C_{X}\sqrt{n\log p}\}$
    for some absolute constant $C_{X}$. Equation (A.14) in the current paper implies that $\Pr\{\bigcup_{t=1}^{n}\mathcal{X}_{t}^{c}\}\to 0$ as $n\to\infty$. Therefore, the arguments below MS24’s Equation (B.24) are still valid. We can thus show that the upper bound of $T_{3}$ defined above into MS24’s Equation (B.12) still holds.
Proof of (A.16). Note that when $j\in\mathcal{M}_{x}$, we have the following AR(1) representation
$\zeta_{j,t}=\rho_{\zeta}\zeta_{j,t-1}+e_{j,t}.$ (A.23)
Thus, for any $k\in\mathcal{M}_{x},$
$x_{k,t-1}\zeta_{j,t}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}+x_{k,t-1}e_{j,t}$
where $\rho_{\zeta}$ is defined as (3.5). By $x_{k,t-1}\zeta_{j,t}=(x_{k,t}-e_{k,t})\zeta_{j,t},$ we obtain
$x_{k,t}\zeta_{j,t}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}+x_{k,t-1}e_{j,t}+e_{k,t}\zeta_{j,t}.$ (A.24)
Summing up both sides of (A.24) and using the fact that $\sum_{t=1}^{n}x_{k,t}\zeta_{j,t}=\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{j,n}$, we deduce
$\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{k,n}=\rho_{\zeta}\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}+\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}.$
It can be further arranged into
$(1-\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}=\left(x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right)+\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}.$ (A.25)
By (A.9) and (A.14),
$\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}.$ (A.26)
By (A.10) and (A.15),
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.27)
By (A.25), (A.26), and (A.27),
$(1-\rho_{\zeta})\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}-x_{k,n}\zeta_{j,n}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}+n(\log p)^{1+\frac{1}{2r}}$
$\leq 2n(\log p)^{1+\frac{1}{2r}}.$
Thus,
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{n(\log p)^{1+\frac{1}{2r}}}{1-\rho_{\zeta}}\preccurlyeq n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$
given $1-\rho_{\zeta}=C_{\zeta}n^{-\tau}.$
Proof of (A.17). By the law of large number (LLN) in Phillips and Magdalinos, (2007, Eq. (4)), we have
$\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm var}(e_{j,t})}{2C_{\zeta}}.$ (A.28)
Without loss of generality, assume $\zeta_{j,0}=0$. Summing up both sides of (A.23), we have
$x_{j,n}-x_{j,0}=\sum_{t=1}^{n}e_{j,t}=\sum_{t=1}^{n}\zeta_{j,t}-\rho_{\zeta}\sum_{t=1}^{n}\zeta_{j,t-1}=\zeta_{j,n}+(1-\rho_{\zeta})\sum_{t=1}^{n}\zeta_{j,t-1}.$
Since $1-\rho_{\zeta}=C_{\zeta}n^{-\tau},$we have $\sum_{t=1}^{n}\zeta_{j,t-1}=n^{\tau}C_{\zeta}^{-1}(x_{j,n}-x_{j,0}-\zeta_{j,n}).$ By (A.9) and (A.14),
$\sum_{t=1}^{n}\zeta_{j,t-1}=O_{p}\left[n^{\tau}\cdot\left(\sqrt{n\log p}+n^{\tau/2}(\log p)^{3/2}\right)\right]=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right),$ (A.29)
and thus
$\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}=O_{p}\left(\dfrac{\log p}{n^{1-\tau}}\right)\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.30)
By (A.28) and (A.30),
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm var}(e_{j,t})}{2C_{\zeta}}.$ (A.31)
This completes the proof of Lemma A.3. ∎
With these preparatory lemmas, we will prove Proposition 1 for $j\in\mathcal{M}_{x}$.
Proof of Proposition 1 for $j\in\mathcal{M}_{x}$..
For simplicity of exposition, define $\widetilde{\zeta}_{j}=(\widetilde{\zeta}_{j,1},\dots,\widetilde{\zeta}_{j,n-1})^{\top}$ and recall that $W_{-j,\cdot}=(W_{-j,1},\dots,W_{-j,n-1})^{\top}.$ By the definition of Slasso, we have
$\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j,\cdot}\widehat{\varphi}^{(j)}\|_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j}\varphi\|_{2}^{2}+\mu\|D_{-j}\varphi\|_{1}.$
for an arbitrary $(p-1)$-dimensional vector $\varphi$. We can write the above inequality into
$n^{-1}\lVert W_{-j,\cdot}\left(\widehat{\varphi}^{(j)}-\varphi\right)\rVert_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq$ $\dfrac{2}{n}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\mu\|D_{-j}\varphi\|_{1}.$ (A.32)
Write the coefficient vector as $\varphi=(\varphi_{k})_{k\in[p],k\neq j}$, where $\varphi_{k}$ is the coefficient of $w_{k,t}$. Define $\varphi_{\mathcal{M}_{x}}=(\varphi_{k})_{k\in\mathcal{M}_{x},k\neq j}$ and $\varphi_{\mathcal{M}_{z}}=(\varphi_{k})_{k\in\mathcal{M}_{z}}.$ Take a vector $\varphi$ such that
$\|\varphi_{\mathcal{M}_{x}}\|_{1}$ $=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+\tau\wedge(1-\tau)}}}\text{ and }\|\varphi_{\mathcal{M}_{z}}\|_{1}=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.33)
for a positive number $C_{\varphi}=O(1)$. Without loss of generality, suppose that $s\geq 1$. We will show that
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\leq\frac{\mu}{2}\left(1-\frac{1}{2s^{2}}\right),$ (A.34)
Under (A.34), we can deduce by (A.32) that
$\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\|D_{-j}\varphi\|_{1}\right]$
$\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}+2\|D_{-j}\varphi\|_{1}\right],$
which implies $\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq 4s^{2}\|D_{-j}\varphi\|_{1}.$ In addition, by (A.11) and (A.12) in Lemma A.2, there exists a absolute constant $C_{\sigma}$ such that
$\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\leq C_{\sigma}\sqrt{n\log p}\text{ and }\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\leq C_{\sigma}$ (A.35)
w.p.a.1. Therefore, by (A.33) and (A.35),
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq 4s^{2}\|D_{-j}\varphi\|_{1}$
$\leq 4s^{2}\left(\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}\right)$
$\leq\dfrac{4C_{\varphi}\cdot C_{\sigma}s^{2}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.36)
w.p.a.1, which implies (3.11).
It is thus sufficient to prove (A.34). Note that
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}\leq(n\widehat{\varsigma}_{j})^{-1}\left(\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|+\sup_{k\in[p_{z}]}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|\right).$
Repeatedly using the bounds in (A.11), (A.12) in Lemma A.2, and (A.15), (A.16), and (A.17) in Lemma A.3,
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(n^{1+\tau/2})^{-1}\left(\dfrac{n^{1+\tau}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n/\log p}}+n(\log p)^{1+\frac{1}{2r}}\right)$
$\leq\ 2(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}.$ (A.37)
In addition,
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$
$\leq$ $\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}$
$\leq$ $\ \max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (A.38)
where the second inequality applies (A.33). By the sup-exponential distributions of the stationary components, we can deduce that
$\sup_{k\in\mathcal{M}_{z},t\in[n]}|w_{k,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p.$ (A.39)
Therefore,
$\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$ $\leq\max_{j\in[p],k\in\mathcal{M}_{x}}\max_{t\in[n]}\left|\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{x},t\in[n]}\left|w_{k,t}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p\cdot\sqrt{n\log p}=\sqrt{n(\log p)^{3}},$ (A.40)
where the fourth row applies (A.9), (A.11), (A.12), and (A.39). Following similar arguments we can deduce
$\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$ $\leq\max_{j\in[p],k\in\mathcal{M}_{z}}\max_{t\in[n]}\left|\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{z},t\in[n]}\left|w_{k,t}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p\cdot\log p=(\log p)^{2}.$ (A.41)
Thus, by (A.38), (A.40), and (A.41),
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n(\log p)^{3}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+(\log p)^{2}\cdot\dfrac{(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{2(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$ (A.42)
(A.37) and (A.42) yield
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$
Therefore, (A.34) holds as $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ with a sufficiently large $C_{{\rm a}}$. This completes the proof of Proposition 1 for $j\in\mathcal{M}_{x}$. ∎
A.2.2 Stationary Regressor
Recall $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ as defined in (3.5), and observe the following decomposition
$\zeta_{j,t}=w_{j,t}+\dfrac{C_{\zeta}}{n^{\tau}}\phi_{j,t},\ \phi_{j,t}:=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}w_{j,s-1}.$ (A.43)
In addition, we define
$\eta_{j,t}=\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$ (A.44)
where $\varphi_{0z}^{(j)*}$ is defined in (3.8). Compare $\eta_{j,t}$ to its standardized version $\widetilde{\eta}_{j,t}$ defined in (3.10), we have
$\eta_{j,t}=\widehat{\varsigma}_{j}\widetilde{\eta}_{j,t}.$ (A.45)
Finally, define the standardized regressors
$\widetilde{W}_{-j,t}=D_{-j}^{-1}W_{-j,t},$ (A.46)
and $\widetilde{W}_{-j,\cdot}=(W_{-j,0},W_{-j,1},\dots,W_{-j,n-1})^{\top}.$
Lemma A.4.
Under Assumptions 15, for $j\in\mathcal{M}_{z}$
$\sup_{t}|\phi_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.47)
$\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1,$ (A.48)
and
$\lVert n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\widetilde{\eta}_{j,t}\rVert_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{-\tau/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}},$ (A.49)
with $\widetilde{\eta}_{j,t}$ defined in (3.10), and $\widetilde{W}_{-j,t}$ defined in (A.46).
Proof of Lemma A.4.
By definition of $\phi_{j,t}$ in (A.43), we can easily deduce the following recursive formula for $\phi_{j,t}$ that
$\phi_{j,t}=\rho_{\zeta}\phi_{j,t-1}+w_{j,t-1}$
Note that $\phi_{j,t-1}\in\sigma(w_{j,0},\dots,w_{j,t-2}).$ Then $\phi_{j,t}$ is an AR(1) process with coefficient $1-C_{\zeta}/n^{\tau}$ and innovation $w_{j,t-1}$. Recall that $w_{j,t}$ is $\alpha$-mixing and sub-exponential; we obtain (A.47) following the same arguments for (A.14).
In addition, we know $n^{-1}\sum_{t=1}^{n}\phi_{j,t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}$ from (A.17). By (A.15), the cross-product between a mildly integrated $\phi_{j,t}$ and a stationary $w_{j,t}$ is bounded by
$\sum_{t=1}^{n}w_{j,t}\phi_{j,t}=o_{p}(n^{1+\tau}).$
Thus by (A.43),
$n^{-1}\sum_{t=1}^{n}\zeta_{j,t}^{2}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}+n^{-(1+2\tau)}\sum_{t=1}^{n}\phi_{j,t}^{2}+n^{-(1+\tau)}\sum_{t=1}^{n}w_{j,t}\phi_{j,t}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}+o_{p}(1).$ (A.50)
Furthermore, (A.29) implies that the mildly integrated time series $\phi_{j,t}$ satisfies
$\sum_{t=1}^{n}\phi_{j,t}=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right)=o_{p}(n^{1+\tau}).$
We thus have by (A.43) that
$n^{-1}\sum_{t=1}^{n}\zeta_{j,t}=n^{-1}\sum_{t=1}^{n}w_{j,t}+o_{p}(1).$ (A.51)
(A.50), (A.51), and a standard law of large number imply
$\widehat{\varsigma}_{j}^{2}=n^{-1}\sum_{t=1}^{n}\zeta_{j,t}^{2}-\left(n^{-1}\sum_{t=1}^{n}\zeta_{j,t}\right)^{2}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}-\left(n^{-1}\sum_{t=1}^{n}w_{j,t}\right)^{2}+o_{p}(1)\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}\left(w_{j,t}\right),$ (A.52)
which verifies (A.48).
We then show (A.49). Note that
$\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-(1+\tau)}\sum_{t=1}^{n}Z_{-j,t-1}\phi_{j,t}\|_{\infty}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{-(1+\tau)}\cdot n\sqrt{\log p}\cdot n^{\tau/2}(\log p)^{3/2}$
$=n^{-\tau/2}(\log p)^{2},$
where the first row applies the decomposition (A.43), and the second row applies the bounds (A.39) and (A.47). Also,
$\|n^{-1}\sum_{t=1}^{n}X_{t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-(1+\tau)}\sum_{t=1}^{n}X_{t-1}\phi_{j,t}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}},$
where the second inequality applies the rate in (A.16) of the cross product between a unit root and a mildly integrated series. Thus,
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\|n^{-1}\sum_{t=1}^{n}X_{t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}+\dfrac{(\log p)^{2}}{n^{\tau/2}}\leq\dfrac{2(\log p)^{2}}{n^{\tau/2}}$
with $\tau\in(0,1)$ and $n$ large enough, where the second inequality applies by (A.11) and (A.12). Further by (A.17),
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2}}{n^{\tau/2}}.$ (A.53)
In addition, define
$\eta_{j,t}^{(1)}=w_{j,t}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*}.$ (A.54)
By the definition of $\eta_{j,t}$, we have
$\eta_{j,t}^{(1)}=(\zeta_{j,t}-w_{j,t})+\eta_{j,t}.$ (A.55)
note that the time series $\eta_{j,t}^{(1)}$ is stationary and $\mathbb{E}\left[Z_{-j,t-1}\eta_{j,t}^{(1)}\right]=0$. We then have
$\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{\dfrac{\log p}{n}}$
by standard concentration inequalities; e.g. MS24’s Eq. (B.31). Also, following the same way to prove (A.10), we have
$\|n^{-1}\sum_{t=1}^{n}X_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}}.$
Thus,
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\|n^{-1}\sum_{t=1}^{n}X_{t-1}\eta_{j,t}^{(1)}\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}+\sqrt{\dfrac{\log p}{n}}\leq\dfrac{2(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}.$ (A.56)
Furthermore, by (A.48) and (A.56) we deduce that
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\leq\dfrac{(\log p)^{2}}{n^{\tau/2}}.$ (A.57)
By (A.55) and (A.45), we have the following decomposition
$\widetilde{\eta}_{j,t}=\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}+\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}.$
Then by the triangular inequality, we have
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\widetilde{\eta}_{j,t}\|_{\infty}$ $\leq\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}+\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2}}{n^{\tau/2}}.$
where the second row applies (A.53) and (A.57). We complete the proof of Lemma A.4. ∎
Proof of Proposition 1 for $j\in\mathcal{M}_{z}$..
According to MS24’s Proposition 3(c), the RE is bounded from below by $\frac{c}{s(\log p)^{4}}$ for some absolute constant $c$ w.p.a.1. In addition, by (A.49), the DB is $O_{p}\left(\dfrac{\log p}{n^{\tau/2}}\right)$ when $\mu$ follows the order in Proposition 1. Then by MS24’s Lemma 1
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}s\cdot\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}/(s(\log p)^{4})=\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}.$
This completes the proof of Proposition 1 for $j\in\mathcal{M}_{z}$. ∎
A.3 Proofs for Section 3.2
Lemma A.5.
Suppose Assumptions 1-5 hold. Then for any $j\in[p]$ ,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm sgn}(G_{j}^{*})$ (A.58)
where
[…]
with $\mathcal{B}_{j}$ being a Brownian motion of variance ${\rm var}\left(e_{j,t}\right)$, and $\eta_{j,t}^{(1)}$ defined in (A.54). In addition,
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}H_{j}$ (A.60)
where
$H_{j}=\begin{cases}1,&j\in\mathcal{M}_{x},\\ \frac{{\rm var}(\eta_{j,t}^{(1)})}{{\rm var}\left(w_{j,t}\right)}&j\in\mathcal{M}_{z}.\end{cases}$
Proof of Lemma A.5.
We first prove (A.58). The first step is to show
$\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.61)
By definition of the Slasso residual,
$\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}=u_{t}+W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}).$
Thus,
$\widehat{\sigma}_{u}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}\widehat{u}_{t}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}+\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}+\dfrac{2}{n}\sum_{t=1}^{n}u_{t}W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}).$ (A.62)
By MS24’s Theorem 3, we have
$\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}=\|n^{-1}W(\theta^{*}-\widehat{\theta}^{{\rm S}})\|_{2}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.63)
By the Cauchy-Schwartz inequality,
$\dfrac{2}{n}\sum_{t=1}^{n}u_{t}W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}})\leq 2\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}}\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.64)
Combining (A.62), (A.63), and (A.64), we have
$\widehat{\sigma}_{u}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}+o_{p}(1).$ (A.65)
Using a standard law of large number, we deduce
$\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}.$ (A.66)
(A.65) and (A.66) imply that $\widehat{\sigma}_{u}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}$. Then for (A.58) it suffices to show
$\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|/\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm sgn}(G_{j}^{*}).$
CASE I: $j\in\mathcal{M}_{x}$. Define $\check{r}_{j,t}=\zeta_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)},$ and by definition we have
$\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j}.$ (A.67)
Then
$\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}-\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}.$ (A.68)
Note that $w_{j,t}$ is unit root and $\zeta_{j,t}$ is mildly integrated. By the functional CLT for the case of unit root in Phillips and Lee, (2016, Lemma 3.2), we have
[…]
Note that in (A.36), we allow for a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(4C_{\sigma}s^{2}n^{1-\tau})$ with the absolute constant $C_{\sigma}$ in (A.36). Then
$\left|n^{-(1+\tau)}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\|D_{-j}^{-1}\widehat{\varphi}^{(j)}\|_{1}$
$\leq\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{\sup_{k\in[p],t\in[n]}|w_{k,t-1}|\cdot|w_{j,t-1}|}{n^{\tau}\min_{k\in[p]}\widehat{\sigma}_{k}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{n(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{n^{\tau}\cdot n^{1-\tau}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}=\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\to 0,$
where the second inequality applies (A.36), the fourth inequality applies (A.11), (A.12), (A.9), and (A.39). Thus,
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}G_{j}^{*}.$ (A.70)
By $\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j}$ and the continuous mapping theorem,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}=\dfrac{\left|\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}{\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\dfrac{\left|G_{j}^{*}\right|}{G_{j}^{*}}={\rm sgn}(G_{j}^{*}).$ (A.71)
Then (A.58) is implied by (A.61) and (A.71).
For (A.60), we have $\check{r}_{j,t}^{2}=\zeta_{j,t}^{2}+(W_{-j,t}^{\top}\widehat{\varphi}^{(j)})^{2}-2\zeta_{j,t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ When $j\in\mathcal{M}_{x}$,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $\leq\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}^{2}\cdot\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$
$\ \ \ \ +2\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\cdot\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}.$ (A.72)
By (A.11), (A.12), (A.40), and (A.41),
$\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$ $\leq\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\dfrac{1}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\dfrac{1}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}$
$=O_{p}\left((\log p)^{2}\right).$ (A.73)
Furthermore,
$\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}$ $\leq\dfrac{\|\dfrac{1}{n}\sum_{t=1}^{n}X_{-j,t-1}\zeta_{j,t}\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|\dfrac{1}{n}\sum_{t=1}^{n}Z_{-j,t-1}\zeta_{j,t}\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}+(\log p)^{1+\frac{1}{2r}}$
$=O\left(n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}\right).$ (A.74)
where the second inequality applies (A.15), (A.16), (A.11), and (A.12). Combining (A.72), (A.73), (A.74), and Proposition 1, we have
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $=O_{p}\left(n^{-\tau\wedge(1-\tau)}(\log p)^{4+1/r}+n^{-\frac{\tau\wedge(1-\tau)}{2}+(\tau-1/2)}(\log p)^{\frac{5}{2}+\frac{1}{r}}\right)$
$=o_{p}\left(n^{\tau}\right).$ (A.75)
In addition, (A.28) implies $\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}$. Then by (A.75),
$\left|\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}-1\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$
as $n\to\infty$, or equivalently
$\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}\bigg{/}\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.76)
Recall from (A.28) and (A.31) that $\frac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}$ and $\frac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}$ have the same probability limit, and thus
$\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\bigg{/}\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.77)
Recall that $\widehat{r}_{j,t}=\check{r}_{j,t-1}/\widehat{\varsigma}_{j}$ as shown in (A.67). Thus, (A.76) and (A.77) imply
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}/\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$
Then (A.60) is verified for CASE I.
CASE II: $j\in\mathcal{M}_{z}$. The definition of $\widehat{r}_{j,t}$ gives
$\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}+\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\left(\widehat{\varphi}^{(j)}-\varphi^{*(j)}\right),$ (A.78)
where $\widetilde{\eta}_{j,t}$ is defined below (3.10). Note that
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}(w_{j,t-1}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*}).$
By (A.52), we have
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}$ (A.79)
where $\eta_{j,t}=w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$. In addition, we deduce that
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\|_{\infty}$ $\leq\dfrac{\max_{k\in\mathcal{M}_{x}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|}{\min_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}}+\dfrac{\max_{k\in\mathcal{M}_{z}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}$
$=O_{p}\left(\sqrt{\dfrac{\log p}{n}}\cdot(\log p)^{1+\frac{1}{2r}}\right)+O_{p}(1)=O_{p}(1),$
where the second step follows (A.10) bounding the cross product between unit roots and a stationary component, (A.11) and (A.12) bounding the standard deviations, and the fact that
$\max_{k\in\mathcal{M}_{z}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|=O_{p}(1)$
following MS24’s Eq.(B.28). Thus, by Proposition 1 in our main text,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\left(\widehat{\varphi}^{(j)}-\varphi^{*(j)}\right)\right|$ $\leq\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{*(j)})\|_{1}$
$=o_{p}(1).$ (A.80)
Combining (A.78), (A.79), and (A.80),
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}.$ (A.81)
Thus,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{\left|{\rm cov}\left(w_{j,t},\eta_{j,t}\right)\right|}{{\rm cov}\left(w_{j,t},\eta_{j,t}\right)}={\rm sgn}\left({\rm cov}\left(w_{j,t},\eta_{j,t}\right)\right),$
which together with (A.61) implies (A.58) for $j\in\mathcal{M}_{z}$.
For (A.60), note that $\widehat{r}_{j,t}$ is the LASSO residual of regression (2.9). Recall that $\widetilde{\eta}_{j,t}$ is the error term of he pseudo-true model (3.10). Following the ideas in the proof of (A.65), we can show that $n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ and $n^{-1}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}$ share the same probability limit, which is
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}={\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}={\rm plim}_{n\to\infty}\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{\widehat{\zeta}_{j}^{2}}=\dfrac{{\rm var}\left(\eta_{j,t}\right)}{{\rm var}\left(w_{j,t}\right)},$
where the probability limit of the denominator follows (A.52). This completes the proof of Lemma A.5. ∎
We then prove the main results in Section 3.2
Proof of Theorem 1.
By the definition of the XDlasso estimator,
$\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}$ $=\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}.$
Then the $t$-statistic can be decomposed as
$\dfrac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}$ $=\mathcal{Z}_{j}+\varDelta_{j},\text{ where }$
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}},$
$\varDelta_{j}$ $=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}.$
We first bound $\varDelta_{j}$. By the Karush-Kuhn-Tucker condition, we can establish
$\|\sum_{t=1}^{n}D_{-j}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$ (A.82)
w.p.a.1 as in (3.12). Thus,
$\left|\varDelta_{j}\right|$ $\leq\sqrt{n}\dfrac{\left|n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}D_{-j}^{-1}D_{-j}(\theta_{-j}^{*}-\widehat{\theta}_{-j})\right|}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}$
$\leq\dfrac{\sqrt{n}\cdot\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\cdot\|D_{-j}(\theta{}_{-j}^{*}-\widehat{\theta}_{-j})\|_{1}}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\sqrt{n}}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\left(\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right)\cdot\left(\dfrac{s^{2}}{\sqrt{n}}(\log p)^{6+\frac{1}{2r}}\right)$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0$
where the third inequality applies (A.82) and Lemma 1, and the fourth inequality applies
$n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ (A.83)
implied by (A.60).
It then suffices to show that $\mathcal{Z}_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ Given the limit established (A.58), when $j\in\mathcal{M}_{z}$ it suffices to show
$\mathcal{Z}_{j}^{(1)}:=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.84)
When $j\in\mathcal{M}_{x}$, we need to additionally show the asymptotic distribution in (A.84) is independent of the Brownian motion in the $G_{j}^{*}$ of limit (A.58), so that
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\mathcal{Z}_{j}^{(1)}$
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}{\rm sgn}(G_{j}^{*})\cdot\mathcal{N}(0,1)=\mathcal{MN}\left(0,{\rm sgn}(G_{j}^{*})^{2}\right)=\mathcal{N}(0,1),$
where $\mathcal{MN}$ denotes a mixed normal distribution.
CASE I. When $j\in\mathcal{M}_{x}$,
$\mathcal{Z}_{j}^{(1)}=-\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\cdot$
We bound the first term by
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}}\to 0,$
where the first step applies (A.83), the second step applies (A.13) and Proposition 1.
As for the second term, by the central limit theorem in Kostakis et al., (2015, Lemma B4(ii)), the law of large numbers Eq.(21) in the supplement of the same reference, and the Slutsky’s Theorem, we have
$\frac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.85)
Besides, Phillips and Magdalinos, (2009)’s Lemma 3.2 shows that the asymptotic distribution in (A.85) is independent of the Brownian motion in the expression (A.59) of $G_{j}^{*}$. By (A.76) and the Slutsky’s Theorem, we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\cdot\dfrac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$
This completes the proof of (A.84) when $j\in\mathcal{M}_{x}.$
CASE II. When $j\in\mathcal{M}_{z}$, recall that we have defined $\eta_{j,t}=\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0,z}^{(j)*}$ in (A.44), with $\varphi_{0,z}^{(j)*}$ defined as (3.8). Then
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t}}{\widehat{\varsigma}_{j}}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}+\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}}{\widehat{\varsigma}_{j}},$
where the first equality is by the definition of $\widehat{r}_{j,t}$ in (2.9), the second row applies the pseudo-true regression model (3.10) and the equality (A.45), the third row applies (A.55), and the fourth row applies (A.43). Then by the definition of $\mathcal{Z}_{0,j}$ in (A.84), we have the following decomposition
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\left(\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}\right)u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}.$ (A.86)
We first bound the first term. By (A.60), we have $\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}\widetilde{W}_{-j,t-1}^{\top}\left(\widetilde{\varphi}-\widetilde{\varphi}^{*}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-h,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}\cdot\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0,$ (A.87)
where the second step applies MS24’s (B.63), and Proposition 1 in this current paper.
We then show the central limit theorem for the second term. Recall that $u_{t}$ is m.d.s. by Assumption 1, and $\eta_{j,t}^{(1)}$ is stationary and strong mixing. By a standard martingale central limit theorem we have $\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ By (A.60) and (A.52), we have $\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}(\eta_{j,t}^{(1)})$. Thus, by the Slutsky’s Theorem we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.88)
Finally, note that $\phi_{j,t-1}$ is mildly integrated. Again by Kostakis et al., (2015, Lemma B4(ii)) we have $\sum_{t=1}^{n}\phi_{j,t-1}u_{t}=O_{p}\left(n^{(\tau+1)/2}\right),$ and thus
$n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$
By (A.60) and (A.52), we further have $\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\dfrac{n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.89)
By (A.86), (A.87), (A.88), and (A.89), we deduce (A.84) when $j\in\mathcal{M}_{z}.$ This completes the proof of Theorem 1. ∎
Proof of Theorem 2.
Recall that
$\widehat{\omega}_{j}^{{\rm XD}}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}$
where $\check{r}_{j,t}=\widehat{r}_{j,t}\cdot\widehat{\varsigma}_{j}.$
CASE I. When $j\in\mathcal{M}_{x}$, the weak convergence (A.70) implies $(n^{-(1+\tau)}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1})^{-1}=O_{p}(1),$ and (A.76) and (A.28) implies $\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}=O_{p}(n^{1+\tau}).$ Then
$\widehat{\omega}_{j}^{{\rm XD}}=O_{p}\left(\dfrac{\sqrt{n^{1+\tau}}}{n^{1+\tau}}\right)=O_{p}\left(\dfrac{1}{n^{(1+\tau)/2}}\right).$
CASE II. When $j\in\mathcal{M}_{z}$, by (A.81) and (A.60),
$\widehat{\omega}_{j}^{{\rm XD}}=O_{p}\left(\dfrac{\sqrt{n}}{n}\right)=O_{p}\left(\dfrac{1}{\sqrt{n}}\right).$
We complete the proof of Theorem 2. ∎
Appendix B Additional Empirical Results
Figure B.1: S&P 500 monthly Return and Log Earnings-price Ratio
Table B.1: Persistence of S&P 500 Monthly Return and Log Earnings-Price Ratio
Sample Period S&P 500 Monthly Return Log Earnings-Price Ratio
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2024) 0.224 $<$0.01 0.993 0.081
Pre-1994 (Jan. 1960 - Dec. 1993) 0.255 $<$0.01 0.994 0.776
Post-1994 (Jan. 1994 - Apr. 2024) 0.193 $<$0.01 0.979 $<$0.01
Notes: The lag order for the ADF test is set as $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.
Figure B.2: Inflation and Unemployment Rate
Table B.2: Persistence of Inflation Rate and Unemployment Rate
Sample Period Inflation Rate Unemployment Rate
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2024) 0.621 0.021 1.000 0.081
Pre-Volcker (Jan. 1960 - July 1979) 0.640 0.220 0.986 0.406
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.623 $<$0.01 0.992 0.071
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2024) 0.543 $<$0.01 0.938 0.365
Notes: The lag order for the ADF test is set as $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.