### v3

# Zhan Gaoa, Ji Hyung Leeb, Ziwei Meic, Zhentao Shid

Abstract
LASSO inflicts shrinkage bias on estimated coefficients, which undermines asymptotic normality and invalidates standard inferential procedures based on the t-statistic. Given cross sectional data, the desparsified LASSO has emerged as a well-known remedy for correcting the shrinkage bias. In the context of high dimensional predictive regression, the desparsified LASSO faces an additional challenge: the Stambaugh bias arising from nonstationary regressors modeled as local unit roots. To restore standard inference, we propose a novel estimator called IVX-desparsified LASSO (XDlasso). XDlasso simultaneously eliminates both shrinkage bias and Stambaugh bias and does not require prior knowledge about the identities of nonstationary and stationary regressors. We establish the asymptotic properties of XDlasso for hypothesis testing, and our theoretical findings are supported by Monte Carlo simulations. Applying our method to real-world applications from the FRED-MD database, we investigate two important empirical questions: (i) the predictability of the U.S. stock returns based on the earnings-price ratio, and (ii) the predictability of the U.S. inflation using the unemployment.
Key words: Data-rich Environment, Forecast, Hypothesis testing, LASSO, Local Unit root, Shrinkage
JEL code: C22, C53, C55
1 Introduction
The evaluation of economic and financial predictability has attracted widespread interest from theoretical and applied researchers for many decades. In today’s era of big data, we have unprecedented access to a vast amount of digital information about the economy. Recent advancements in inference with high dimensional data have uncovered new empirical patterns in predictive practices using large-scale datasets with temporal features.
This paper aims at a plain quest: in a high dimensional linear predictive regression model, where the number of potential regressors is larger than the sample size, how can one conduct valid inference for a regressor of primary interest? No research has solved this question before. The challenges are twofold. First, predictive regressions were mainly studied in the low-dimensional context. A defining feature of predictive regression theory lies in persistent regressors (stambaugh1999predictive), which upend standard inference based on the standard $t$-statistics. Second, one must estimate the coefficient via some regularization methods to cope with high dimensionality. For example, when the underlying true regression model is sparse, LASSO (tibshirani1996regression) is the off-the-shelf method. It is well known that the LASSO estimator is biased toward zero due to absolute-value shrinkage and exhibits a nonstandard asymptotic distribution that is distinct from the normal distribution. If we intend to provide an asymptotically normally distributed estimator to facilitate standard statistical inference, we must simultaneously combat two evils: the Stambaugh bias due to persistent regressors and the shrinkage bias caused by the LASSO penalty.
The above diagnosis hints at a plausible solution path. In low-dimensional predictive regressions, phillips2009econometric’s IVX method leverages a self-generated instrument to alleviate the regressor persistence, thereby overcoming the Stambaugh bias. In high dimensional cross sectional regressions, zhang2014confidence’s desparsified LASSO (Dlasso) constructs a score vector for the parameter of interest and removes the shrinkage bias via an auxiliary regression. Each method provides an asymptotic normal estimator in its respective environment.
Can we combine these two methods into a single procedure to address inference in high dimensional predictive regressions? We find that the answer is both no and yes. “No” is in the sense that a naive combination of the two does not lead to desirable results. “Yes”, on the other hand, is established upon a deep understanding of the mechanisms of both components and their adaptation to the context. This research culminates in a new IVX-Desparsified LASSO (XDlasso) estimator that is free from both biases and has an asymptotic normal distribution.
With a predictor of interest in mind, the construction of XDlasso, detailed in Algorithm 1 in Section 2.3, is summarized as follows. First, a workhorse estimator is needed to lay the groundwork for a high dimensional predictive regression. When both nonstationary and stationary predictors are present in the regression, mei2022lasso has recently established the consistency of the standardized LASSO (Slasso), making it a natural candidate for the workhorse estimator. Beyond the pure unit roots considered in MS24, we extend the characterization of nonstationary regressors by generalizing our framework to allow for local unit root (LUR) processes. The convergence rates of Slasso for LURs align with those for pure unit roots. However, the technical proofs for LURs are more involved than those for pure unit roots. We address the complexity arising from high dimensional predictors with both LURs and stationary regressors, and derive the convergence rates of the initial Slasso estimator.
Second, the common practice of generating the instrument in IVX is insufficient — the IV must be scale-standardized to have the stochastic order aligned with all other predictors in Slasso. The standardized IV serves as the target variable for the auxiliary Slasso regression in Dlasso to estimate the shrinkage bias. The XDlasso estimator of the parameter of interest is defined as the initial Slasso estimator plus the bias-correction term, and the companion $t$-statistic is employed for statistical inference by comparing it to critical values from the standard normal distribution.
We further establish the asymptotic normality of our proposed XDlasso estimator and the convergence rate of its standard error. Specifically, the XDlasso estimator is $\sqrt{n}$-consistent for a stationary regressor while its convergence accelerates for an LUR regressor. Moreover, to conduct simultaneous inference for multiple parameters of interest, we develop a Wald statistic with an asymptotic $\chi^{2}$ distribution based on XDlasso. This Wald test is valid even when the parameters involve both stationary and nonstationary regressors.
To tackle persistent regressors, the self-generated IVX instrument in the second step is the key ingredient. The generated IV is less persistent than the nonstationary regressors modeled as LURs. This important feature enables us to decorrelate the IV from other covariates in the auxiliary LASSO regression, so that the resulting XDlasso estimator possesses these two properties: (i) It is free from the Stambaugh bias and thus enjoys asymptotic normality; (ii) It reduces the order of shrinkage bias to make it correctable. In contrast, the ordinary Dlasso encounters a spurious auxiliary regression, failing to correct the bias arising from persistent regressors (see Section 2.4 for details). More importantly, XDlasso inference does not require a priori knowledge of the persistence of the regressor of interest and is thus immune to pretesting bias. To the best of our knowledge, this is the first methodology to handle the inferential problem in high dimensional predictive regressions ($p\gg n$) with nonstationary predictors. This is also the first paper that extends the IVX technique into the high dimensional framework.
Monte Carlo simulations show that XDlasso successfully removes the bias for inference on the coefficient of a nonstationary regressor, but the ordinary Dlasso fails to do so. Our procedure is applied to the high dimensional macroeconomic FRED-MD dataset (mccracken2016fred) with both stationary and persistent variables, to study two important macro-finance problems: financial market return predictability and the Phillips curve in macroeconomics.
Literature review. With the advent of big data, machine learning methods have spread to time series topics such as nonstationarity (phillips2021boosting; smeekes2021automated; mei2024boosted), cointegration testing (onatski2018alternative; zhang2019identifying; bykhovskaya2022cointegration), and structural breaks (deshpande2023online; tu2023penetrating). This paper builds on several strands of literature. First, LASSO is one of the most studied methods in recent years, with well-developed theory in high dimension (bickel2009simultaneous). It is well received and used for economic applications; see belloni2012sparse, shi2016estimation, caner2018asymptotically, and babii2022machine, to name a few. In recent years, the properties of LASSO are studied in various topics in high dimensional time series, including nonstationary time series models (koo2020high; lee2022lasso) and inference based on the heteroskedasticity and autocorrelation consistent (HAC) estimation (babii2021highdimensionalgrangercausalitytests; babii2024high). None of these works has considered hypothesis testing problem for high dimensional predictive regressions with both LURs and stationary regressors.
Hypothesis testing after LASSO is challenging because of the shrinkage bias. To validate hypothesis testing in high dimensions, zhang2014confidence, van2014asymptotically, and javanmard2014confidence have developed the desparsified (debiased) LASSO estimators under the independently and identically distributed (i.i.d.) setting. adamek2022lasso generalize the Dlasso inference to high dimensional stationary time series. We follow this line of desparsified LASSO literature thanks to its convenience, which requires a baseline regression and an auxiliary regression only. On the other hand, chernozhukov2018’s double machine learning (DML) is a more general theoretical framework of debiased inference, widely used in cross sectional data where sample-splitting is readily implementable. However, none of the aforementioned works has devised any inferential procedure for high dimensional nonstationary time series. hecq2023inference apply a post-double selection procedure to test the Granger causality in high dimensional nonstationary vector autoregressive models with cointegrated data. In contrast, our procedure relies on desparsified LASSO without variable selection.
The other strand is the vast literature on predictive regressions. As highlighted by campbell2006efficient and jansson2006optimal, non-standard distortion in the asymptotic distribution arises from persistent regressors. The peculiar asymptotic distributions invalidate the standard inferential procedures. There have been multiple proposals for valid inference, for example, the Bonferroni method (campbell2006efficient), the conditional likelihood method (jansson2006optimal), the linear projection method (cai2014testing), the weighted empirical likelihood approach (zhu2014predictive; liu2019unified; yang2021unified), and the implication-based inference (xu2020testing). Some of these methods are designed for univariate predictive regressions; it would be difficult to extend them to the high dimensional case, where regularization is required to handle many parameters. On the other hand, phillips2009econometric’s IVX estimator gained its popularity by recovering asymptotic normality, enabling valid inference for mean regressions (kostakis2015robust; kostakis2018taking; phillips2013predictive; phillips2016robust; yang2020testing; demetrescu2023extensions) and quantile regressions (lee2016predictive; fan2019predictive; cai2023new; liu2023unified) with low dimensional regressors. IVX recovers asymptotic normality by projecting the persistent regressor onto a self-generated IV.
Layout. The rest of the paper is organized as follows. Section 2 introduces the high dimensional predictive regression model with a mixture of stationary and nonstationary regressors and proposes XDlasso. Section 3 establishes the theoretical results, justifying the size and power of the XDlasso inference procedure. Section 4 carries out simulation studies that corroborate the theory. Section 5 applies XDlasso inference to two macro-finance empirical examples. Technical proofs are relegated to the Online Appendices.
Notations. We set up the notation before the formal discussion. We define $\boldsymbol{1}\left\{\cdot\right\}$ as the indicator function, and $\Delta$ as the difference operator so that $\Delta x_{t}=x_{t}-x_{t-1}$. The set of natural numbers, integers, and real numbers are denoted as $\mathbb{N}$, $\mathbb{Z}$, and $\mathbb{R}$, respectively. For some $n\in\mathbb{N}$, the integer set $\{1,2,\cdots,n\}$ is denoted as $[n]$, and the space of $n$-dimensional vectors is denoted as $\mathbb{R}^{n}.$ For $x=(x_{t})_{t\in[n]}\in\mathbb{R}^{n}$, the $L_{1}$-norm is $\left\|x\right\|_{1}=\sum_{t=1}^{n}\left|x_{t}\right|$, and the sup-norm is $\|x\|_{\infty}=\sup_{t\in[n]}|x_{t}|$. Let $0_{n}$ be an $n\times 1$ zero vector, $1_{n}$ be an $n\times 1$ vector of ones, and $I_{n}$ be the $n\times n$ identity matrix. For a generic matrix $B,$ let $B_{ij}$ be the $(i,j)$-th element, and $B^{\top}$ be its transpose. Let $\|B\|_{\infty}=\max_{i,j}|B_{ij}|$, and $\lambda_{\min}(B)$ and $\lambda_{\max}(B)$ be the minimum and maximum eigenvalues, respectively. Define $a\wedge b:=\min\left\{a,b\right\}$, and $a\vee b:=\max\left\{a,b\right\}$. An absolute constant is a positive, finite constant that is invariant with the sample size. The abbreviation “w.p.a.1” is short for “with probability approaching one”. We use $\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}$ and $\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ to denote convergence in probability and in distribution, respectively. For any time series $\{a_{t}\}_{t=1}^{n}$, we use $\bar{a}$ to denote its sample mean $n^{-1}\sum_{t=1}^{n}a_{t}$. For any time series $\{a_{t}\}$ and $\{b_{t}\}$, we say they are asymptotically uncorrelated if their sample correlation coefficient $\frac{\sum_{t=1}^{n}(a_{t}-\bar{a})(b_{t}-\bar{b})}{\sqrt{\sum_{t=1}^{n}(a_{t}-\bar{a})^{2}\sum_{t=1}^{n}(b_{t}-\bar{b})^{2}}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$ as $n\to\infty.$
2 Model and Procedure
2.1 High Dimensional Predictive Regression
Suppose that a time series of the outcome $y_{t}$ is generated by the following linear predictive regression: [^1] For simplicity of exposition, an intercept in (2.1) is omitted, without loss of generality. As explained by MS24, the intercept in LASSO can be handled by the well-known Frisch-Waugh-Lovell theorem. In practical implementation — throughout all simulations and empirical exercises in this paper — we keep an unpenalized intercept in the model.
$y_{t}$ $=W_{t-1}^{\top}\theta^{*}+u_{t}=X_{t-1}^{\top}\beta^{*}+Z_{t-1}^{\top}\gamma^{*}+u_{t},$ (2.1)
where the error term $u_{t}$ is a stationary martingale difference sequence (m.d.s.) with mean zero and conditional variance $\sigma_{u}^{2}$. We consider two types of regressors with different stochastic properties. Firstly, the $p_{x}\times 1$ vector $X_{t}=(x_{1,t},\cdots,x_{p_{x},t})^{\top}$ collects the LURs:
$x_{j,t}=\rho_{j}^{*}x_{j,t-1}+e_{j,t}\text{ for }j=1,2,...,p_{x},$ (2.2)
where $e_{t}=(e_{1,t},\cdots,e_{p_{x},t})^{\top}$ is a $p_{x}$-dimensional vector of stationary time series. The AR(1) coefficient $\rho_{j}^{*}$ in (2.2) is close to 1 when the sample size $n$ is large, specified as
$\rho_{j}^{*}=1+\frac{c_{j}^{*}}{n}\text{ for }j=1,2,...,p_{x},$ (2.3)
where $c_{j}^{*}\in\mathbb{R}$ is allowed to be negative, positive, or zero. Therefore, our framework accommodates nonstationary regressors that are locally integrated ($c_{j}^{*}<0$), unit roots ($c_{j}^{*}=0$), and locally explosive ($c_{j}^{*}>0$). Secondly, stationary regressors are stored in the $p_{z}\times 1$ vector $Z_{t}=(z_{1,t},\cdots,z_{p_{z},t})^{\top}$. The two types of regressors are combined into a long vector $W_{t}=(X_{t}^{\top},Z_{t}^{\top})^{\top}=(w_{1,t},\cdots,w_{p,t})^{\top}$ of $p$ ($p_{x}+p_{z}$) elements, and the associated coefficients are placed into $\theta^{*}=(\beta^{*\top},\gamma^{*\top})^{\top}\in\mathbb{R}^{p}$. Following the literature, we refer to $W_{t}$, which has multiple degrees of persistence, as mixed root regressors. For simplicity, let the initial value $\|W_{t=0}\|_{\infty}=O_{p}(1)$. Define the sparsity index $s=\sum_{j=1}^{p}\boldsymbol{1}\{\theta_{j}^{*}\neq 0\}$ as the number of nonzero components in the coefficient vector $\theta^{*}$.
As in the default R program option glmnet::glmnet(x,y), it is a common practice in LASSO to scale-standardize each regressor $w_{j,t}$ by its sample standard deviation (s.d.) $\widehat{\sigma}_{j}=\sqrt{\frac{1}{n}\sum_{t=1}^{n}(w_{j,t-1}-\bar{w}_{j})^{2}},$ where $\bar{w}_{j}=n^{-1}\sum_{t=1}^{n}w_{j,t-1}$ is the sample mean. Let the diagonal matrix $D={\rm diag}\left(\widehat{\sigma}_{1},\widehat{\sigma}_{2},\cdots,\widehat{\sigma}_{p}\right)$ store the sample standard deviations. The standardized LASSO (Slasso) estimator is
$\widehat{\theta}^{{\rm S}}:=\arg\min_{\theta\in\mathbb{R}^{p}}\dfrac{1}{n}\sum_{t=1}^{n}(y_{t}-W_{t-1}^{\top}\theta)^{2}+\lambda\|D\theta\|_{1}.$ (2.4)
The Slasso estimator is scale-invariant: if the regressor $w_{j,t-1}$ is multiplied by a nonzero constant $m$, then the $j$-th coefficient estimator changes proportionally to $\widehat{\theta}_{j}^{{\rm S}}/m.$ The standardization renders the magnitudes of LURs into the same order as those of stationary regressors, so that the same LASSO tuning parameter $\lambda$ in (2.4) is valid for both stationary and persistent regressors. In contrast, the plain LASSO (Plasso) with the matrix $D$ in (2.4) replaced with the identity matrix is scale-variant. What is worse, equipped with a single tuning parameter $\lambda$, Plasso favors the LURs of a larger order, and shrinks the coefficients of stationary regressors with a smaller order all the way to zero — thus becoming inconsistent.
Remark 1.
We follow the default option of the statistical software to use the s.d. for scaling in the Slasso estimator (2.4). The purpose of scaling is to ensure consistency of Slasso (2.4). The long-run variance, which requires tuning a bandwidth in its estimation, provides no additional benefit. We therefore prefer and stick to the vanilla s.d.
2.2 Two Types of Biases
When data are i.i.d. or stationary, LASSO is subject to shrinkage bias as well as a nonstandard asymptotic distribution, which cannot be used for standard inference (fu2000asymptotics). This motivates zhang2014confidence to bring forth the desparsified LASSO to correct the bias and recover asymptotic normality. With high dimensional LURs, $\widehat{\theta}^{{\rm S}}$ is subject to not only the shrinkage bias, but also the Stambaugh bias due to persistence, which further distorts the standard t-statistics inference. In this section, we examine both the shrinkage bias and the Stambaugh bias, and propose XDlasso for correcting both biases.
We are interested in inference on a null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ for a $j\in[p]$, a prevalent practice in empirical studies. In a low dimensional linear regression where $p$ is fixed, the Frisch-Waugh-Lovell theorem yields the following formulation of the ordinary least squares (OLS) estimator:
$\widehat{\theta}_{j}^{{\rm OLS}}=\frac{\sum_{t=1}^{n}w_{j,t-1}^{\perp}y_{t}}{\sum_{t=1}^{n}w_{j,t-1}^{\perp}w_{j,t-1}},$
where $w_{j}^{\perp}=(w_{j,0}^{\perp},\dots,w_{j,n-1}^{\perp})^{\top}$ is the OLS residual from regressing $w_{j,t}$ on all other regressors $W_{-j,t}=(w_{k,t})_{k\neq j}$. OLS induces a large variance as $p$ gets large, and becomes infeasible when $p>n$. Now, consider replacing the OLS residual $w_{j}^{\perp}$ by a generic score vector $r_{j}=(r_{j,0},\dots,r_{j,n-1})^{\top}$ to construct an estimator of $\theta_{j}^{*}$ that is linear in $y_{t}$ in the form of $\widehat{\theta}_{j}^{({\rm lin})}=\frac{\sum_{t=1}^{n}r_{j,t-1}y_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}.$ Since
$y_{t}=w_{j,t-1}\theta_{j}^{*}+W_{-j,t-1}^{\top}\theta_{-j}^{*}+u_{t}$
where $\theta_{-j}^{*}=(\theta_{k}^{*})_{k\neq j}$ is the vector of coefficients excluding the $j$-th entry, the generic estimator $\widehat{\theta}_{j}^{({\rm lin})}$ can be decomposed into
$\widehat{\theta}_{j}^{({\rm lin})}=\theta_{j}^{*}+\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\theta_{-j}^{*}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}=:\theta_{j}^{*}+{\rm N}_{j}+{\rm B}_{j},$
where ${\rm N}_{j}$ is the noise component that determines the asymptotic distribution of $\widehat{\theta}_{j}^{({\rm lin})}$, and ${\rm B}_{j}$ is the potential bias due to the choice of $r_{j}$.
For OLS, the score vector $r_{j}=w_{j}^{\perp}$ is orthogonal to the column space of $W_{-j,\cdot}:=(W_{-j,0},\dots,W_{-j,n-1})^{\top}$, under which ${\rm B}_{j}=0$ and no bias is present. The bias term ${\rm B}_{j}$ pops up whenever $r_{j}$ is not orthogonal to $W_{-j,\cdot}$, which happens if we add a penalty to the OLS objective function. With the LASSO penalty at place, we call ${\rm B}_{j}$ shrinkage bias.
Following zhang2014confidence, we replace the unknown parameter $\theta_{-j}^{*}$ by the feasible workhorse estimator $\widehat{\theta}_{-j}^{{\rm S}}=(\widehat{\theta}_{k}^{{\rm S}})_{k\neq j}$ to obtain
$\widehat{\theta}_{j}=\widehat{\theta}_{j}^{({\rm lin})}-\widehat{{\rm B}}_{j}\text{ where }\widehat{{\rm B}}_{j}=\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}\widehat{\theta}_{-j}^{{\rm S}}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}$ (2.5)
to compensate for ${\rm B}_{j}$. Equivalently, $\widehat{\theta}_{j}$ can be written as
$\widehat{\theta}_{j}=\widehat{\theta}_{j}^{{\rm S}}+\frac{\sum_{t=1}^{n}r_{j,t-1}\widehat{u}_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}},$ (2.6)
where $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ is the Slasso residual. Though LASSO may shrink $\widehat{\theta}_{j}^{{\rm S}}$ all the way to exactly zero, the second term in (2.6) is continuously distributed and therefore $\widehat{\theta}_{j}$ is a desparsified version of $\widehat{\theta}_{j}^{{\rm S}}$. A straightforward calculation yields
$\widehat{\theta}_{j}-\theta_{j}^{*}=\frac{\sum_{t=1}^{n}r_{j,t-1}u_{t}}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}r_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j}^{{\rm S}})}{\sum_{t=1}^{n}r_{j,t-1}w_{j,t-1}}={\rm N}_{j}-(\widehat{{\rm B}}_{j}-{\rm B}_{j}),$ (2.7)
where $\widehat{{\rm B}}_{j}-{\rm B}_{j}$ is the approximation error of the shrinkage bias.
Let $\omega_{j}$ denote the standard deviation of ${\rm N}_{j}$. To secure asymptotic normality for $\widehat{\theta}_{j}$, we need an appropriate score vector $r_{j}$ such that as $n\to\infty$:
(R1) ${\rm N}_{j}/\omega_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1),$ and (R2) $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$.
These two results will furnish $\widehat{\theta}_{j}$ with asymptotic normality, validating inference based on the standard $t$-statistic.
The result (R1) is well established by zhang2014confidence under i.i.d. data when the score $r_{j,t}$ is taken as the LASSO residual of regressing $w_{j,t}$ on all other regressors. However, when $w_{j,t}$ is an LUR, this practice leads to a highly persistent $r_{j,t}$ and ruins the asymptotic normality of ${\rm N}_{j}$. This is the Stambaugh bias — a highly persistent regressor produces an asymptotically non-normally distributed OLS estimator skewed away from zero. To safeguard (R1), we must seek a score that is less persistent than an LUR.
To retain (R2), we again examine the estimation error of the shrinkage bias
${\rm\widehat{B}}_{j}-{\rm B}_{j}$ $=g_{j}^{\top}(\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}),\ \text{where}\ g_{j}=\dfrac{\sum_{t=1}^{n}W_{-j,t-1}r_{j,t-1}}{\sum_{t=1}^{n}w_{j,t-1}r_{j,t-1}}\cdot$ (2.8)
Note that $|{\rm\widehat{B}}_{j}-{\rm B}_{j}|\leq\|g_{j}\|_{\infty}\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}.$ Slasso’s $L_{1}$ estimation error $\|\widehat{\theta}_{-j}^{{\rm S}}-\theta_{-j}^{*}\|_{1}$ is invariant with the choice of score vector $r_{j,t}$, and diminishes if $\widehat{\theta}^{{\rm S}}$ is consistent. Thus, it suffices to control the order of $\|g_{j}\|_{\infty}$ to achieve (R2). The explicit expression of $g_{j}$ in (2.8) suggests that a weak correlation in sup-norm between $W_{-j,t}$ and the score $r_{j,t}$ (relative to the correlation between $w_{j,t}$ and $r_{j,t}$) helps.
With these routes in mind, we devise XDlasso in the following section. It proceeds with two key steps: (i) conducting an IVX transformation to get a new variable less persistent than the LUR regressors to eliminate the Stambaugh bias; (ii) running an auxiliary LASSO regression to construct an $r_{j,t}$ to remove the shrinkage bias.
2.3 IVX-Desparsified LASSO
For inference of $\theta_{j}^{*}$ in low dimensions, phillips2009econometric’s IVX method generates an instrument by quasi-differencing $w_{j,t}$. When $w_{j,t}$ is an LUR, this self-generated instrument is mildly integrated. The mitigation of persistence will remove the Stambaugh bias when the sample size passes to infinity, and thus recovers asymptotic normality of the IVX estimator. Due to the coexistence of Stambaugh bias and shrinkage bias, the wisdom of IVX cannot be directly transplanted into the high dimensional case. Instead, we integrate IVX with the idea of desparsified LASSO after comprehending the mechanism of the biases illustrated in Section 2.2.
Specifically, IVX adopts the following instrumental variable:
$\zeta{}_{j,t}=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}\Delta w_{j,s}$ (2.9)
where $\rho_{\zeta}\in(0,1)$ is a user-determined tuning parameter. Define the s.d. of the instrument as $\widehat{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=0}^{n-1}\left(\zeta_{j,t}-\bar{\zeta}_{j}\right)^{2}.}$ We unify the scale for LURs and stationary regressors by standardizing the instrument with its s.d.:
$\widetilde{\zeta}_{j,t}=\zeta_{j,t}/\widehat{\varsigma}_{j}.$ (2.10)
For low dimensional predictive regressions, the IVX literature has established asymptotic normality of ${\rm N}_{j}$ in (2.7) taking $r_{j,t}=\widetilde{\zeta}_{j,t}$; see phillips2009econometric. When $w_{j,t}$ is an LUR, the IV $\zeta_{j,t}$ is mildly integrated and less persistent than an LUR. Furthermore, recall that the vector of regressors $W_{-j,t}$ includes either LURs or stationary regressors. Due to different degrees of persistence, the mildly integrated IV $\widetilde{\zeta}_{j,t}$ and the regressors $W_{-j,t}$ are asymptotically uncorrelated. Thus, when $w_{j,t}$ is an LUR, we can choose the score vector as $r_{j,t}=\widetilde{\zeta}_{j,t}$ to deliver a small order of $\|g_{j}\|_{\infty}$. If $w_{j,t}$ is stationary, however, this score vector fails. When $n$ is large, the instrumental variable $\zeta_{j,t}$ behaves similarly as the stationary regressor $w_{j,t}$, and thus its correlation to the high dimensional stationary regressors in $W_{-j,t}$ is not negligible.
The above analysis implies that we must decorrelate the score vector with the other regressors $W_{-j,t}$ to reduce the order of $g_{j}$ to control for the approximation error $\widehat{{\rm B}}_{j}-{\rm B}_{j}$. This decorrelation will produce a unified testing approach for both LURs and stationary regressors. To this end, we construct a residual score vector $\widehat{r}_{j}=(\widehat{r}_{j,0},\cdots,\widehat{r}_{j,n-1})^{\top}$ by the following auxiliary LASSO regression
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)},\ \text{where}$ (2.11)
$\widehat{\varphi}^{(j)}$ $=\arg\min_{\varphi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{\zeta}_{j,t-1}-W_{-j,t-1}^{\top}\varphi)^{2}+\mu_{j}\|D_{-j}\varphi\|_{1}$ (2.12)
with the LASSO tuning parameter $\mu_{j}$ and $D_{-j}={\rm diag}(\{\widehat{\sigma}_{k}\}_{k\neq j})$. In low dimensional multivariate predictive regressions, IVX transforms each regressor into a less persistent instrumental variable parallel to (2.9), and constructs a two-stage least squares estimator using all the self-generated instrumental variables. In contrast, we only transform the variable of interest $w_{j,t}$ and estimate one auxiliary regression (2.12).
The score vector $\widehat{r}_{j}$ in (2.11) accommodates stationary and LUR regressors. Recall that when $w_{j,t}$ is stationary, the magnitude of the instrument $\zeta{}_{j,t}$ behaves similarly as $w_{j,t}$. Thus, the score $\widehat{r}_{j}$ is asymptotically equivalent to the standardized residual of the LASSO regression of $w_{j,t}$ on $W_{-j,t}$. The latter is proportional to the score in zhang2014confidence for cross-sectional data, which is also used in adamek2022lasso for stationary time series. When $w_{j,t}$ is an LUR, the instrument $\zeta_{j,t}$ is mildly integrated and has a different degree of persistence from $W_{-j,t}$. Therefore, $\zeta_{j,t}$ is asymptotically uncorrelated to each regressor in $W_{-j,t}$. For notational conciseness, define $\widetilde{W}_{-j,t}=(D_{-j})^{-1}W_{-j,t}$ and $\widetilde{\varphi}^{(j)}=D_{-j}\widehat{\varphi}^{(j)}$. The analysis above suggests $\widetilde{\varphi}^{(j)}\approx 0$ so that $\widehat{r}_{j,t}=\widetilde{\zeta}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\varphi}^{(j)}\approx\widetilde{\zeta}_{j,t}$. Recall from the discussion right after (2.10) that the standardized instrument $\widetilde{\zeta}_{j,t}$ is a valid score process for an LUR regressor, and thus the score $\widehat{r}_{j,t}$ can also remove the biases asymptotically. As a result, the residual of the auxiliary LASSO regression provides a unified construction of the score for either stationary or nonstationary $w_{j,t}$. It allows practitioners to conduct hypothesis testing on the coefficient in high dimensional predictive regression regardless of the order of integration of $w_{j,t}$. We maintain an agnostic attitude about the persistence of the regressors; in practical implementation, we need not distinguish these two types.
Remark 2.
The dependent variable of the auxiliary LASSO regression (2.12) is the standardized instrumental variable (2.10). It is possible to use its original form (2.9). However, the s.d. of $\zeta_{j,t}$ passes to infinity if $w_{j,t}$ is an LUR. Thus, without standardizing the instrument, the theoretical order of the tuning parameter $\mu_{j}$ in (2.12) for an LUR $w_{j,t}$ would be much larger than the order for a stationary regressor. This difference complicates the theoretical justifications. The standardization in (2.10) unifies the convergence rate of $\mu_{j}$ regardless of $w_{j,t}$ being an LUR or a stationary regressor.
Remark 3.
The central idea of debiasing technique via orthogonalization in high dimensional regressions is shared by the van2014asymptotically, javanmard2014confidence and chernozhukov2018. DML by chernozhukov2018 provides a more general framework by allowing nonlinear and semiparametric models. Given cross sectional data, DML advocates using sample splitting to eliminate $c^{*}$ (in the notation of chernozhukov2018), even though empirical process theory implies that $c^{*}$ should vanish asymptotically. In other words, the quality of finite sample approximation using the sample splitting technique is much better than without it. Justifying sample splitting in time series is not as straightforward as that in cross sectional data (beutner2021justification; adamek2022lasso, Remark 4). Without sample splitting, the theory of empirical processes with nonstationary data will be challenging, and this topic deserves thorough investigations in future research.
Following (2.6), XDlasso is constructed as
$\widehat{\theta}_{j}^{{\rm XD}}=\widehat{\theta}_{j}^{{\rm S}}+\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\widehat{u}_{t}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ (2.13)
with the standard error
$\widehat{\omega}_{j}^{{\rm XD}}=\dfrac{\widehat{\sigma}_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}$ (2.14)
where $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. For simplicity, we focus on homoskedastic errors in the main text. In Section B.3 of the appendix, we provide a formula of the heteroskedasticity-robust standard error, and show that our method is robust to conditional heteroskedasticity in both Monte Carlo simulations (Section B.3) and empirical applications (Section C.2).
With the point estimator (2.13) and the associated standard error in (2.14), we perform the $t$-test for the null hypothesis $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$. We summarize the testing procedure using the XDlasso estimator below.
Algorithm 1 (XDlasso Inference for $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$).
Step1 Obtain $\widehat{\theta}^{{\rm S}}$ from the Slasso regression (2.4). Save the residual $\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}$ and $\widehat{\sigma}_{u}^{2}=n^{-1}\sum_{t=1}^{n}\widehat{u}_{t}^{2}$. Step2 Obtain the IV $\zeta_{j,t}$ by the transformation (2.9), and standardize it by (2.10). Step3 Run the auxiliary LASSO regression (2.12), and save the residual $\widehat{r}_{j,t}$ in (2.11). Step4 Compute the XDlasso estimator (2.13) and the standard error (2.14). Step5 Obtain the $t$-statistic $t_{j}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{0,j})\big/\widehat{\omega}_{j}^{{\rm XD}}.$ (2.15) Reject $\mathbb{H}_{0}$ under the significance level $\alpha$ if $|t_{j}^{{\rm XD}}|>\Phi_{1-\alpha/2}$, where $\Phi_{1-\alpha/2}$ is the $100(1-\alpha/2)$-th percentile of the standard normal distribution.
The testing procedure in Algorithm 1 refines the conventional predictive regression inference using modern high dimensional inference techniques. In the following, we elaborate on the necessity of the IVX transformation by explaining the drawback of Dlasso under high dimensional LURs.
2.4 Necessity of IVX Transformation
Given i.i.d. data, Dlasso proceeds with the following residual as the score vector:
$r_{j,t}$ $=\widetilde{w}_{j,t}-W_{-j,t}^{\top}\widehat{\psi}^{(j)}\text{ where }\text{$\widetilde{w}_{j,t}=w_{j,t}/\widehat{\sigma}_{j}$,}$ (2.16)
$\widehat{\psi}^{(j)}$ $=\arg\min_{\psi\in\mathbb{R}^{p-1}}\dfrac{1}{n}\sum_{t=1}^{n}(\widetilde{w}_{j,t-1}-W_{-j,t-1}^{\top}\psi)^{2}+\mu_{j}\|D_{-j}\psi\|_{1}.$ (2.17)
where the dependent variable in (2.16) is also scaled to be kept in line with (2.11) and (2.12). For simplicity, in this subsection we temporarily restrict all nonstationary regressors to be pure unit roots ($\rho_{j}^{*}=1$). We show that the standard Dlasso procedure fails to correct the bias in this special case of LURs, thereby invalidating the inference in general cases.
First, (2.17) is a spurious regression as in granger1974spurious. In a low dimensional regression where all regressors have unit roots, the standardized least squares counterpart of the regression (2.17) follows a functional central limit theorem (FCLT)
$\widetilde{\psi}^{(j,{\rm OLS})}$ $=$ $D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{w}_{j,t-1}$ (2.18)
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}$ $\widetilde{\psi}^{*}:={\rm diag}(\{\sigma_{k}^{*}/\sigma_{j}^{*}\}_{k\neq j})\left(\int_{0}^{1}\mathcal{B}_{-j}\mathcal{B}_{-j}^{\top}\right)^{-1}\int_{0}^{1}\mathcal{B}_{-j}\mathcal{B}_{j},$
where $\mathcal{B}=(\mathcal{B}_{1},\dots,\mathcal{B}_{p})^{\top}$ is a $p$-dimensional Brownian motion, $\sigma_{j}^{*}=\sqrt{\int_{0}^{1}\mathcal{B}_{j}^{2}-(\int_{0}^{1}\mathcal{B}_{j})^{2}}$, and $\mathcal{B}_{-j}=(\mathcal{B}_{k})_{k\neq j}$. In other words, the OLS estimator converges in distribution to a nondegenerate random variable. This is the well-known spurious regression phenomenon in unit root regressions. While (2.18) suggests that the limit target coefficients are random and continuously distributed, the asymptotics for the LASSO estimator (2.17) in high dimensions depends on the sparsity of the target coefficients. The randomness of coefficients is contradictory to the sparsity required in LASSO.
Second, the score vector in (2.16) cannot remove the Stambaugh bias. Even in low dimensions, the least squares residual $\widehat{r}_{j,t}^{{\rm OLS}}=\widetilde{w}_{j,t}-\widetilde{W}_{-j,t}^{\top}\widetilde{\psi}^{(j,{\rm OLS})}$remains highly persistent due to spurious regression. Using $\widehat{r}_{j,t}^{{\rm OLS}}$ as the score vector in low dimensions would therefore keep the Stambaugh bias in the noise component ${\rm N}_{j}$. This issue in low dimension is also present in the score vector $r_{j,t}$ construction in (2.16) by high dimensional LASSO.
Figure 1: Histograms of $t$-statistics from XDlasso and Dlasso
In contrast, with the help of the IVX transformation, the score vector in (2.11) achieves (R1) and (R2). Figure 1 provides an illustrative simulation to compare Dlasso with XDlasso. We set $\beta_{1}^{*}$ (the coefficient of the first unit root) as zero, $n=300$, $(p_{x},p_{z})=(150,300)$, and use i.i.d. innovations (See Eq. (4.1)). We generate the data following the DGP (2.1), with a mixture of stationary and nonstationary regressors as in Section 4.1. Figure 1 displays the histograms of the $t$-statistics over 2000 replications. The density of XDlasso $t$-statistic is well approximated by $\mathcal{N}(0,1)$, whereas the Dlasso $t$-statistic suffers from a substantial bias.
2.5 Joint Inference for Low-Dimensional Coefficients
We have devised the XDlasso inference for a scalar coefficient $\theta_{j}^{*}$. In low dimensional predictive regression, the IVX estimator is applicable to multiple coefficients, which jointly follows an asymptotic multivariate normal distribution. Therefore, a Wald statistic by IVX is available to jointly test the predictability of multiple regressors, provided there are a finite number of them. This test statistic is shown to be valid, even when the parameters of interest involve both stationary and nonstationary regressors.
The validity of Wald test extends to XDlasso in high dimensional predictive regression. Specifically, suppose that we are interested in a subset of regressors indexed by $\mathcal{J}\subset[p]$ with a fixed cardinality $|\mathcal{J}|$. In this case, the XDlasso estimators $\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}=(\widehat{\theta}_{j}^{{\rm XD}})_{j\in\mathcal{J}}$ defined in (2.13) asymptotically follow a multivariate normal distribution. Let the null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$ involve $|\mathcal{J}|$ restrictions, where $\theta_{0,\mathcal{J}}$ is a $|\mathcal{J}|$-dimensional vector. We construct the following Wald statistic
${\rm Wald}_{\mathcal{J}}^{{\rm XD}}=(\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}-\theta_{0,\mathcal{J}})^{\top}[\widehat{\Omega}_{\mathcal{J}}^{{\rm XD}}]^{-1}(\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}-\theta_{0,\mathcal{J}}),$ (2.19)
where $\widehat{\Omega}_{\mathcal{J}}^{{\rm XD}}$ estimates the covariance matrix of $\widehat{\theta}_{\mathcal{J}}^{{\rm XD}}$, with its $(j,k)$-th entry
$\widehat{\Omega}_{j,k}^{{\rm XD}}=\widehat{\sigma}_{u}^{2}\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\widehat{r}_{k,t-1}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\cdot\sum_{t=1}^{n}\widehat{r}_{k,t-1}w_{k,t-1}}$
measuring the covariance between $\widehat{\theta}_{j}^{{\rm XD}}$ and $\widehat{\theta}_{k}^{{\rm XD}}$. Under the null hypothesis, the Wald statistic in (2.19) will follow an asymptotic $\chi^{2}$ distribution with the degree of freedom $|\mathcal{J}|$, enabling joint inference on $\theta_{\mathcal{J}}^{*}$.
3 Asymptotic Theory
This section develops the limit theory to shed light on the asymptotic behaviors of the XDlasso estimator. Unsurprisingly, this paper’s assumptions share similarities with those in MS24. We state our theoretical assumptions and then highlight and explain the differences between the assumptions in these two papers. Regarding the asymptotic framework, we define the number of regressors $p=p\left(n\right)$ and the sparsity index $s=s\left(n\right)$ as deterministic functions of the sample size $n$. In asymptotic statements, we will explicitly send $n\to\infty$, and it is understood that $p(n)\to\infty$ as $n\to\infty$, while $s\left(n\right)$ is allowed to be either fixed or divergent. Recall that the stationary vector $e_{t}\in\mathbb{R}^{p_{x}}$ is the innovation of the LUR regressors. We assume that the stationary high dimensional vector $v_{t}=(e_{t}^{\top},Z_{t}^{\top})^{\top}$ is generated by the innovations $\varepsilon_{t}=(\varepsilon_{k,t})_{k\in[p]}$ via a linear transformation
$v_{t}=\Phi\varepsilon_{t},$ (3.1)
where $\Phi$ is a $p\times p$ deterministic matrix. Let $\mathcal{F}_{t}$ denote the $\sigma$-field generated by $\{u_{s},\varepsilon_{s}\}_{s\leq t}$.
Assumption 1.
Suppose that $u_{t}$ and $\varepsilon_{t}$ are strictly stationary. Moreover, $u_{t}$ is a martingale difference sequence (m.d.s.) such that $\mathbb{E}\left(u_{t}|\mathcal{F}_{t-1}\right)=0$ and $\mathbb{E}(u_{t}^{2}|\mathcal{F}_{t-1})=\sigma_{u}^{2}>0$. There exist absolute constants $C_{u}$, $b_{u}$, $C_{\varepsilon}$, and $b_{\varepsilon}$ such that for all $t\in\mathbb{Z}$ and $a>0$,
$\Pr\left(|u_{t}|>a\right)$ $\leq C_{u}\exp(-a/b_{u}),$ (3.2)
$\Pr\left(|\varepsilon_{k,t}|>a\right)$ $\leq C_{\varepsilon}\exp(-a/b_{\varepsilon}),\ \ \forall k\in[p].$ (3.3)
Furthermore, $\{\varepsilon_{k,t}\}_{t\in\mathbb{Z}}$ and $\{\varepsilon_{\ell,t}\}_{t\in\mathbb{Z}}$ are independent for all $k\neq\ell.$
Remark 4.
In Assumption 1, Eqs. (3.2) and (3.3) impose the sub-exponential tails for the innovations, which includes the familiar sub-Gaussian tail as a special case. The sub-exponential condition is also imposed by MS24 and it is needed in Section 3.1 to deduce the restricted eigenvalue and deviation bound under high dimensional nonstationary data. The heavy-tail features of financial data like extreme returns are not covered in the current paper, and will be an important extension in future studies.
The following Assumption 2 imposes restrictions on the $\alpha$-mixing coefficients that characterize the time dependence of the innovations $u_{t}$ and $\varepsilon_{t}$. For any two $\sigma$-fields $\mathcal{A}$ and $\mathcal{B}$, define $\alpha(\mathcal{A},\mathcal{B})=\sup_{A\in\mathcal{A},B\in\mathcal{B}}|\Pr\left(A\cap B\right)-\Pr(A)\Pr(B)|$ and $\alpha(d)=\sup_{s\in\mathbb{Z}}\alpha(\sigma(\{u_{t},\varepsilon_{t}\}_{t\leq s}),\sigma(\{u_{t},\varepsilon_{t}\}_{t\geq s+d}))$.
Assumption 2.
There exist some absolute constants $C_{\alpha}$, $c_{\alpha}$, $r$, $c_{\varepsilon}$ such that
$\alpha(d)\leq C_{\alpha}\exp\left(-c_{\alpha}d^{r}\right),\ \ \forall d\in\mathbb{Z},$ (3.4)
and the long-run variance $\mathbb{\mathbb{E}}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]\geq c_{\varepsilon}$ for all $k\in[p].$
The following Assumption 3 depicts the contemporary correlation of $v_{t}$ defined as (3.1), as well as the constants in the local-to-unity AR coefficients specified in (2.3). Define $\Omega=\Phi\Phi^{\top}$ where $\Phi$ has appeared in (3.1).
Assumption 3.
There are absolute constants $\underline{c}$ and $\overline{C}$ such that: (a) $\underline{c}\leq\lambda_{\min}(\Omega)\leq\lambda_{\max}(\Omega)\leq\overline{C}$; (b) $\max_{j\in[p]}\sum_{\ell=1}^{p}\left|\Phi_{j\ell}\right|\leq\overline{C}$; (c) $\max_{j\in[p_{x}]}|c_{j}^{*}|\leq\overline{C}.$
We specify the user-determined parameter in (2.9) as
$\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ (3.5)
with absolute constants $C_{\zeta}>0$ and $\tau\in(0,1)$. The choice of $\tau$ determines the persistence of the IV $\zeta_{j,t}$, which will be elaborated in Remark 8. The following Assumption 4 characterizes the number of regressors $p$ and the sparsity index $s$ relative to the sample size $n$.
Assumption 4.
(a) $p=O(n^{\nu})$ for an arbitrary $\nu>0$ and (b) $s=O(n^{\frac{1}{4}(\tau\wedge(1-\tau))-\xi}\wedge p^{1-\xi})$ for an arbitrary small $\xi>0$.
In the following, we expound the differences between the assumptions above and those in MS24. First, Assumption 1 imposes the m.d.s. and conditional homoskedasticity conditions for the error term $u_{t}$. Three remarks are in order to justify these two conditions.
Remark 5.
Although the m.d.s. assumption is not required in MS24, in this paper it is essential for the asymptotic normality of XDlasso. Without the m.d.s. assumption, we would need to use long-run covariances that not only complicate the procedures but also rule out stationary regressors in the theory. See phillips2016robust for detailed discussions.
Remark 6.
In empirical finance, the m.d.s. assumption is commonly imposed on the error term, especially when testing asset return predictability (zhu2014predictive; kostakis2015robust). It indicates that the dependent variable (financial asset return) is not predictable if the null hypothesis of zero regression coefficients is not rejected, which aligns with the Efficient Market Hypothesis. In macroeconomic applications, the high dimensional covariates with different degrees of persistence alleviate the concern of variable omission, which makes the m.d.s. assumption plausible. [^2] Mild model misspecification with approximate sparsity can be accommodated by our framework; see belloni2012sparse and mei2022lasso.
Remark 7.
It is possible to extend our methodology and theoretical results to conditional heteroskedastic errors. Under low dimensional predictive regressions, kostakis2015robust show that the homoskedastic-only standard error of the IVX estimator is robust to conditional heteroskedastic error terms when the regressor of interest is persistent. We conjecture that this result applies to XDlasso, and the expression of our standard error (2.14) is robust to conditional heteroskedasticity in our two empirical applications, where each predictor of interest is persistent. Simulation results in Tables B.5 and B.6 provide supportive evidence on the conjecture. In Appendix B.3, we also consider a heteroskedasticity-robust standard error (B.5), and verify its validity by simulations. A complete theory of conditional heteroskedasticity in high dimensional predictive regression deserves a standalone paper for future research.
Second, MS24 assume the innovations follow linear processes, and impose the mixing condition and the lower bounded long run variances through the coefficients in the linear process. In contrast, our Assumption 2 does not assume any specific form of the linear process for $u_{t}$ and $\varepsilon_{t}$, but directly imposes the exponentially decaying rate for the mixing coefficient and the lower bound of the long run variances.
Third, Assumption 4(a) follows MS24 by allowing $p$ to diverge at a polynomial rate of $n$; it can be extended to an exponential rate of $n$ at the cost of expositional complications. Assumption 4(b) imposes a more restrictive condition for the sparsity index $s$, compared to $s=o(n^{1/4})$ in MS24. This is understandable as asymptotic normality is more delicate and demanding than consistency. This condition ensures that the shrinkage bias is accurately estimated to achieve the result (R2) in Section 2.2, so that XDlasso asymptotically follows a normal distribution centered at the true coefficient; also see the second term of (3.17) below.
Remark 8.
In practice, the sparsity index $s$ is unknown. We therefore recommend $\tau=1/2$ for practitioners, under which the quantity $(\tau\wedge(1-\tau))/4$ achieves its maximum $1/8$ and thus permits the weakest sparsity condition. This is different from the conventional wisdom of IVX (phillips2016robust; kostakis2015robust) where $\tau$ is recommended to be as large as 0.95 to minimize the loss of rate efficiency (or local power). Their context is different from our setting of high dimensional predictive regression, under which an excessively large $\tau$ will permit a very small sparsity index $s$ relative to $n$ and thus cause a severe distortion in the size of the test under finite sample. As discussed in Theorem 2 below, the hypothesis testing based on XDlasso is consistent for a wide class of local-to-zero $\theta_{j}^{*}$, despite a slower convergence rate compared with the case when $\tau$ is close to 1 (campbell2006efficient).
Remark 9.
Our theoretical results do not cover cointegrated regressors. Theorem 4 in MS24 shows that in the presence of cointegration, Slasso over-penalizes the coefficients of cointegrated regressors and shrinks them all the way to zero, regardless of their true values. We are unaware of any regularization method that achieves consistent estimation in $p>n$ regime with cointegrated variables mixed with LURs and stationary ones. Without consistency, inferential theory on the cointegrated variable’s true parameter is beyond reach at this moment. Our Appendix B.2 uses a numerical example to show that under some specifications XDlasso can remain robust despite the presence of cointegrated control variables.
Assumptions 14 will be sufficient for the consistency of Slasso with high dimensional LURs and stationary regressors. As the bias correction of XDlasso is mounted on the workhorse estimator Slasso, the consistency of the latter is a prerequisite for the ensuing maneuver.
3.1 Consistency of Slasso
The leading case of persistent regressors is LUR in the low-dimensional predictive regressions (campbell2006efficient). LUR includes the unit root as a special case, and is thus more general in modeling nonstationary behaviors. lee2022lasso study the variable selection properties by the adaptive LASSO under a finite number of LUR regressors, and MS24 cover the consistency of Slasso under high dimensional unit roots. In this paper, the first theoretical result extends Slasso’s consistency in the latter paper to incorporate the LUR processes in the former one. This generalization calls for sophisticated arguments, as briefed in Remark 10.
The consistency of Slasso is founded on two building blocks. The first one, which is essential and challenging, is the restricted eigenvalue (RE) condition of the Gram matrix of the standardized regressors. For any $L>1$, the RE of any $p\times p$ matrix $\Sigma$ is defined as
$\kappa_{H}(\Sigma,L,s):=\inf_{\delta\in\mathcal{R}(L,s)}\dfrac{\delta^{\top}H^{-1}\Sigma H^{-1}\delta}{\delta^{\top}\delta},$ (3.6)
where $\mathcal{R}(L,s)=\{\delta\in\mathbb{R}^{p}\backslash\{0_{p}\}:\|\delta_{\mathcal{M}^{c}}\|_{1}\leq L\|\delta_{\mathcal{M}}\|_{1},\ \text{for all }|\mathcal{M}|\leq s\}.$ The generic matrix $H$ is a placeholder and varies in different contexts. Let $\widehat{\Sigma}=W^{\top}W/n$ be the sample Gram matrix of all regressors. In the context of Slasso, we consider $\Sigma=\widehat{\Sigma}$ and $H=D$ along with the scale standardization in (2.4). The choice of the constant $L$ is related to the procedures of technical proofs and does not impact the rate of convergence. Following the common practice (bickel2009simultaneous), we set $L=3$ as a convenient choice, and simplify the notation as $\widehat{\kappa}_{D}=\kappa_{D}(\widehat{\Sigma},3,s)$. The quantity $\widehat{\kappa}_{D}$ will appear, according to Lemma 1 in MS24, in the denominator of Slasso’s convergence rates. Therefore, a lower bound for $\widehat{\kappa}_{D}$ is essential for the consistency of Slasso.
The second condition for Slasso’s consistency is the deviation bound (DB) of the cross-product between the error term $u_{t}$ in (2.1) and the standardized regressors. The theoretical order of the tuning parameter $\lambda$ must be no smaller than that of $\|n^{-1}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\|_{\infty}$ to avoid overfitting. On the other hand, an excessively large $\lambda$ causes over shrinkage and damages consistency. A tight upper bound of $\|n^{-1}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\|_{\infty}$ is therefore indispensable.
Next, we establish the RE and DB conditions for high dimensional mixed roots, and highlight their similarities to and differences from those in MS24. We then leverage them to derive the convergence rates of Slasso.
Lemma 1.
Under Assumptions 14, there exists an absolute constant $c_{\kappa}$ such that
$\widehat{\kappa}_{D}\geq\frac{c_{\kappa}}{s(\log p)^{4}}\cdot$ (3.7)
w.p.a.1 as $n\to\infty.$ In addition, there exists some absolute constant $C_{{\rm DB}}$ sch that
$4\,\bigg\|\dfrac{1}{n}\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\bigg\|_{\infty}\leq\dfrac{C_{{\rm DB}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}$ (3.8)
w.p.a.1 as $n\to\infty$, where $r$ is defined in Assumption 2.
Since LURs share similar asymptotic behavior with unit roots, the orders of the RE (Eq. (3.7)) and DB (Eq. (3.8)) are the same as those in Proposition 3 of MS24. Nevertheless, the technical proofs for LURs are challenging.
Remark 10.
For illustration, in this remark we suppose that all regressors are LURs. In low dimensions, the Gram matrix of LURs, after scaled by $n^{-1}$, converges in distribution to a non-degenerate stochastic integral
$n^{-1}\widehat{\Sigma}\stackrel{{\scriptstyle{\rm d}}}{{\to}}\int_{0}^{1}\mathcal{\mathcal{U}}_{{\rm{\bf C}}^{*}}(r)\mathcal{\mathcal{\mathcal{U}}}_{{\rm{\bf C}}^{*}}(r)^{\top}{\rm d}r,$ (3.9)
where $\mathcal{\mathcal{\mathcal{U}}}_{{\rm{\bf{\rm{\bf C}}^{*}}}}(t):=\int_{0}^{t}{\rm e}^{{\rm{\bf{\rm{\bf C}}^{*}}}(t-r)}{\rm d}\mathcal{B}(r)$ is a vector of Ornstein–Uhlenbeck processes, with ${\rm{\bf C}}^{*}:={\rm diag}(c_{1}^{*},c_{2}^{*},\cdots,c_{p_{x}}^{*})$ storing the constants in the AR(1) coefficients of LURs in (2.3), and $\mathcal{B}(r)$ being a multivariate Brownian motion. The diagonal entries of the stochastic integral on the right-hand side of (3.9) are nonnegative and continuously distributed, with a non-trivial probability in a neighborhood of zero. Consequently, the minimum diagonal entry of $n^{-1}\widehat{\Sigma}$ diminishes to zero as the dimension of LUR regressors passes to infinity. Eq. (3.7) establishes a lower bound of RE that shrinks to zero in a sufficiently slow speed, thereby still ensuring the consistency of Slasso. Under LUR, the linear coefficients in $x_{j,t}=\sum_{\ell\geq 0}(1+c_{j}^{*}/n)^{\ell}e_{j,t-\ell}$ depends on $n$ and $\ell$, which is much more complicated to deal with than the special case $c_{j}^{*}=0$ in MS24, where all the linear coefficients become $1$ and the representation of $x_{j,t}$ becomes a simple partial sum of stationary components. Interested readers may refer to Lemma A.4 in Section A.1.2 of the Appendix for details.
With RE and DB, we formulate the consistency of the initial workhorse Slasso estimator in the following lemma.
Lemma 2.
Under Assumptions 14, there exists some absolute constant $C_{{\rm m}}$ such that when the tuning parameter in (2.4) for the main regression satisfies $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$, as $n\to\infty$, we have
$\|D(\widehat{\theta}^{{\rm S}}-\theta^{*})\|_{1}$ $=O_{p}\left(\dfrac{s^{2}}{\sqrt{n}}(\log p)^{6+\frac{1}{2r}}\right).$ (3.10)
Lemma 2 provides the consistency of the standardized coefficients, which is necessary to establish the asymptotic normality of XDlasso. The rates of tuning parameters are technical devices for proofs. In practice, we recommend cross-validation to select the tuning parameter $\lambda$.
3.2 The Auxiliary Regression
While Lemma 2 generalizes the convergence of Slasso from unit roots to LUR regressors, brand new theory must be developed for the auxiliary regression (2.12), again by the workhorse Slasso when $w_{j,t}$ is either LUR or stationary. Importantly, as we maintain an agnostic attitude about the persistence of the variable of interest, which is the dependent variable of the auxiliary regression, the theory must be applicable in a unified manner to accommodate two very different types of time series. Without loss of generality, let $\mathcal{M}_{x}=[p_{x}]$ and $\mathcal{M}_{z}=[p]\backslash[p_{x}]$ denote the integer sets indexing the locations of LURs and stationary regressors, respectively.
We first examine the case when $j\in\mathcal{M}_{x}.$ The standardized instrument $\widetilde{\zeta}_{j,t}$ has a different degree of persistence from $X_{-j,t}$ and $Z_{t}$, due to the different orders of integration. In the low dimensional framework, the FCLT in phillips2016robust yields the following rate of convergence in OLS:
$D_{-j}\widehat{\varphi}^{(j){\rm OLS}}=D_{-j}\left(\sum_{t=1}^{n}W_{-j,t-1}W_{-j,t-1}^{\top}\right)^{-1}\sum_{t=1}^{n}W_{-j,t-1}\widetilde{\zeta}_{j,t}=O_{p}\left(1/\sqrt{n^{\tau\wedge(1-\tau)}}\right).$ (3.11)
In high dimensional models, the sample Gram matrix is rank-deficient and the FCLT no longer works. Thanks to the $L_{1}$ penalization, Slasso has a comparable local-to-zero order as (3.11), which is shown in Proposition 1 below.
We then turn to $j\in\mathcal{M}_{z}$, and slightly abuse the notation $Z_{-j,t}$ to denote the vector of stationary regressors excluding $w_{j,t}$. Define
$\varphi_{0z}^{(j)*}:=\mathbb{E}\left(Z_{-j,t}Z_{-j,t}^{\top}\right)^{-1}\mathbb{E}\left(Z_{-j,t}w_{j,t}\right)$ (3.12)
as the linear projection of $w_{j,t}$ on $Z_{-j,t}$, and normalize it by the standard deviation of the IV as
$\varphi_{z}^{(j)*}=\widehat{\varsigma}_{j}^{-1}\varphi_{0z}^{(j)*}.$ (3.13)
The “pseudo-true” model for the LASSO regression (2.12) is
$\widetilde{\zeta}_{j,t}=X_{t}^{\top}0_{p_{x}}+Z_{-j,t}^{\top}\varphi_{z}^{(j)*}+\widetilde{\eta}_{j,t},$ (3.14)
where $\text{$\widetilde{\eta}$}_{j,t}=(\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$. Note that when $w_{j,t}$ is stationary, the IV $\zeta_{j,t}$ is close to $w_{j,t}$ under a large sample size, and thus asymptotically uncorrelated with the LUR regressors $X_{t}$. As a result, the coefficients associated with $X_{t}$ in the pseudo-true model (3.14) are zero.
In addition, the error term $\widetilde{\eta}_{j,t}$ is close to the stationary time series $(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})/\widehat{\varsigma}_{j}$ and thus asymptotically uncorrelated to the nonstationary regressors $X_{t}$ due to different persistence. Furthermore, the coefficient $\varphi_{0z}^{(j)*}$ satisfies $\mathbb{E}\left(Z_{-j,t}(w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*})\right)=0$. Therefore, $\widetilde{\eta}_{j,t}$ is also asymptotically uncorrelated to the the stationary regressors $Z_{-j,t}$, thereby ensuring the consistency of the Slasso estimator $\widehat{\varphi}^{(j)}$ in (2.12). We impose the following assumption on the coefficient $\varphi_{0z}^{(j)*}$.
Assumption 5.
$\|\varphi_{0z}^{(j)*}\|_{0}\leq s$ with $\varphi_{0z}^{(j)*}$ defined in (3.12) and $s$ specified in Assumption 4. Moreover, $\|\varphi_{0z}^{(j)*}\|_{1}\leq C_{1}$ for some absolute constant $C_{1}.$
To bound the LASSO estimation errors, we need sparsity of not only the main regression (2.1), but also the auxiliary regression (3.14). This is similar to the high dimensional sparse instrumental variable regression; see zhu2018sparse and gold2020inference. In Assumption 5, we slightly abuse the sparsity index $s$ to bound the number of nonzero coefficients in the vector $\varphi_{0z}^{(j)*}$. For simplicity, we directly impose this high-level sparsity assumption on $\varphi_{0z}^{(j)*}$, which can be deduced under the commonly used sparsity restriction on the precision matrix $\left(\mathbb{E}\left(Z_{t}Z_{t}^{\top}\right)\right)^{-1}$; see the definition of $s_{j}$ and the conditions in Theorem 2.1 of zhang2017simultaneous. Finally, the upper bound of the $L_{1}$-norm of $\varphi_{0z}^{(j)*}$ controls the variance of the error term in the pseudo-true model (3.14). Parallel conditions naturally hold for $j\in\mathcal{M}_{x}$ as the pseudo-true coefficients are zero in view of the local-to-zero OLS estimate displayed in (3.11).
The following proposition formally lays out the convergence rate of the auxiliary estimator (2.12).
Proposition 1.
Suppose that Assumptions 15 hold. Then, there exists some absolute constant $C_{{\rm a},j}>0$ such that when $\mu_{j}=C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, we have
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right),$ (3.15)
where $\varphi^{(j)*}={\bf 1}\{j\in\mathcal{M}_{x}\}\cdot 0_{p-1}+{\bf 1}\{j\in\mathcal{M}_{z}\}\cdot(0_{p_{x}}^{\top},\varphi_{z}^{(j)*\top})^{\top}$.
Remark 11.
The rate of tuning parameter $\mu$ specified in Proposition 1 induces the following Karush–Kuhn–Tucker (KKT) condition
$\|n^{-1}\sum_{t=1}^{n}(D_{-j})^{-1}W_{-j,t}\widehat{r}_{j,t}\|_{\infty}\leq\dfrac{\mu_{j}}{2}\leq\dfrac{C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}}{2\sqrt{n^{(1-\tau)\wedge\tau}}},$ (3.16)
which is sufficient to bound the approximation error $(\widehat{{\rm B}}_{j}-{\rm B}_{j})$ in (2.7). Nevertheless, it is still necessary to establish the consistency of $\widehat{\varphi}^{(j)}$. This consistency result not only helps us show the asymptotic normality of the $t$-statistic $t_{j}^{{\rm XD}}$ defined in (2.15) to guarantee the asymptotic size of the test, but is also critical for the convergence rate of the standard error that governs the asymptotic power.
The consistency in the main equation and the auxiliary regression has paved the way for statistical inference. We will analyze the asymptotic size and power of the concerning test statistics in the next section.
3.3 Asymptotic Distributions
The desirable rate of the auxiliary regression built in Proposition 1 guarantees the following decomposition of the $t$-statistic:
$\frac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}={\rm sgn}_{j}\cdot\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}+O_{p}\left(\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right),$ (3.17)
where ${\rm sgn}_{j}=\frac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}$ is either 1 or 1-1 with probability one as $\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}$ is continuously distributed. The first term of the above expression is a counterpart of $\mathrm{N}_{j}/\omega_{j}$ in the discussion of generic desparsifying argument in (2.7). The second term is the convergence rate of the approximation error of the shrinkage bias analogous to $(\widehat{{\rm B}}_{j}-{\rm B}_{j})/\omega_{j}$. Under Assumption 4, the second term on the right-hand side of (3.17) is asymptotically negligible, thereby yielding the following asymptotic normality for any $j\in[p]$.
Theorem 1.
Suppose Assumptions 1-5 hold. There exist absolute constants $C_{{\rm m}}$ and $C_{{\rm a},j}$ such that when $\lambda=C_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\mu_{j}=C_{{\rm a},j}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$, as $n\to\infty$ we have
$(\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*})\big/\widehat{\omega}_{j}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (3.18)
The asymptotic normality in Theorem 1 is our main theoretical result that delivers the valid asymptotic size of the hypothesis testing for $\mathbb{H}_{0}:\theta_{j}^{*}=\theta_{0,j}$ using the $t$-statistic $t_{j}^{{\rm XD}}$ in (2.15). The following Theorem 2 provides the convergence rate of the estimated standard error, which characterizes the asymptotic power of the test.
Theorem 2.
Conditions in Theorem 1 yield
$\widehat{\omega}_{j}^{{\rm XD}}=\boldsymbol{1}\left\{j\in\mathcal{M}_{x}\right\}\cdot O_{p}(1/\sqrt{n^{1+\tau}})+\boldsymbol{1}\left\{j\in\mathcal{M}_{z}\right\}\cdot O_{p}\left(1/\sqrt{n}\right).$
The convergence rates displayed in Theorem 2 for the two types of regressors are coherent with the results in low dimensional predictive regressions. When $j\in\mathcal{M}_{x}$, the standard error converges faster than the rate $1/n^{\delta_{j}}$ for any $\delta_{j}\in(0,(1+\tau)/2)$. Thus, for the null hypothesis $\mathbb{H}_{0}:\beta_{j}^{*}=0$, the hypothesis testing based on the $t$-statistic $t_{j}^{{\rm XD}}$ is consistent under the alternative with $\beta_{j}^{*}=c/n^{\delta_{j}}$ for some $c>0$ over this class of $\delta_{j}$. The range $\delta_{j}\in(0,(1+\tau)/2)$ includes the important $1/\sqrt{n}$ rate of coefficients for the LUR regressor $X_{j,t}$, under which ${\rm var}(X_{j,t}\beta_{j}^{*})=O(1).$ The $1/\sqrt{n}$ factor thus balances the larger order of LUR regressors $X_{t}$ and the standard $O_{p}(1)$ stochastic order of $y_{t}$ (phillips2015halbert; lee2022lasso), and the test by XDlasso is consistent under this class of alternatives. When $j\in\mathcal{M}_{z}$, XDlasso achieves the standard $\sqrt{n}$-consistency.
Remark 12.
Unlike the convergence rates of Slasso in Lemma 2, the convergence rate of XDlasso in Theorem 2 is independent of the variable dimension $p$ and the sparsity index $s$. In view of (2.7), in XDlasso the order of the approximation error of the shrinkage bias $(\widehat{{\rm B}}_{j}-{\rm B}_{j})$ depends on $p$ and $s$, but it is dominated by the order of the noise component ${\rm N}_{j}$. The convergence rate of XDlasso therefore follows the order of ${\rm N}_{j}$, which relates only to the sample size $n$.
Finally, when we are interested in a joint null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$ for the coefficients indexed by $\mathcal{J}\subset[p]$ with a fixed cardinality, we can use the Wald statistic in (2.19) for this test.
Theorem 3.
Suppose that $|\mathcal{J}|$ is fixed as $n\to\infty$, and the conditions in Theorem 1 hold for all $j\in\mathcal{J}$. Under the null hypothesis $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}}$, we have
${\rm Wald}_{\mathcal{J}}^{{\rm XD}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\chi_{|\mathcal{J}|}^{2}.$
Theorem 3 is a natural extension from the $t$-statistic for a one-dimensional univariate hypothesis to a simultaneous multivariate hypothesis. Notice that the Wald statistic is valid even if $\mathcal{J}$ includes a mixture of LUR regressors and stationary ones. It allows the researcher to maintain the agnostic attitude when conducting the joint hypothesis testing.
Let us summarize the insights gained from the theoretical development. The fundamental principle behind the Dlasso method, as discussed in zhang2014confidence, is based on the Frisch-Waugh-Lovell theorem. This theorem purges the influence of other control variables to overcome shrinkage bias. In the context of persistent predictors, the two-stage least squares approach, as developed in magdalinos2009limit is used to eliminate Stambaugh bias. Each piece of XDlasso is a machine learning version of a classical idea. To adapt these existing procedures to high dimensional predictive regressions, we must rely on the consistency of Slasso, with its extension to LURs. This consistency is crucial for both the main regression and the auxiliary regression. Moreover, as mentioned in Remark 8, the construction of the IV and the choice of $\tau$ must be tailored to balance the two types of predictors. This approach is original in that such a restriction is unique to high dimensional models and has not been studied before even in conventional low dimensional settings.
4 Monte Carlo Simulations
In this section, we evaluate the performance of the proposed XDlasso inference procedure by comparing its test size and power with those of Dlasso. Although Dlasso and other existing LASSO bias-correction procedures are not designed to handle persistent regressors, we include this comparison to highlight the value added by the IVX transformation. Furthermore, to demonstrate the robustness of the XDlasso approach, we benchmark it using tests based on infeasible estimators, where an oracle reveals the locations of the nonzero coefficients.
4.1 Setup
We consider the linear predictive regression model in (2.1). The vector of the stationary components, denoted as $v_{t}=\left(u_{t},e_{t}^{\top},Z_{t}^{\top}\right)^{\top}$, is generated by the two different processes:
Case I (IID Innovations): $v_{t}\sim\text{i.i.d. }\mathcal{N}\left(0,\Sigma\right),$ (4.1)
Case II (AR(1) Innovations): $v_{t}=R_{n}v_{t-1}+\xi_{t},\ \ \xi_{t}\sim\text{i.i.d. }\mathcal{N}\left(0,\Sigma\right),$ (4.2)
where $R_{n}=\text{diag}(0,0.3,0.3,\ldots,0.3)$. Under this $R_{n}$ the error term $u_{t}$ in the main regression (2.1) remains i.i.d., satisfying the m.d.s. condition in Assumption 1, while the (local) unit root innovations $e_{t}$ and the stationary regressors $Z_{t}$ are AR(1) processes. The covariance matrix $\Sigma=(\Sigma_{ij})_{i,j\in[p]}$ is specified as
$\Sigma_{ij}=\begin{cases}0,&\text{if }(i,j)\text{ is associated with }Z_{t}\text{ and }u_{t};\\ 0.5^{|i-j|},&\text{otherwise}.\end{cases}$
The persistent regressors $X_{t}$ are generated by
$X_{t}=\text{diag}(\rho^{*})X_{t-1}+e_{t},$ (4.3)
where $\rho^{*}=(1,1-1/n,1+1/n,1,1-1/n,1+1/n,\cdots)^{\top}\in\mathbb{R}^{p_{x}}$. Recall $0_{p}$ is a $p$-dimensional zero vector and $1_{p}$ is a $p$-dimensional vector of ones. The true coefficient vectors are:
$\beta^{*}=\left(\beta_{1}^{*},\dfrac{0.5}{\sqrt{n}}\times 1_{4}^{\top},0_{p_{x}-5}^{\top}\right)^{\top},\ \ \gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},0_{p_{z}-5}^{\top})^{\top}.$ (4.4)
The specification involves four active LUR regressors and four active stationary regressors. [^3] In Section B.1, we consider a setting with more nonzero coefficients and find robust performance of XDlasso. The $1/\sqrt{n}$ scaling balances the regression by normalizing the coefficients of the unit root regressors. We test the hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=0$, $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, and the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$, respectively, under the sample sizes $n\in\{200,300,400,500,600\}$ and the dimensionality pairs $(p_{x},p_{z})\in\{(50,100),(100,150),(150,300)\}.$ We conduct 2000 replications in each setting.
We compare the finite sample performance of XDlasso, as described in Algorithm 1, and Dlasso with the score vector in (2.16) and (2.17). In addition, we consider two infeasible testing procedures as benchmarks. Using the known active set of regressors, we conduct IVX inference (IVX oracle) and the standard $t$-test based on the OLS estimator (OLS oracle), employing only the regressor of interest and the active regressors, which form a low-dimensional predictive regression model. We set $C_{\zeta}=5$ and $\tau=0.5$ for the parameter $\rho_{\zeta}$ specified as (3.5). As discussed in Remark 8, the choice $\tau=$ 0.5 admits the weakest sparsity condition, and thus effectively improves the finite sample performance.
Both XDlasso and Dlasso involve Slasso, where the selection of tuning parameters $\lambda$ and $\mu$ affects finite sample performance. In our experiments, we employ the block 10-fold cross-validation (CV), splitting the sample into 10 equally sized chronologically ordered consecutive blocks for validations. Though the unconditional variances of nonstationary regressors vary in different chronological blocks, the standardization in the Slasso estimators like (2.4) and (2.12) account for such variation. This explains the robustness of the block CV to nonstationary time series, as shown by our simulation results.
In Theorem 1, the tuning parameters are specified as constants multiplied by the appropriate rates of convergence determined by the sample size $n$, dimensionality $p$ and the mixing condition constant $r$. As a benchmark, we also calibrate the tuning parameters following lee2022lasso to examine the validity of the theoretical orders of tuning parameters specified in Theorem 1. Specifically, we perform $500$ pilot replications for each DGP, with $n_{0}=400$ , $\left(p_{x0},p_{z0}\right)=\left(100,150\right)$, and $p_{0}=p_{x0}+p_{z0}$. In each replication $q=1,2,\dots,500$, we use the 10-fold cross-validation to choose the tuning parameters $\lambda^{(q)}$ and $\mu^{(q)}$, and calibrate the constants as
$C_{{\rm m}}^{\left(q\right)}=\lambda^{(q)}n_{0}^{1/2}/(\log p_{0})^{\frac{3}{2}+\frac{1}{2r}},\ \ C_{{\rm a}}^{\left(q\right)}=\mu^{(q)}n_{0}^{[(1-\tau)\wedge\tau]/2}/(\log p_{0})^{2+\frac{1}{2r}},$
where $r=1$ and $\tau=0.5$ are chosen in the simulation. We then fix $\widehat{C}_{\star}=\mathrm{median}(C_{\star}^{\left(1\right)},\ldots,C_{\star}^{\left(500\right)})$ for $\star\in\left\{{\rm m},{\rm a}\right\}$ in the full-scale experiments. The tuning parameters are then set as $\widehat{\lambda}=\widehat{C}_{{\rm m}}(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}$ and $\widehat{\mu}=\widehat{C}_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ as in Theorem 1.
4.2 Results
We first investigate the empirical size of different testing methods at a 5% nominal significance level when the true coefficients $\beta_{1}^{*}=0$ and $\gamma_{1}^{*}=0$. Tables 1 and 2 report the empirical sizes under the IID and AR(1) innovations, respectively. Foremost among the findings is that XDlasso effectively controls the empirical size for both $\beta_{1}^{*}$, associated with a unit root regressor, and $\gamma_{1}^{*}$, associated with a stationary regressor. This performance stands in sharp contrast to that of Dlasso and OLS oracle, which exhibit severe size distortions for $\beta_{1}^{*}$. Such distortions can be attributed to the failure to account for the Stambaugh bias arising from nonstationarity. Furthermore, the results yield noteworthy insights regarding the tuning parameter selection. The empirical size of XDlasso with both cross-validated and calibrated tuning parameters is close to the nominal level. This result not only validates the asymptotic rates of tuning parameters specified in Theorem 1 but also supports CV as a feasible data-driven tuning parameter selection method in practice. In addition, we investigate the efficiency and robustness of XDlasso in comparison with alternatives. When compared to the unbiased but infeasible “IVX oracle” estimator, the confidence intervals produced by XDlasso are only slightly wider. In contrast to the estimators solely for low-dimensional data, XDlasso demonstrates robustness by accommodating high dimensional covariates without compromising much in efficiency. Lastly, the empirical sizes for testing the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$, as reported in Table 3, are also well controlled around the nominal level across setups. The results validate the theoretical result in Theorem 3 and is consistent with the findings for testing $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ individually.
Table 1: Empirical size and length of confidence interval: IID innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.035 0.216 0.150 0.098 0.050 0.222 0.354 0.104 0.060 0.227 0.438 0.159
300 0.044 0.153 0.145 0.066 0.052 0.163 0.436 0.079 0.060 0.168 0.527 0.122
400 0.055 0.122 0.148 0.050 0.049 0.132 0.474 0.064 0.061 0.136 0.572 0.096
500 0.049 0.102 0.149 0.040 0.056 0.112 0.512 0.054 0.070 0.114 0.591 0.076
600 0.046 0.087 0.141 0.033 0.049 0.098 0.553 0.047 0.061 0.100 0.602 0.065
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.041 0.217 0.158 0.099 0.052 0.220 0.382 0.102 0.060 0.225 0.484 0.160
300 0.047 0.154 0.158 0.066 0.059 0.163 0.487 0.077 0.077 0.169 0.601 0.126
400 0.048 0.123 0.156 0.050 0.051 0.130 0.540 0.062 0.067 0.135 0.681 0.106
500 0.047 0.103 0.159 0.040 0.045 0.109 0.599 0.053 0.065 0.115 0.714 0.088
600 0.042 0.088 0.156 0.033 0.045 0.096 0.632 0.046 0.065 0.099 0.745 0.077
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.047 0.219 0.157 0.100 0.052 0.213 0.349 0.097 0.064 0.222 0.508 0.140
300 0.046 0.154 0.138 0.065 0.046 0.157 0.421 0.073 0.055 0.166 0.585 0.115
400 0.046 0.122 0.134 0.049 0.043 0.126 0.500 0.059 0.051 0.134 0.665 0.097
500 0.045 0.101 0.136 0.039 0.057 0.105 0.547 0.050 0.068 0.114 0.694 0.084
600 0.045 0.087 0.144 0.033 0.052 0.093 0.593 0.044 0.066 0.099 0.742 0.075
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.043 0.373 0.064 0.324 0.068 0.324 0.065 0.288 0.071 0.323 0.066 0.288
300 0.042 0.294 0.050 0.264 0.062 0.265 0.065 0.240 0.061 0.265 0.065 0.240
400 0.039 0.251 0.045 0.229 0.053 0.229 0.060 0.210 0.056 0.229 0.060 0.211
500 0.041 0.222 0.045 0.203 0.050 0.204 0.052 0.189 0.047 0.204 0.054 0.190
600 0.043 0.200 0.045 0.185 0.049 0.186 0.049 0.174 0.049 0.187 0.050 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.041 0.371 0.055 0.323 0.075 0.324 0.074 0.287 0.080 0.320 0.080 0.284
300 0.049 0.294 0.057 0.263 0.070 0.264 0.066 0.239 0.072 0.263 0.065 0.239
400 0.058 0.251 0.057 0.228 0.074 0.228 0.073 0.209 0.074 0.228 0.075 0.209
500 0.051 0.222 0.063 0.203 0.071 0.203 0.074 0.189 0.072 0.204 0.071 0.189
600 0.051 0.201 0.057 0.186 0.062 0.186 0.067 0.173 0.061 0.186 0.067 0.174
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.040 0.375 0.056 0.324 0.060 0.329 0.056 0.291 0.066 0.317 0.065 0.282
300 0.037 0.297 0.040 0.264 0.054 0.265 0.055 0.241 0.059 0.260 0.056 0.237
400 0.036 0.251 0.046 0.228 0.053 0.228 0.047 0.210 0.053 0.227 0.050 0.208
500 0.030 0.222 0.048 0.204 0.051 0.203 0.052 0.188 0.050 0.203 0.049 0.188
600 0.043 0.201 0.049 0.186 0.057 0.185 0.057 0.172 0.052 0.185 0.054 0.173
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ ($\approx 1.96$) is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table 2: Empirical size and length of confidence: AR(1) innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.052 0.162 0.162 0.072 0.064 0.167 0.391 0.077 0.089 0.172 0.489 0.134
300 0.045 0.113 0.154 0.048 0.058 0.122 0.458 0.058 0.091 0.125 0.551 0.096
400 0.052 0.088 0.160 0.036 0.056 0.098 0.507 0.047 0.081 0.100 0.582 0.070
500 0.050 0.073 0.145 0.029 0.056 0.081 0.547 0.040 0.082 0.084 0.617 0.056
600 0.049 0.063 0.147 0.024 0.053 0.071 0.574 0.035 0.075 0.073 0.625 0.047
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.052 0.161 0.149 0.073 0.059 0.164 0.410 0.075 0.095 0.172 0.541 0.130
300 0.053 0.113 0.149 0.048 0.059 0.120 0.511 0.057 0.098 0.126 0.646 0.105
400 0.047 0.088 0.158 0.036 0.046 0.093 0.564 0.046 0.082 0.099 0.711 0.078
500 0.046 0.074 0.154 0.028 0.055 0.080 0.620 0.039 0.076 0.084 0.735 0.063
600 0.046 0.063 0.157 0.023 0.056 0.068 0.641 0.034 0.082 0.072 0.762 0.056
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.051 0.161 0.157 0.072 0.050 0.159 0.368 0.071 0.081 0.170 0.536 0.113
300 0.046 0.112 0.146 0.047 0.049 0.113 0.461 0.054 0.075 0.122 0.624 0.086
400 0.052 0.088 0.142 0.035 0.057 0.091 0.532 0.043 0.076 0.099 0.662 0.072
500 0.051 0.073 0.140 0.028 0.058 0.077 0.587 0.037 0.076 0.084 0.718 0.062
600 0.039 0.062 0.150 0.023 0.056 0.065 0.619 0.032 0.076 0.072 0.756 0.057
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.044 0.380 0.057 0.312 0.067 0.333 0.065 0.275 0.067 0.331 0.068 0.273
300 0.051 0.298 0.055 0.254 0.065 0.269 0.061 0.229 0.068 0.269 0.066 0.228
400 0.048 0.252 0.053 0.219 0.059 0.231 0.060 0.200 0.061 0.231 0.064 0.200
500 0.043 0.222 0.044 0.195 0.052 0.205 0.049 0.180 0.053 0.205 0.049 0.180
600 0.041 0.200 0.046 0.178 0.055 0.186 0.050 0.165 0.056 0.187 0.048 0.165
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.049 0.381 0.054 0.313 0.080 0.334 0.078 0.274 0.084 0.330 0.080 0.272
300 0.050 0.299 0.056 0.254 0.071 0.268 0.074 0.228 0.074 0.268 0.083 0.227
400 0.053 0.252 0.055 0.219 0.073 0.230 0.070 0.200 0.075 0.230 0.074 0.199
500 0.050 0.221 0.058 0.195 0.069 0.204 0.067 0.179 0.069 0.205 0.066 0.179
600 0.053 0.200 0.056 0.178 0.063 0.185 0.065 0.165 0.060 0.186 0.064 0.165
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.037 0.382 0.052 0.314 0.066 0.334 0.068 0.276 0.068 0.326 0.067 0.268
300 0.036 0.300 0.050 0.254 0.058 0.268 0.061 0.228 0.061 0.265 0.060 0.225
400 0.042 0.252 0.050 0.218 0.064 0.229 0.062 0.199 0.065 0.228 0.065 0.197
500 0.041 0.222 0.045 0.196 0.048 0.203 0.060 0.179 0.048 0.203 0.058 0.178
600 0.047 0.200 0.053 0.178 0.058 0.184 0.061 0.164 0.056 0.185 0.061 0.164
Notes: The data generating process corresponds to (4.2). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}$ ($\approx$ 1.96) is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table 3: Empirical size: Joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$
Case I: IID Innovations
$\left(p_{x},p_{z}\right)=\left(50,100\right)$ $\left(p_{x},p_{z}\right)=\left(100,150\right)$ $\left(p_{x},p_{z}\right)=\left(150,300\right)$
$n$ Calibrated CV Calibrated CV Calibrated CV
200 0.064 0.077 0.067 0.076 0.053 0.069
300 0.057 0.066 0.064 0.081 0.052 0.063
400 0.058 0.069 0.067 0.080 0.047 0.056
500 0.054 0.064 0.061 0.069 0.053 0.063
600 0.051 0.055 0.060 0.069 0.056 0.059
Case II: AR(1) Innovations
$\left(p_{x},p_{z}\right)=\left(50,100\right)$ $\left(p_{x},p_{z}\right)=\left(100,150\right)$ $\left(p_{x},p_{z}\right)=\left(150,300\right)$
$n$ Calibrated CV Calibrated CV Calibrated CV
200 0.070 0.090 0.064 0.098 0.061 0.084
300 0.070 0.095 0.065 0.096 0.055 0.079
400 0.056 0.075 0.063 0.090 0.054 0.070
500 0.059 0.079 0.064 0.086 0.055 0.067
600 0.053 0.071 0.067 0.079 0.060 0.076
Notes: The data generating processes for Case I and Case II correspond to (4.1) and (4.2), respectively. The table reports the empirical size of testing the joint null hypothesis $\mathbb{H}_{0}:\beta_{1}^{*}=\gamma_{1}^{*}=0$ at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[\text{Wald}_{\mathcal{J}}^{\text{XD}(r)}>\chi_{0.95,2}^{2}\right]$ across $R=2,000$ replications, where $\text{Wald}_{\mathcal{J}}^{\text{XD}(r)}$ is computed based on (2.19) for the $r$-th replication with $\mathcal{J}=\left\{1,p_{x}+1\right\}$, and the critical value $\chi_{0.95,2}^{2}(\approx 5.99)$ is the 95-th percentile of the chi-squared distribution with 2 degrees of freedom. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
We now turn to the empirical power of the XDlasso inference. Figure 2 plots power curves for the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ under varying true coefficient values. In our analysis, we vary either $\beta_{1}^{*}$ or $\gamma_{1}^{*}$ from 0 to $0.5$. All remaining coefficients are held fixed as specified in (4.4). Across various configurations, XDlasso exhibits increasingly high power against the null hypothesis as the true coefficient moves away from 0 and the sample size $n$ increases. Furthermore, Figure 3 compares the power for testing $\beta_{1}^{*}$ and $\gamma_{1}^{*}$, which reveals that the power associated with a unit root regressor surpasses that of a stationary regressor. This observation provides finite sample evidence to support the theoretical results in Theorem 2. Specifically, it corroborates the faster convergence of standard error for unit root regressors compared with stationary regressors, thereby inducing higher power in the hypothesis testing.
Figure 2: Power curves of XDlasso inference
(a) IID Innovations
(b) AR(1) Innovations
Notes: The left and right panels correspond to DGPs (4.1) and (4.2), respectively. In each subplot, the first row depicts the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, across various $\left(p_{x},p_{z}\right)$ configurations, while the second row pertains to $\gamma_{1}^{*}$, associated with a stationary regressor. The empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution.
Figure 3: Power of XDlasso inference for nonstationary and stationary regressors
Notes: This figure plots the power curves under $n=600.$ The first and second rows correspond to DGPs (4.2) and (4.1), respectively. In each subplot, blue lines represent the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, while red lines represent that for $\gamma_{1}^{*}$, associated with a stationary regressor. Empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution.
5 Empirical Applications
In recent years, high dimensional macroeconomic data have been extensively used to forecast key variables of interest; see smeekes2018macroeconomic, medeiros2021forecasting, and giannone2021economic, for example. Researchers have primarily focused on the point estimation of forecast. There has been limited empirical exploration of statistical inference on the predictive power of specific predictors, due to a lack of suitable toolkits. This section showcases two empirical applications using our proposed XDlasso inference method. We utilize the monthly data of 112 U.S. macroeconomic variables spanning from January 1960 to April 2025, sourced from the FRED-MD dataset by mccracken2016fred.
5.1 Predictability of Stock Return Using Earnings-Price Ratio
In financial economics, testing the predictive power of valuation ratios, particularly the earnings-price ratio, on stock returns has been subject to widespread discussions. Much of the literature tests the predictability using univariate predictive regressions (e.g., welch2008comprehensive, zhu2014predictive, and goyal2024comprehensive), but inference can be sensitive to model misspecification arising from omitted variables. Controlling for high dimensional covariates is therefore necessary not only to enhance out-of-sample prediction but also to mitigate the omitted variables for credible and accurate inference. The literature on predictive regression for stock returns has focused on identifying which variables possess significant predictive power for future returns. For better out-of-sample prediction, recent literature has documented gains in forecasting performance from incorporating high-dimensional covariates into predictive regressions; see, for instance, smeekes2018macroeconomic, gu2020empirical and medeiros2021forecasting. The increasing popularity of statistical learning methods with high dimensional predictors calls for suitable inference methods beyond univariate predictive regressions. The goal is to provide investors with signals to adjust their portfolios based on changes in these predictive variables. In this section, we investigate the predictability of stock returns using the log earnings-price ratio in a data-rich environment with high dimensional mixed-root control variables.
5.1.1 Data
Our analysis focuses on predicting the monthly return of the S&P 500 index, calculated as $\text{Return}_{t}=\log(P_{t})-\log(P_{t-1}),$ where $P_{t}$ refers to S&P 500 (S&P’s Common Stock Price Index: Composite). The primary predictor of interest is the log earnings-price ratio. We obtain it from the variable S&P PE ratio (S&P’s Composite Common Stock: Price-to-Earnings Ratio), denoted as $\text{PE}_{t}$. The log earnings-price ratio is calculated by inverting the original price-earnings ratio as $\text{logEP}_{t}=\log(1/\text{PE}_{t}).$
Figure 4 displays the time series plot of the monthly return of the S&P 500 index and the log earnings-price ratio from January 1960 to April 2025. The log earnings-price ratio exhibits persistent patterns, while the S&P 500 return appears stationary. We further report the AR(1) coefficient estimates and Augmented Dickey–Fuller (ADF) test $p$-values of both series under different sample periods in Table 4. The S&P 500 return is evidenced to be stationary under the full sample with an ADF test $p$-value below 1%, rejecting the null hypothesis of nonstationarity. Conversely, the log earnings-price ratio shows high persistence with an AR(1) coefficient estimate equal to $0.993$. Given the $p$-value of $0.074$, nonstationarity is not rejected at the 5% significance level. Note that the nonstationary log earnings-price ratio can predict the stationary monthly return of S&P 500 in our model, since we allow for a local-to-zero coefficient to balance the different scales between a stationary outcome and a nonstationary regressor (phillips2015halbert). Theorem 2 and the paragraph that follows illuminate that inference by XDlasso has the power to detect a wide range of local-to-zero alternatives.
Figure 4: S&P 500 monthly Return and Log Earnings-price Ratio
Table 4: Persistence of S&P 500 Monthly Return and Log Earnings-Price Ratio
Sample Period S&P 500 Monthly Return Log Earnings-Price Ratio
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2025) 0.227 <<0.01 0.993 0.074
Pre-1994 (Jan. 1960 - Dec. 1993) 0.255 <<0.01 0.994 0.776
Post-1994 (Jan. 1994 - Apr. 2025) 0.200 <<0.01 0.979 <<0.01
Notes: The lag order for the ADF test is set to $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.
Return predictability can be time-varying and sporadic (tu2023penetrating). campbell2006efficient discover that the return predictability test results can vary depending on the inclusion of the samples after 1994, and the break is revisited by zhu2014predictive. Following this empirical finding, we divide the sample into two periods: pre-1994 (January 1960 - December 1993) and post-1994 (January 1994 - April 2025). The S&P 500 return demonstrates stationarity across both subperiods with ADF $p$-values below 1%. In contrast, the persistence of the log earnings-price ratio depends on the sample period considered. Nonstationarity is evident in the Pre-1994 period with a large ADF test $p$-value $0.776$, while the null hypothesis of nonstationarity is rejected at the 1% significance level in the Post-1994 period. The ambiguity in stationarity of the log earnings-price ratio motivates the use of the XDlasso procedure, a unified approach for both stationary and nonstationary regressors without prior knowledge of their persistence.
Besides the monthly return of the S&P 500 index and the log earnings-price ratio, our analysis incorporates high dimensional covariates including all other 110 macroeconomic variables from the FRED-MD dataset. These variables comprise a mixture of stationary and nonstationary time series. In practice, it is common for empirical analysts to transform potentially nonstationary time series into stationary ones to avoid challenges arising from nonstationarity. To facilitate such stationarization, FRED assigns a transformation code (TCODE) to each variable denoting the recommended transformation.
In our analysis, we follow the common practice of using the TCODE to transform the 110 time series, and subsequently use the transformed variables as covariates. Nevertheless, we highlight that these elementary transformations are not a silver bullet in taming nonstationarity. We perform the ADF test for each of the transformed variables for different sample periods. A nontrivial proportion (9.1%) of the transformed variables still demonstrate nonstationarity based on the ADF test at the 5% significance level for both Pre-1994 and Post-1994 subperiods. The high persistence of the log earnings-price ratio and other covariates suggests the necessity of XDlasso.
As highlighted by smeekes2020unit, the predictive performance in regressions using FRED-MD data is sensitive to the transformations. To assess the robustness of our results, we also conduct our analysis using the original (untransformed) FRED-MD time series as covariates. To reduce the impact of highly nonstationary series, we exclude I(2) variables that require second differencing for stationarity according to their TCODE classification. [^4] Appendix C.1 further investigates robustness by considering: (i) excluding nonstationary variables based on integrated orders determined by the bootstrap sequential testing procedure of smeekes2015bootstrap, as reported in smeekes2020unit, and (ii) applying only logarithmic transformations, as indicated by TCODE, without differencing.
5.1.2 Results
We study the one-month-ahead regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes a high dimensional vector that collects all control variables. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the log earnings-price ratio in forecasting the S&P 500 monthly return. Given the stationary pattern of the dependent variable, we additionally consider specifications where $W_{-1,t-1}$ includes the lagged dependent variable $\text{Return}_{t-1}$. For each model, we apply the wild bootstrapped automatic variance ratio test (kim2009automatic) to the Slasso residuals as a diagnostic test for the martingale difference sequence (m.d.s.) condition in Assumption 1.555This practice serves as a heuristic diagnostic. Demonstrated by our simulation results in Appendix B.4, the variance ratio test on the Slasso residuals $\widehat{u}_{t}$ tends to severely over-reject the m.d.s. condition for $u_{t}$. To the best of our knowledge, at present the literature has no valid testing procedure yet for m.d.s. in high dimensional predictive regressions. It is an open question for future research.
In addition to XDlasso and Dlasso, we conduct the IVX inference in a simple regression setup, omitting all control variables, to demonstrate the necessity of incorporating high dimensional control variables in practice. We compare the results of XDlasso to those of Dlasso for the high dimensional regression and IVX for a simple regression $\text{Return}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{logEP}_{t-1}+u_{t}.$
Table 5: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ across sample periods and specifications in stock return prediction
(a) TCODE Transformed Data
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period IVX Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.023** 0.009 0.003 0.074 0.009 0.005 0.216
(Jan. 1960 - Apr. 2025) (0.010) (0.006) (0.015) (0.006) (0.015)
Pre-1994 -0.059** 0.025** 0.059* 0.227 0.024** 0.062* 0.296
(Jan. 1960 - Dec. 1993) (0.026) (0.010) (0.033) (0.010) (0.032)
Post-1994 -0.022 0.002 -0.001 0.053 0.002 -0.001 0.049
(Jan. 1994 - Apr. 2025) (0.015) (0.007) (0.017) (0.007) (0.017)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period IVX Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.023** 0.013 -0.008 0.001 0.019 0.012 0.811
(Jan. 1960 - Apr. 2025) (0.010) (0.014) (0.011) (0.014) (0.010)
Pre-1994 -0.059** 0.064** -0.312 0.046 0.055* 0.096 0.467
(Jan. 1960 - Dec. 1993) (0.026) (0.031) (0.290) (0.033) (0.070)
Post-1994 -0.022 -0.003 -0.022 0.016 -0.000 -0.004 0.280
(Jan. 1994 - Apr. 2025) (0.015) (0.009) (0.017) (0.008) (0.016)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The upper and lower panels corresponds to the scenarios where we use TCODE transformed or untransformed time series as covariates, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table 5 presents the point estimates and testing results for $\theta_{1}^{*}$. IVX delivers negative point estimates across all sample periods and detects significant effects in the full sample period and Pre-1994 subperiod at the 5% significance level. However, the resulting negative relation between stock return and earning performance contradicts the economic mechanism between the two variables. A large number of potential confounding variables are present in the dataset, which may be the culprit in producing the counterintuitive result from the simple regression.
Dlasso, which accounts for high dimensionality but not nonstationarity, generally reverses the sign of the estimates compared to the simple regression. For the full sample, Dlasso consistently yields positive but statistically insignificant coefficients. Zooming into the Pre-1994 period, Dlasso detects strong evidence of predictability with transformed data, with coefficients significant at the 5% level. This result persists with the untransformed data, though the significance weakens to the 10% level when the lagged dependent variable is included. Accounting for the potential nonstationarity in the predictors, XDlasso provides notably different estimates and standard errors from Dlasso. XDlasso reveals no significant predictive power of log earnings-price ratio for stock return in the full sample and the Post-1994 period. Moreover, for the Pre-1994 period, the evidence of predictability is weakened to the 10% significance level with transformed data, and becomes entirely insignificant with untransformed data. Our empirical findings with XDlasso are consistent with the general recognition that there is little predictability in the financial market, thereby supporting the efficient market hypothesis. The divergence between Dlasso and XDlasso underscores our theoretical prediction: XDlasso mitigates the Stambaugh bias arising from highly persistent regressors while Dlasso does not. Despite the short confidence intervals, Dlasso results can be misleading when nonstationary time series is present, as clearly shown by the illustrative simulation in Section 2.4.
The variance ratio tests on the Slasso residuals provide a diagnostic check on the m.d.s. condition required by our asymptotic theory. The results exhibit stark differences across data specifications: with untransformed data and without the lagged dependent variable, the tests strongly reject the m.d.s. condition ($p$-values ranging from 0.001 to 0.046), suggesting potential model misspecification. Including the lagged return or using transformed data substantially improves the diagnostic results, with $p$-values generally exceeding 0.05. This pattern suggests that both data transformation and including dynamics help satisfy the underlying assumptions, thereby boosting credibility for the XDlasso results under these specifications.
5.2 Predictability of Inflation Using Unemployment Rate
It is essential for monetary policymakers to understand the relationship between unemployment and inflation. As engemann2020phillips pointed out, “The Federal Reserve has a dual mandate to promote maximum sustainable employment and price stability.” First alluded to by fisher1926statistical; fisher1973discovered though, phillips1958relation popularized the Phillips curve — a negative relationship between the level of unemployment and the change rate of money wage rates. There has been a prolonged debate about whether the unemployment is a credible barometer for inflation among not only modern economic studies, but also policymakers. [^6] Mary Daly, San Francisco Fed President, delivered at daly2019new a negative view on the Phillips curve that “As for the Phillips curve… most arguments today center around whether it’s dead or just gravely ill. Either way, the relationship between unemployment and inflation has become very difficult to spot.” John Williams, New York Fed President, expressed a different opinion that “The Phillips curve is the connective tissue between the Federal Reserve’s dual mandate goals of maximum employment and price stability. Despite regular declarations of its demise, the Phillips curve has endured. It is useful, both as an empirical basis for forecasting and for monetary policy analysis.” See engemann2020phillips for more details. Empirical findings suggest that inflation rate can be either positively or negatively correlated with unemployment, depending on the shocks to the economies, the policies, and the lag orders (niskanen2002death; gordon2011history; gordon2013phillips). Given the ongoing debate, we revisit the Phillips curve in a predictive regression framework utilizing the FRED-MD dataset.
5.2.1 Data
The inflation rate, as the outcome variable in the predictive regression, is calculated by $\text{Inflation}_{t}=\left(\log({\rm CPI}_{t})-\log({\rm CPI}_{t-1})\right)\times 100,$ where ${\rm CPI}_{t}$ denotes the Consumer Price Index for All Urban Consumers: All Items (CPI). The unemployment rate, as the predictor of interest, denoted as $\text{Unrate}_{t}$, is retrieved in its original form under the name UNRATE.
Similar to our first empirical application in Section 5.1, we follow benati2015long to delineate three subperiods in addition to the full sample: Pre-Volcker (January 1960 - July 1979), Volcker and Greenspan (August 1979 - January 2006), and Bernanke, Yellen, and Powell (February 2006 - April 2025). These subperiods correspond to different eras in U.S. monetary policy, each named after the Federal Reserve chairperson who presided during that time. [^7] Pre-Volcker (before August 1979): This period was characterized by high and volatile inflation, with the Federal Reserve lacking a clear nominal anchor. Volcker and Greenspan (August 1979–January 2006): Volcker, who served as the Federal Reserve Chairman from August 1979 to August 1987, implemented aggressive anti-inflation measures, notably raising interest rates to historically high levels. Greenspan succeeded Volcker and continued to focus on maintaining price stability during his tenure, which contributed to a period of low and stable inflation known as the “Great Moderation”. Bernanke, Yellen, and Powell (February 2006–December 2019): Bernanke’s tenure as Fed Chairman was marked by the Great Recession and the implementation of unconventional monetary policies, such as quantitative easing, aimed at stimulating the economy and preventing deflation. Yellen and Powell carried on these policies. This periodization allows us to examine how the relationship between inflation and unemployment may have evolved across different policy regimes and economic conditions.
Figure 5 plots the inflation and unemployment rates over our sample period. Visual inspection suggests that the unemployment rate is more persistent than the inflation rate. Table 7 further reports the AR(1) coefficient estimates and ADF test $p$-values of the inflation and the unemployment rate under different sample periods. The inflation rate appears stationary in most periods, with AR(1) coefficient estimates ranging from 0.54 to 0.64. However, the inflation rate during the pre-Volcker period shows a slight upward trend, and nonstationarity is indicated by the ADF test with a $p$-value of 0.22. In contrast, the inflation rate is found to be stationary in the other two subperiods. The unemployment rate, on the other hand, appears highly persistent with AR(1) coefficient estimates close to 1. The ADF test rejects the null hypothesis of nonstationarity at a 10% significance level for both the full sample and the Volcker-Greenspan period. However, during the Pre-Volcker and Bernanke-Yellen-Powell periods, there is strong evidence of nonstationarity in the unemployment rate. The stationarity of both inflation and the unemployment rate in different periods is again ambiguous, which prompts the use of XDlasso from an agnostic perspective.
Figure 5: Inflation and Unemployment Rate
Table 7: Persistence of Inflation Rate and Unemployment Rate
Sample Period Inflation Rate Unemployment Rate
AR(1) ADF $p$-value AR(1) ADF $p$-value
Full Sample (Jan. 1960 - Apr. 2025) 0.620 0.0183 0.999 0.074
Pre-Volcker (Jan. 1960 - July 1979) 0.640 0.220 0.986 0.406
Volcker-Greenspan (Aug. 1979 - Jan. 2006) 0.623 <<0.01 0.992 0.071
Bernanke-Yellen-Powell (Feb. 2006 - Apr. 2025) 0.537 <<0.01 0.939 0.325
Notes: The lag order for the ADF test is set to $\lfloor n^{1/3}\rfloor$ where $n$ is the effective sample size. The exact start and end dates of subperiods are provided in the first column of the table.
In addition to the unemployment rate, we incorporate all other 110 macroeconomic variables from the FRED-MD dataset as controls after the TCODE transformation. Still, a significant proportion, from 11% to 17% across subperiods, of the transformed variables exhibit nonstationarity according to the ADF test at the 5% significance level. Consistent with Section 5.1, we also perform the analysis using the original (untransformed) FRED-MD time series as covariates. To further address potential concerns regarding the high correlation between control variables and the unemployment rate, we additionally present results excluding variables from the labor market group, as classified by FRED-MD.
5.2.2 Results
We study, in our predictive regression framework, a one-month-ahead regression $\text{Inflation}_{t}=\alpha^{*}+\theta_{1}^{*}\times\text{Unrate}_{t-1}+W_{-1,t-1}^{\top}\theta_{-1}^{{}^{*}}+u_{t},$ where $W_{-1,t-1}$ denotes high-dimensional covariates. Given the ambiguous persistence of the inflation rate discussed in the previous section, including the lagged dependent variable in our model may introduce further technical complications. Therefore, we do not recommend its inclusion in this analysis, and will focus on the traditional predictive regression setting. We carry out hypothesis testing on the key parameter of interest $\theta_{1}^{*}$ that measures the predictive power of the unemployment rate in forecasting inflation. The predictive form is of particular interest among policymakers in leveraging the relationship as a practical tool.
Table 8: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ across sample periods and specifications in inflation prediction
(a) TCODE Transformed Data
Sample Period IVX Dlasso XDlasso VR Test
Full Sample -0.014 0.018*** -0.024 0.000
(Jan. 1960 - Apr. 2025) (0.018) (0.006) (0.077)
Pre-Volcker 0.080 0.074*** 0.013 0.125
(Jan. 1960 - Jul. 1979) (0.069) (0.017) (0.224)
Volcker-Greenspan 0.036 -0.020 0.161 0.025
(Aug. 1979 - Jan. 2006) (0.063) (0.020) (0.118)
Bernanke/Yellen/Powell -0.043 -0.002 -0.054 0.002
(Feb. 2006 - Apr. 2025) (0.027) (0.011) (0.120)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.077 0.068 0.080 -0.050*** 0.033 0.065
(Jan. 1960 - Apr. 2025) (0.069) (0.218) (0.015) (0.032)
Pre-Volcker -0.129 -0.014 0.522 0.007 0.113 0.548
(Jan. 1960 - Jul. 1979) (0.117) (0.276) (0.051) (0.098)
Volcker-Greenspan 0.094 -0.272 0.741 -0.092* -0.259 0.339
(Aug. 1979 - Jan. 2006) (0.217) (0.287) (0.054) (0.228)
Bernanke-Yellen-Powell 0.550 0.585 0.283 0.001 0.040 0.230
(Feb. 2006 - Apr. 2025) (0.442) (0.975) (0.055) (0.073)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The upper and lower panels corresponds to the scenarios where we use TCODE transformed or untransformed time series as covariates, respectively. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table 8 reports the point estimates and standard errors for $\theta_{1}^{*}$ using IVX, Dlasso, and XDlasso across sample periods and specifications. In the benchmark setup with the transformed data, the diagnostic check rejects the m.d.s. condition except for the Pre-Volcker period. Nevertheless, Dlasso, which ignores nonstationarity, delivers significantly positive coefficients for both the full sample and the Pre-Volcker period, in striking contrast to XDlasso and IVX. Using untransformed data significantly alleviates concerns about model misspecification, with $p$-values of the diagnostic test greater than 5% in all cases. Both XDlasso and IVX consistently find no significant predictive power of unemployment for inflation across all specifications and time periods. The empirical findings add new insight to the recent debates on the Phillips curve and echo Mankiw2024’s latest pessimistic remark: “The large confidence intervals for the natural rate, together with the apparent futility of this Holy Grail search, lead me to think that we should not expect much from the Phillips curve as a guide for forecasting inflation or for judging the stance of policy.” On the other hand, with untransformed data, Dlasso yields significantly negative coefficients at the 1% level for the full sample and at the 10% level in the Volcker-Greenspan period, without controlling for other labor market variables. The unstable results of Dlasso across setups highlight the necessity of accounting for nonstationarity in the inference.
6 Conclusion
This paper proposes XDlasso to overcome the difficulties in hypothesis testing for high dimensional predictive regressions with stationary and nonstationary regressors. XDlasso fuses the IVX technique from time series econometrics and the debiasing technique from the high dimensional statistics, thereby reducing the order of biases to make them readily correctable. We establish the asymptotic normality and convergence rate of XDlasso. The validity of our methods is further evidenced by simulation studies and empirical applications.
Appendix A Technical Proofs
In the proofs, we use $c$ and $C$, without superscripts or subscripts, to denote generic positive constants that may vary place to place. For any positive sequences $\{a_{n}\}$ and $\{b_{n}\}$, $a_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}b_{n}$” means that there is an absolute constant, say $c$, such that the event $\left\{a_{n}\leq cb_{n}\right\}$ holds with probability approaching one (w.p.a.1.). Symmetrically, $a_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\succcurlyeq}}b_{n}$” means $b_{n}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}a_{n}$”. The integer floor function is denoted as $\left\lfloor\cdot\right\rfloor$. For an $n$-dimensional vector $x=(x_{t})_{t\in[n]}$, the $L_{2}$-norm is $\left\|x\right\|_{2}=\sqrt{\sum_{t=1}^{n}x_{t}^{2}}$. For notational simplicity, in the proofs we assume $p\geq n^{\nu_{1}}$ for some absolute constant $\nu_{1}$, which is reasonable as we focus on the high dimensional case with a larger $p$ relative to $n$.A.1A.1A.1There is no technical difficulty in allowing $p$ to grow either slowly at a logarithmic or fast at an exponential rate of $n$, but we have to compare $\log p$ and $\log n$ in many places, and in many conditions and rates the term $\log p$” has to be changed into $\log(np)$.We use $\mathbf{I}_{n}$ to denote the $n$-dimensional identity matrix, where the subscript may be omitted when there is no ambiguity with matrix dimensions.
Section A.1 proves the results in Section 3.1 for the consistency of Slasso at the presence of both LUR and stationary regressors. Section A.2 includes the proofs for the results in Section 3.2 that constructs consistency of the auxiliary LASSO regression adapted for bias correction. Section A.3 includes the proofs for the theorems in Section 3.3 about the asymptotic distribution of the XDlasso estimator and the order of its standard error, which determine the size and power of the XDlasso inference.
A.1 Proofs for Section 3.1
A.1.1 Technical Lemmas of Gaussian Approximation
The assumptions imposed in Section 3 slightly differ from those in mei2022lasso (MS24, hereafter). Specifically, the linear process assumption in MS24 is extended to more general $\alpha$-mixing and sub-exponential conditions. One of the main modifications of the proof is the Gaussian approximation error deduced in MS24’s Lemma B.4. The following lemma re-establishes Gaussian approximation errors for the stationary components $v_{t}$ defined as (3.1). The results of Gaussian approximation will be useful for the RE condition required for the Slasso’s consistency.
Lemma A.1.
Under Assumptions 14, there exist standard Brownian motions $\{\mathcal{B}_{k}(t)\}$ with independent increment $\mathcal{B}_{k}(t)-\mathcal{B}_{k}(s)\sim\mathcal{N}(t-s)$ for $t\geq s\geq 0$ such that
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{k}^{*})\right)\right|=O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
where $V_{k}^{*}=\mathbb{E}\left[\sum_{d=-\infty}^{\infty}\varepsilon_{k,t}\varepsilon_{k,t-d}\right]$ is the long-run variance of $\{\varepsilon_{k,t}\}.$
Remark A.1.
The convergence rate in Lemma A.1 is less sharp than that in MS24’s Lemma B.4, since we work with general $\alpha$-mixing sequences without specifying linear processes.
Proof of Lemma A.1.
Define $V_{k,t}=\mathbb{E}\left[\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}\right)^{2}\right].$ The proof includes the following two steps:
  1. 1.
    The variance of the partial sum $V_{k,t}$ is well approximates by the long run variance $V_{k}^{*}$ scale by $t$, in the sense that
    $\sup_{k\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|$ $=O\left(\dfrac{1}{n}\right).$ (A.1)
  2. 2.
    Strong Gaussian approximation: There exist standard Brownian motions $\{\mathcal{B}_{k}(t)\}$ with independent increment $\mathcal{B}_{k}(t)-\mathcal{B}_{k}(s)\sim\mathcal{N}(t-s)$ for $t\geq s\geq 0$ such that
    $\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|=O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right).$ (A.2)
Given the two steps above, Lemma A.1 follows by the triangular inequality
$\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|$
$\leq$ $\sup_{j\in[p],t\in[n]}\left|\dfrac{1}{\sqrt{n}}\left(\mathcal{B}_{k}(V_{k,t})-\mathcal{B}_{k}(t\cdot V_{j}^{*})\right)\right|+\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|$
$=$ $O_{p}\left(\sqrt{\log(np)\cdot\sup_{k\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|}\right)+\sup_{k\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|$
$=$ $O_{p}\left(\sqrt{\dfrac{\log(np)}{n}}\right)+O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right)$
$=$ $O_{p}\left(\dfrac{(\log p)^{3/2}}{n^{1/4}}\right)$
where the second row applies the fact that $n^{-1/2}\left(\mathcal{B}_{k}(V_{k,t})-\mathcal{B}_{k}(t\cdot V_{k}^{*})\right)$ for each $(j,t)$ follows a normal distribution of mean zero and variance $\left|\left(t\cdot V_{k}^{*}-V_{k,t}\right)/n\right|$, the third row applies (A.1), (A.2), and the last row applies the assumption $p\geq n^{\nu_{1}}$.
Step 1. Verifying (A.1). Define ${\rm cov}_{k}(d):={\rm cov}(\varepsilon_{k,t},\varepsilon_{k,t-d})$ as the autocovariance function of $\{\varepsilon_{k,t}\}$. Then by some fundamental calculations,
$V_{k,t}=t\cdot{\rm cov}_{k}(0)+2\sum_{d=1}^{t}(t-d)\cdot{\rm cov}_{k}(d).$ (A.3)
In addition,
$t\cdot V_{k}^{*}=t\cdot{\rm cov}_{k}(0)+2t\cdot\sum_{d=1}^{\infty}{\rm cov}_{k}(d).$
Then
$t\cdot V_{k}^{*}-V_{k,t}$ $=2t\sum_{d=1}^{\infty}{\rm cov}_{k}(d)-2\sum_{d=1}^{t}(t-d)\cdot{\rm cov}_{k}(d)$
$=2t\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)+2\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d).$ (A.4)
Recall that Assumption 2 imposes an upper bound for the $\alpha$-mixing coefficient. With $p=q=3$ in Equation (2.2) in the Corollary of davydov1968convergence, we have
$\sup_{j\in[p]}|{\rm cov}_{k}(d)|\leq 12\left(\mathbb{E}|\varepsilon_{k,t}|^{3}\right)^{2/3}\sqrt{\alpha(d)}\leq C\cdot\exp\left(-c_{\alpha}d^{r}/2\right).$ (A.5)
where $C=12\left(\mathbb{E}|\varepsilon_{k,t}|^{3}\right)^{2/3}\sqrt{C_{\alpha}}$ and $C_{\alpha}$ is in Assumption 2. By Equation (B.78) in MS24’s supplement,
$\sup_{t\in[n]}\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\leq\dfrac{2}{c_{\alpha}}\exp\left(-c_{\alpha}d^{r}/4\right),$
and thus uniformly for all $t,$ there exists some absolute constant $C_{1}$ such that
$\sup_{j\in[p],t\in[n]}t\cdot\sum_{d=t+1}^{\infty}|{\rm cov}_{k}(d)|\leq C\cdot\sup_{t\in[n]}t\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\leq\sup_{t\in[n]}\dfrac{4Ct}{c_{\alpha}}\exp\left(-c_{\alpha}t^{r}/4\right)<C_{1}.$ (A.6)
In addition,
$\sup_{j\in[p],t\in[n]}\sum_{d=1}^{t}d\cdot|{\rm cov}_{k}(d)|<C\sum_{d=1}^{\infty}d\cdot\exp\left(-c_{\alpha}d^{r}/2\right)<C_{2}$ (A.7)
for some absolute constant $C_{2}$. Then
$\sup_{j\in[p],t\in[n]}\left|\dfrac{t\cdot V_{k}^{*}-V_{k,t}}{n}\right|$ $\leq\dfrac{2}{n}\cdot\left(\sup_{j\in[p],t\in[n]}t\cdot\sum_{d=t+1}^{\infty}|{\rm cov}_{k}(d)|+\sup_{j\in[p],t\in[n]}\sum_{d=1}^{t}d\cdot|{\rm cov}_{k}(d)|\right)$
$\leq\dfrac{2(C_{1}+C_{2})}{n},$
which implies (A.1).
Step 2. Verifying (A.2). We use the strong Gaussian approximation from lin1997limit’s Theorem 9.3.1. Specifically, define $g(x)=\exp(x)$. By the sub-exponential tail imposed by Assumption 1, the sub-exponential norm of $\varepsilon_{k,t}$, denoted as $\|\varepsilon_{k,t}\|_{g}$ in lin1997limit, is uniformly bounded by an absolute constant. It then suffices to verify the following two conditions required in the aforementioned theorem:
  • (i)
    $V_{k,t}\geq ct$ for some absolute constant $c$.
  • (ii)
    $\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))<\infty$, where the parameter $\delta$ in lin1997limit is taken as 2.
Then by Theorem 9.3.1 of lin1997limit, for any $j\in[p]$
$\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})=O_{a.s.}\left(V_{k,t}^{1/4}(\log V_{k,t})^{3/2}\right).$ (A.8)
By (A.3) and (A.5),
$V_{k,t}$ $=t\left[{\rm cov}_{k}(0)+2\sum_{d=1}^{t}(1-d/t)\cdot{\rm cov}_{k}(d)\right]$
$\leq t\left[{\rm cov}_{k}(0)+2C\cdot\sum_{d=1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)\right]=O(t)$
uniformly for all $(k,t)$. Then by (A.8),
$\sup_{j\in[p],t\in[n]}\left|\frac{1}{\sqrt{n}}\left(\sum_{s=0}^{t-1}\varepsilon_{k,s}-\mathcal{B}_{k}(V_{k,t})\right)\right|=O_{a.s.}\left(\sup_{t\in[n]}\dfrac{t^{1/4}(\log t)^{3/2}}{\sqrt{n}}\right)=O_{a.s.}\left(\dfrac{(\log n)^{3/2}}{n^{1/4}}\right),$
which leads to (A.2). It then suffices to show the Conditions (i) and (ii) above.
Proof of (i). By (A.4) and (A.5),
$\lim_{t\to\infty}\sup_{k\in[p]}\left|\dfrac{V_{k,t}}{t}-V_{k}^{*}\right|$ $\leq\lim_{t\to\infty}2\sup_{k\in[p]}\left|\sum_{d=t+1}^{\infty}{\rm cov}_{k}(d)+\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\right|$
$\leq\lim_{t\to\infty}C\left[\sum_{d=t+1}^{\infty}\exp\left(-c_{\alpha}d^{r}/2\right)+\dfrac{2}{t}\sum_{d=1}^{t}d\cdot{\rm cov}_{k}(d)\right]$
$=0,$
where $C$ is an absolute constant, and the limit applies (A.6) and (A.7). By Assumption 2, the long-run variance $V_{k}^{*}$ is bounded from below by some absolute constant. This result implies that $V_{k,t}/t$ is lower bounded by some absolute constant uniformly for all $(k,t)$.
Proof of (ii). This is a direct corollary of the exponential decaying mixing coefficient imposed by Assumption 2, in the sense that
$\sum_{d=1}^{\infty}\alpha(d)^{1/4}\cdot\log(1/\alpha(d))\leq\sum_{d=1}^{\infty}C_{\alpha}^{1/4}\exp\left(-c_{\alpha}d^{r}/4\right)\cdot c_{\alpha}d^{r}<\infty.$
This completes to proof of Lemma A.1. ∎
The next lemma establishes that result that the LUR regressors $X_{t}$ with general weakly dependent innovations can be approximated by another vector LUR processes with normally distributed innovations.
Lemma A.2.
Suppose that Assumptions 14 hold. There exists independent normally distributed variables $\eta_{j,t}$ for all $j\in\mathcal{M}_{x}$, such that the LUR processes $\xi_{j,t}=\rho_{j}^{*}\xi_{j,t-1}+\sum_{k=1}^{p}\Phi_{j,k}\eta_{k,t}$ satisfy
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|x_{j,t}-\xi_{j,t}\right|=O_{p}(n^{1/4}\log p),$
where $\Phi_{j,k}$ is the $(j,k)$-th entry of the matrix $\Phi$ in (3.1).
Proof of Lemma A.2.
Note that $x_{j,t}-x_{j,t-1}=e_{j,t}+\dfrac{c_{j}^{*}}{n}x_{j,t-1}$. Without loss of generality, assume that $x_{j,0}=0.$ Then
$x_{j,t}=\sum_{s_{1}=1}^{t}(x_{j,t}-x_{j,t-1})=\sum_{s_{1}=1}^{t}e_{j,s_{1}}+\dfrac{c_{j}^{*}}{n}\sum_{s_{1}=1}^{t}x_{j,s_{1}-1}.$
By induction, we have for any fixed integer $M$, whenever $t>M$
$x_{j,t}$ $=\sum_{s_{1}=1}^{t}e_{j,s_{1}}+\dfrac{c_{j}^{*}}{n}\sum_{s_{1}=1}^{t}\left(\sum_{s_{2}=2}^{s_{1}}e_{j,s_{2}-1}+\dfrac{c_{j}^{*}}{n}\sum_{s_{2}=2}^{s_{1}}x_{j,s_{2}-2}\right)$
$=\sum_{s_{1}=1}^{t}e_{j,s_{1}}+\dfrac{c_{j}^{*}}{n}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}e_{j,s_{2}-1}+\dfrac{c_{j}^{*2}}{n^{2}}\sum_{s=1}^{t}\sum_{s_{2}=2}^{s_{1}}x_{j,s_{2}-2}$
$=...=\sum_{\ell=1}^{M}\left(\dfrac{c_{j}^{*}}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}e_{j,s_{\ell}-\ell+1}+\left(\dfrac{c_{j}^{*}}{n}\right)^{M}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}x_{j,s_{M}-M},$
where we define $s_{0}=t$. By Assumption 3, we have $e_{j,t}=\sum_{k=1}^{p}\Phi_{j,k}\varepsilon_{k,t}$, and therefore
$x_{j,t}=\sum_{k=1}^{p}\Phi_{j,k}\sum_{\ell=1}^{M}\left(\dfrac{c_{j}^{*}}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}\varepsilon_{k,s_{\ell}-\ell+1}+\left(\dfrac{c_{j}^{*}}{n}\right)^{M}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}x_{j,s_{M}-M}.$ (A.9)
Let $\mathcal{B}_{j}$ denote the Brownian motion in the Gaussian approximation of Lemma A.1, and define $\eta_{k,t}=\mathcal{B}_{k}(tV_{k}^{*})-\mathcal{B}_{j}((t-1)V_{k}^{*}).$ Then $\{\eta_{j,t}\}$ are i.i.d. distributed, and
$\sup_{k\in[p],t\in[n]}\dfrac{1}{\sqrt{n}}\left|\sum_{s=0}^{t-1}\varepsilon_{k,s}-\sum_{s=0}^{t-1}\eta_{k,s}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\log p}{n^{1/4}}.$ (A.10)
Let $\xi_{j,t}$ be an LUR satisfying $\xi_{j,t}=\rho_{j}^{*}\xi_{j,t-1}+\sum_{k=1}^{p}\Phi_{j,k}\eta_{k,t}$, where $\rho_{j}^{*}=1-c_{j}^{*}/n$ is the same as the AR coefficient of the LUR regressor $x_{j,t}$. Following the same arguments for (A.9), we have
$\xi_{j,t}=\sum_{k=1}^{p}\Phi_{j,k}\sum_{\ell=1}^{M}\left(\dfrac{c_{j}^{*}}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}\eta_{k,s_{\ell}-\ell+1}+\left(\dfrac{c_{j}^{*}}{n}\right)^{M}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}\xi_{j,s_{M}-M}.$
Thus,
$x_{j,t}-\xi_{j,t}=A_{j,t}^{(1)}+A_{j,t}^{(2)},$ (A.11)
$A_{j,t}^{(1)}:=\sum_{k=1}^{p}\Phi_{j,k}\sum_{\ell=1}^{M}\left(\dfrac{c_{j}^{*}}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}(\varepsilon_{k,s_{\ell}-\ell+1}-\eta_{k,s_{\ell}-\ell+1}),$
$A_{j,t}^{(2)}:=\left(\dfrac{c_{j}^{*}}{n}\right)^{M}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}(x_{j,s_{M}-M}-\xi_{j,s_{M}-M}).$
We first bound $A_{j,t}^{(1)}.$ Recall that $|c_{j}^{*}|\leq\bar{c}$ for all $j$ by Assumption 4., and thus $|c_{j}^{*}|^{\ell-1}\leq\max\{\bar{C}^{M},1\}.$ Therefore,
$\left|A_{j,t}^{(1)}\right|\leq$ $\left|\sum_{k=1}^{p}\Phi_{j,k}\right|\cdot\max\{\bar{C}^{M},1\}\cdot\sup_{k\in[p]}\sum_{\ell=1}^{M}\left(\dfrac{1}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{(\ell-1)}=\ell-1}^{s_{(\ell-2)}}\left|\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}(\varepsilon_{k,s_{\ell}}-\eta_{k,s_{\ell}})\right|.$
Each of the summations in $\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{(\ell-1)}=\ell-1}^{s_{(\ell-2)}}$” involves no more than $n^{\ell-1}$ terms. Therefore,
$\left(\dfrac{1}{n}\right)^{\ell-1}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{(\ell-1)}=\ell-1}^{s_{(\ell-2)}}\left|\sum_{s_{\ell}=\ell}^{s_{(\ell-1)}}(\varepsilon_{k,s_{\ell}-\ell+1}-\eta_{k,s_{\ell}-\ell+1})\right|\leq\max_{t\in[n]}\left|\sum_{s=1}^{t}(\varepsilon_{k,s}-\eta_{k,s})\right|,$
which implies
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|A_{j,t}^{(1)}\right|\leq$ $\sup_{j\in[p]}\left|\sum_{k=1}^{p}\Phi_{j,k}\right|\cdot\max\{\bar{c}^{M},1\}\cdot\sup_{k\in[p]}\sum_{\ell=1}^{M}\sup_{t\in[n]}\left|\sum_{s=1}^{t}(\varepsilon_{k,s}-\eta_{k,s})\right|$
$\leq$ $C_{M}\sup_{k\in[p],t\in[n]}\left|\sum_{s=1}^{t}(\varepsilon_{k,s}-\eta_{k,s})\right|,$ (A.12)
where $C_{M}$ is an absolute constant dependent on the integer $M$ only.
We then bound $A_{j,t}^{(2)}.$ Note that
$\left|A_{j,t}^{(2)}\right|\leq\dfrac{\max\{\bar{C}^{M},1\}}{n^{M}}\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}\left|x_{j,s_{M}-1}-\xi_{j,s_{M}-1}\right|.$
The summations $\sum_{s_{1}=1}^{t}\sum_{s_{2}=2}^{s_{1}}\cdots\sum_{s_{M}=M}^{s_{(M-1)}}$” involve no more than $\binom{t}{M}\leq\binom{n}{M}$ terms, and
$\lim_{n\to\infty}\dfrac{1}{n^{M}}\binom{n}{M}=\dfrac{1}{M!}.$
Therefore, when $n$ is large enough,
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|A_{j,t}^{(2)}\right|\leq\dfrac{2\max\{\bar{C}^{M},1\}}{M!}\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}-\xi_{j,t}|.$
Note that the upper bound holds for any fixed $M$. Let $M$ be sufficiently large so that $\dfrac{2\max\{\bar{C}^{M},1\}}{M!}<0.5$. Then
$\sup_{j\in\mathcal{M}_{x},t\in[n]}\left|A_{j,t}^{(2)}\right|\leq 0.5\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}-\xi_{j,t}|.$ (A.13)
By (A.11), (A.12), (A.13),
$\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}-\xi_{j,t}|\leq C_{M}\sup_{k\in[p],t\in[n]}\left|\sum_{s=1}^{t}(\varepsilon_{k,s}-\eta_{k,s})\right|+0.5\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}-\xi_{j,t}|,$
which implies
$\sup_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}-\xi_{j,t}|\leq 2C_{M}\sup_{k\in[p],t\in[n]}\left|\sum_{s=1}^{t}(\varepsilon_{k,s}-\eta_{k,s})\right|.$
Then Lemma A.2 is implied by Lemma A.1. ∎
A.1.2 Technical Lemmas for DB and RE
Define the sample Gram matrices of the LURs and stationary regressors as $\widehat{\Sigma}^{(x)}=\frac{1}{n}\sum_{t=1}^{n}X_{t-1}X_{t-1}^{\top}$ and $\widehat{\Sigma}^{(z)}=\frac{1}{n}\sum_{t=1}^{n}Z_{t-1}Z_{t-1}^{\top}$, respectively. The following lemma shows that after standardization, the Gram matrix of all regressors $\widehat{\Sigma}$ can be approximated by the block-diagonal matrix
$\widehat{\Delta}={\rm diag}\left(\widehat{\Sigma}^{(z)},\widehat{\Sigma}^{(x)}\right).$
Lemma A.3.
Under Assumptions 1-4,
$\|D^{-1}\left(\widehat{\Sigma}-\widehat{\Delta}\right)D^{-1}\|_{\infty}=O_{p}\left(\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\right)$ (A.14)
as $n\to\infty,$ where $r$ is specified in Assumption 2.
Proof of Lemma A.3.
By Lemma A.8,
$\dfrac{1}{\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}\min_{\ell\in\mathcal{M}_{z}}\widehat{\sigma}_{\ell}}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{1}{\sqrt{\log p}/\sqrt{n}}=\sqrt{\dfrac{n}{\log p}}.$ (A.15)
Therefore,
$\|D^{-1}\left(\widehat{\Sigma}-\widehat{\Delta}\right)D^{-1}\|_{\infty}$ $\leq\dfrac{\|n^{-1}\sum_{t=1}^{n}x_{t-1}z_{t-1}^{\top}\|_{\infty}}{\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}\min_{\ell\in\mathcal{M}_{z}}\widehat{\sigma}_{\ell}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{1+\frac{1}{2r}}}{\sqrt{n/\log p}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}=O_{p}\left(\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\right),$
where the second row applies (A.15) and (A.38). ∎
The following Lemma establishes the RE condition for LURs without standardization. We focus on the demeaned regressors $x_{t-1}-\bar{x}$, since the results will be helpful to bound the standard deviations used in Slasso. Following (B.33) and (B.34) of MS, define
$C_{m}(L)=\lceil 4L^{2}\widetilde{C}/\widetilde{c}\rceil\text{ and }m=C_{m}s,$
where $\widetilde{C}$ and $\widetilde{c}$ are absolute constants following the definitions between (B.33) and (B.34) of MS.
Lemma A.4.
Suppose that $(1+C_{m}(L))s=o(n\wedge p)$ as $n\to\infty$, and $k=1.$ Define $\ddot{\Sigma}^{(x)}=n^{-1}\sum_{t=1}^{n}(x_{t-1}-\bar{x})(x_{t-1}-\bar{x})^{\top}.$ Then under Under Assumptions 1-4,
$\frac{\kappa_{\mathbf{I}}(\ddot{\Sigma}^{(x)},L,s)}{n}\geq\dfrac{c_{\kappa}}{L^{2}s\log p}$ (A.16)
holds w.p.a.1. for any $L\geq 1$.A.2A.2A.2Here we use a generic $L\geq 1$ is useful for deducing the lower bound of $\widehat{\kappa}_{D}$ using $\widehat{\kappa}_{\mathbf{I}}$.
Proof of Lemma A.4.
To simplify the proof, we assume all regressors are LURs, and the values of all time series at $t=0$ are zeros.
(a) We first impose the normality assumption $\varepsilon_{t}\sim i.i.d.\ \mathcal{N}(0,\mathbf{I}_{p})$. It implies $e_{t}\sim i.i.d.\ \mathcal{N}(0,\Omega_{e})$ with $\Omega_{e}=\Phi_{e}\Phi_{e}^{\top}.$ Note that for the LUR cases,
$X_{t}-X_{t-1}=\dfrac{\mathbf{C}}{n}X_{t-1}+e_{t}$
for any $t\geq 1$, where $\mathbf{C}={\rm diag}(c_{1}^{*},c_{2}^{*},\dots,c_{p}^{*})$. Define
$e_{t}^{\Delta}=\begin{cases}\dfrac{\mathbf{C}}{n}X_{t-1}+e_{t},&t\geq 1,\\ 0,&t=0,\end{cases}$ (A.17)
and note that $X_{t}=\sum_{s=1}^{t}e_{s}^{\Delta}.$ Let $R$ be an $n\times n$ lower triangular matrix of ones on and below the diagonal. Define $X=(X_{0},X_{1},\dots,X_{n-1})^{\top}$, $e=(e_{0},e_{1},\dots,e_{n-1})^{\top}$ and $e^{\Delta}=(e_{0}^{\Delta},e_{1}^{\Delta},\dots,e_{n-1}^{\Delta})^{\top}$. Note that $\underset{(n\times p)}{X}=\underset{(n\times n)}{R}\underset{(n\times p)}{e^{\Delta}}$. We decompose we write
$\ddot{\Sigma}=n^{-1}X^{\top}X=n^{-1}e^{\Delta}{}^{\top}R^{\top}Re^{\Delta}.$
[…]
Following the technique used to prove Remark 3.5 in zhang2019identifying, which is also used for Theorem B.2 in smeekes2021automated, we diagonalize $R(\mathbf{I}_{n}-\mathbf{J}_{n})R^{\top}=V{\rm diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})V^{\top}$, where $V$ is an orthonormal matrix. For any $\delta\in\mathbb{R}^{p}$, $\delta\neq 0$, the quadratic form
$\delta^{\top}\ddot{\Sigma}\delta$ $=\dfrac{1}{n}e^{\Delta\top}R^{\top}Re^{\Delta}=\frac{1}{n}\delta^{\top}e^{\Delta\top}V{\rm diag}(\lambda_{1},\lambda_{2},\cdots,\lambda_{n})V^{\top}e^{\Delta}\delta$
$\geq\frac{1}{n}\delta^{\top}e^{\Delta\top}V_{\cdot[\ell]}{\rm diag}(\lambda_{1},\cdots,\lambda_{\ell})V_{\cdot[\ell]}^{\top}e^{(n)}\delta\geq\frac{\lambda_{\ell}}{n}\delta^{\top}e^{\Delta\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e^{\Delta}\delta$
$\geq\frac{\ell\widetilde{\lambda}_{\ell+1}}{n}\cdot\delta^{\top}\Gamma_{\ell}^{\Delta}\delta$ (A.18)
for any $\ell\in[n-1]$, where $V_{\cdot[\ell]}$ is the submatrix composed of the first $\ell$ columns of $V$ and $\Gamma_{\ell}=\ell^{-1}e^{\Delta\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e^{\Delta}$.
We first work with the first factor $\ell\lambda_{\ell}/n$ in (A.18). smeekes2021automated provide the exact formula of $\lambda_{\ell}$:
$\widetilde{\lambda}_{\ell+1}=\left[2\left(1-\cos\left(\dfrac{(2\ell+1)\pi}{2n+1}\right)\right)\right]^{-1}\text{ for all }\ell\in[n].$ (A.19)
A Taylor expansion of $\cos\left(x\pi\right)$ around $x=0$ yields
$\widetilde{\lambda}_{\ell+1}=\left(\dfrac{(2\ell+1)\pi}{2n+1}\right)^{2}\left(1+o\left(\frac{\ell+1}{n}\right)\right)=\left(\dfrac{\ell\pi}{n}\right)^{2}\left(1+o\left(\frac{\ell}{n}\right)\right)$
whenever $\ell=o\left(n\right)$. This implies
$\dfrac{\ell\widetilde{\lambda}_{\ell+1}}{n}=\dfrac{n}{\pi^{2}\ell\left(1+o\left(\ell/n\right)\right)}\geq\dfrac{n}{2\pi^{2}\ell}$ (A.20)
for $\ell=o\left(n\right)$ when $n$ is sufficiently large.
Next, we focus on the second factor $\delta^{\top}\Gamma_{\ell}^{\Delta}\delta$ in (A.18). Define $X_{\mathbb{L}}:=\left(0_{p},X_{0},X_{1},\cdots,X_{t-2}\right)^{\top}.$ By definition, we have
$e^{\Delta}=X_{\mathbb{L}}\dfrac{{\rm\mathbf{C}}}{n}+e.$
We deduce that
$\delta^{\top}\Gamma_{\ell}^{\Delta}\delta$ $=\dfrac{\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}+\dfrac{\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}X_{\mathbb{L}}{\rm\mathbf{C}}^{\top}\delta}{n^{2}\ell}+\dfrac{2\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{n\ell}$
$\geq\dfrac{\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}+\dfrac{2\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{n\ell}.$
Recall the generic inequality $2a^{\top}b\leq a^{\top}a+b^{\top}b$ for any vectors $a$ and $b$ of the same dimension. Let $a=\frac{V_{\cdot[\ell]}^{\top}e\delta}{\sqrt{2}}$ and $b=$ $\sqrt{2}n^{-1}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta$, we have
$\dfrac{2\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{n}\leq 0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta+\dfrac{2{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}X_{\mathbb{L}}{\rm\mathbf{C}}^{\top}}{n^{2}}.$
It implies
$\delta^{\top}\Gamma_{\ell}^{\Delta}\delta\geq\dfrac{0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}-\dfrac{2\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}X_{\mathbb{L}}{\rm\mathbf{C}}\delta}{n^{2}\ell}.$
In addition, $\lambda_{\max}(V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top})\leq\|V_{\cdot[\ell]}\|_{2}^{2}\leq 1$ given $V$ is a unitary matrix. Therefore,
$\delta^{\top}\Gamma_{\ell}^{\Delta}\delta$ $\geq\dfrac{0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}-\dfrac{2\delta^{\top}{\rm\mathbf{C}}X_{\mathbb{L}}^{\top}X_{\mathbb{L}}{\rm\mathbf{C}}\delta}{n^{2}\ell}$
$=\dfrac{0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}-\dfrac{2\delta^{\top}{\rm\mathbf{C}}\sum_{t=1}^{n-1}X_{t-1}X_{t-1}^{\top}{\rm\mathbf{C}}\delta}{n^{2}\ell}$
$\geq\dfrac{0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}-\dfrac{2\delta^{\top}{\rm\mathbf{C}}\sum_{t=1}^{n}X_{t-1}X_{t-1}^{\top}{\rm\mathbf{C}}\delta}{n^{2}\ell}$
$=\dfrac{0.5\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}-\dfrac{2\delta^{\top}{\rm\mathbf{C}}\widehat{\Sigma}{\rm\mathbf{C}}\delta}{n\ell}.$ (A.21)
Define $\Gamma_{\ell}=\dfrac{\delta^{\top}e{}^{\top}V_{\cdot[\ell]}V_{\cdot[\ell]}^{\top}e\delta}{\ell}.$ By (A.18), (A.20), and (A.21),
$\delta^{\top}\ddot{\Sigma}\delta$ $\geq\dfrac{n}{2\pi^{2}\ell}\left(0.5\delta^{\top}\Gamma_{\ell}\delta-2n^{-1}\ell^{-1}\delta^{\top}{\rm\mathbf{C}}\widehat{\Sigma}{\rm\mathbf{C}}\delta\right).$ (A.22)
We first lower bound the first term. Let $\ell=(16+C_{\ell})\cdot(s+m)\log p$ for some $C_{\ell}>0$ to be determined later. Following the proof of (B.43) in MS24 utilizing the non-asymptotic bounds for Whishart matrices, we have
$\delta^{\top}\Gamma_{\ell}\delta$ $\geq C_{\kappa}\|\delta\|_{2}^{2},$ (A.23)
w.p.a.1, where the absolute constant $C_{\kappa}$ not dependent on $L$ or $C_{\ell}$. We then bound the second term in (A.22). Note that for any $\delta\in\mathcal{R}(L,s)$ such that for any $|\mathcal{M}|\leq s$ we have $\|\delta_{\mathcal{M}^{c}}\|_{1}\leq L\|\delta_{\mathcal{M}}\|_{1},$
$\|\delta\|_{1}\leq\|\delta_{\mathcal{M}}\|_{1}+\|\delta_{\mathcal{M}^{c}}\|_{1}\leq(1+L)\|\delta_{\mathcal{M}}\|_{1}\leq(1+L)\sqrt{s}\|\delta\|_{2}.$ (A.24)
Therefore,
$\delta^{\top}{\rm\mathbf{C}}\widehat{\Sigma}{\rm\mathbf{C}}^{\top}\delta\leq\left(\|\delta\|_{1}\right)^{2}\cdot\|{\rm\mathbf{C}}\|_{1}^{2}\cdot\|\widehat{\Sigma}\|_{\infty}\leq(1+L)^{2}s\|\delta\|_{2}^{2}\cdot\|{\rm\mathbf{C}}\|_{1}^{2}\cdot\|\widehat{\Sigma}\|_{\infty}$.
Note that $x_{j,t}=\sum_{s=1}^{t}(\rho_{j}^{*})^{s}e_{t-s}$ is a partial sum of a stationary time series. By Lemma B.2 of MS, we have.
$\max_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}.$ (A.25)
Therefore, $\|\widehat{\Sigma}\|_{\infty}^{2}\leq\max_{j,t}|x_{j,t-1}|^{2}\leq C_{{\rm sup}}n\log p$ w.p.a.1 for some absolute constant $C_{{\rm sup}}$, which implies
$\delta^{\top}{\rm\mathbf{C}}\ddot{\Sigma}{\rm\mathbf{C}}^{\top}\delta$ $\leq(1+L)^{2}s\|\delta\|_{2}^{2}\cdot\|{\rm\mathbf{C}}\|_{1}^{2}\cdot\|\widehat{\Sigma}\|_{\infty}$
$\leq(1+L)^{2}s\|\delta\|_{2}^{2}\cdot\bar{C}^{2}\cdot C_{\sup}n\log p$
$\leq 4\bar{C}^{2}L^{2}\cdot s\|\delta\|_{2}^{2}\cdot C_{\sup}n\log p,$
where the second inequality applies $\|{\rm\mathbf{C}}\|_{1}\leq\sup_{j\in[p]}|c_{j}^{*}|\leq\bar{C}^{2}$, and the third inequality applies $L\geq 1$ and $\|\delta_{\mathcal{S}}\|_{1}\leq\sqrt{s}\|\delta\|_{2}.$ Recall that $\ell=(16+C_{\ell})\cdot(s+m)\log p$. Let $C_{\ell}=16\cdot(1\vee(L\bar{C}^{2}C_{\sup}/C_{\kappa}))-16$. Then
$\delta^{\top}{\rm\mathbf{C}}\ddot{\Sigma}{\rm\mathbf{C}}^{\top}\delta\leq\|\delta\|_{2}^{2}\cdot 0.25C_{\kappa}n\cdot(16+C_{\ell})s\log p\leq\|\delta\|_{2}^{2}\cdot 0.25C_{\kappa}n\ell.$ (A.26)
Insert (A.23) and (A.26) into (A.22), we have
$\delta^{\top}\ddot{\Sigma}\delta$ $\geq\dfrac{n}{2\pi^{2}\ell}\left(0.5C_{\kappa}-0.25C_{\kappa}\right)\|\delta\|_{2}^{2}=\dfrac{nC_{\kappa}}{8\pi^{2}\ell}\|\delta\|_{2}^{2}.$
By $\ell=(16+C_{\ell})\cdot(s+m)\log p$, $m=\lceil 4L\widetilde{C}/\widetilde{c}\rceil s,$ and $L\geq 1$
$\frac{\delta^{\top}\text{$\widehat{\Sigma}$}\delta}{n\|\delta\|_{2}^{2}}$ $\geq\dfrac{C_{\kappa}}{8\pi^{2}(16+C_{\ell})\cdot(s+m)\log p}$
$\geq\dfrac{C_{\kappa}}{8\pi^{2}(16+C_{\ell})\cdot(1+\lceil 4L\widetilde{C}/\widetilde{c}\rceil)s\log p}$ (A.27)
$\geq\dfrac{C_{\kappa}}{8\pi^{2}\cdot 16\cdot(1\vee(L\bar{C}^{2}C_{\sup}/C_{\kappa}))\cdot(8L\widetilde{C}/\widetilde{c})\cdot s\log p}$
$\geq\dfrac{\widetilde{c}_{\kappa}}{L^{2}\cdot s\log p}$
w.p.a.1, where $\widetilde{c}_{k}$ is an absolute constant dependent on $C_{\kappa},C_{{\rm sup}},\widetilde{C},\widetilde{c}$, and $\bar{C}$. Then (A.16) holds.
(b) We then extend the result to non-normal errors. Let $\ddot{\Upsilon}=n^{-1}\sum_{t=1}^{n}(\xi_{t-1}-\bar{\xi})(\xi_{t-1}-\bar{\xi})^{\top}$, where $\xi_{t}=(\xi_{j,t})_{j\in\mathcal{M}_{x}}$is the vector of LUR processes with normally distributed errors as in Lemma A.2.
$\delta^{\top}\ddot{\Sigma}\delta$ $\geq\delta^{\top}\ddot{\Upsilon}\delta-\left|\delta^{\top}(\ddot{\Sigma}-\ddot{\Upsilon})\delta\right|$ (A.28)
Notice that $\widehat{\Upsilon}$ is the Gram matrix of the LUR processes $\zeta_{t}$ with normally distributed errors. The procedures as in Part (a) bounds the first term on the right-hand side of the above expression
$\delta^{\top}\ddot{\Upsilon}\delta\geq\dfrac{c_{\kappa}^{\prime}}{L^{2}s\log p}n\|\delta\|_{2}^{2}$ (A.29)
w.p.a.1 for some absolute constant $c_{\kappa}^{\prime}$. We move on to the second term
$\left|\delta^{\top}(\text{$\widehat{\Sigma}$}-\widehat{\Upsilon})\delta\right|$ $\leq\|\delta\|_{1}^{2}\|\ddot{\Sigma}-\ddot{\Upsilon}\|_{\infty}\leq(1+L)^{2}s\|\delta\|_{2}^{2}\|\ddot{\Sigma}-\ddot{\Upsilon}\|_{\infty}$
$\leq 4L^{2}s\cdot\|\delta\|_{2}^{2}\|\ddot{\Sigma}-\ddot{\Upsilon}\|_{\infty}$ (A.30)
whenever $L\geq 1$. Since $X_{t}=\sum_{s=0}^{t}e_{s}=\Phi_{e}\sum_{s=0}^{t}\varepsilon_{s}=\Phi_{e}\xi_{t-1}$, it follows that
$\|\ddot{\Sigma}-\ddot{\Upsilon}\|_{\text{$\infty$}}$ $\leq$ $C_{L}^{2}\left(\|n^{-1}\sum_{t=1}^{n}(x_{t-1}x_{t-1}^{\top}-\zeta_{t-1}\zeta_{t-1}^{\top})\|_{\infty}+\|\bar{x}\bar{x}^{\top}-\bar{\xi}\bar{\xi}^{\top}\|_{\infty}\right).$
Following the proof of Part (b) in Proposition B.4 of MS24, we can show that under Lemma A.2,
$\|\ddot{\Sigma}-\ddot{\Upsilon}\|_{\text{$\infty$}}=O_{p}\left(n^{3/4+\nu^{\prime}}\sqrt{\log p}\right)$
for any arbitrary small absolute value $\nu^{\prime}$. Inserting the above expression into (A.30), we have
$\frac{|\delta^{\top}(\ddot{\Sigma}-\ddot{\Upsilon})\delta|}{n\|\delta\|_{2}^{2}}\leq 4L^{2}s\cdot O_{p}\left(n^{-1/4+\nu^{\prime}}\sqrt{\log p}\right)=o_{p}\left(\frac{L^{-2}}{s\log p}\right)$ (A.31)
given the condition $s^{2}L^{4}(\log p)^{3/2}=o(n^{1/4-\nu^{\prime}})$ implied by Assumption 4. (A.29) and (A.31) then provide
$\frac{\delta^{\top}\ddot{\Sigma}\delta}{n\|\delta\|_{2}^{2}}$ $\geq\dfrac{c_{\kappa}^{\prime}}{L^{2}s\log p}-o_{p}\left(\frac{L^{-2}}{s\log p}\right)\geq\dfrac{c_{\kappa}}{L^{2}s\log p}$
w.p.a.1 when $n$ is large enough, where $c_{\kappa}=0.5c_{\kappa}^{\prime}.$
The following Lemma establishes the RE condition for LUR regressors without standardization nor demeaning.
Lemma A.5.
Under Assumptions 1-4, w.p.a.1
$\kappa_{{\bf I}}(\widehat{\Sigma}^{(x)},3,s)\geq\dfrac{c_{\kappa}^{(x)}n}{s\cdot\log p}.$ (A.32)
Proof of Lemma A.5.
This lemma is a direct corollary of Lemma A.4 by taking $L=3$ and $c_{\kappa}^{(x)}=c_{\kappa}/9$, and the fact that $\kappa_{{\bf I}}(\widehat{\Sigma}^{(x)},3,s)\geq\kappa_{{\bf I}}(\ddot{\Sigma}^{(x)},3,s)$. ∎
The following Lemma establishes the RE condition for stationary regressors without standardization.
Lemma A.6.
Under Assumptions 1-4,
$\kappa_{{\bf I}}(\widehat{\Sigma}^{(z)},3,s)\geq c_{\kappa}^{(z)}$ (A.33)
w.p.a.1 for some absolute constant $c_{\kappa}^{(z)}$.
Proof of Lemma A.6.
The proof follows standard arguments using concentration inequalities for weakly dependent time series, like (B.30) in MS24. ∎
Define $D^{(x)}={\rm diag}(\widehat{\sigma}_{j})_{j\in\mathcal{M}_{x}}$ as the diagonal matrix that stores the standard deviations of the LUR regressors. In the following Lemma A.7, we establish a lower bound of RE for standardized LURs.
Lemma A.7.
Under Assumptions 1-4, there exists an absolute constant $c_{\kappa}^{(1)}$ such that
$\kappa_{D^{(x)}}(\widehat{\Sigma}^{(x)},3,s)\geq\frac{c_{\kappa}^{(1)}}{s(\log p)^{6}}$ (A.34)
w.p.a.1 as $n\to\infty.$
Proof of Lemma A.7.
Define $\widehat{\sigma}_{\max}^{(x)}=\max_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}$, $\widehat{\sigma}_{\min}^{(x)}=\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}$, and $\widehat{\varsigma}^{(x)}=\widehat{\sigma}_{\max}^{(x)}/\widehat{\sigma}_{\min}^{(x)}$. Further define $\widetilde{\delta}^{(x)}:=(D^{(x)})^{-1}\delta=(\widehat{\sigma}_{j}^{-1}\delta_{j}){}_{j\in\mathcal{M}_{x}}$. Note that $\widehat{\sigma}_{\min}\|\widetilde{\delta}_{\mathcal{M}^{c}}\|_{1}\leq\|\delta_{\mathcal{M}^{c}}\|_{1}$ and $\|\delta_{\mathcal{M}}\|_{1}\leq\widehat{\sigma}_{\max}\|\widetilde{\delta}_{\mathcal{M}}\|_{1}$. Therefore, whenever $\delta\in\mathcal{R}(3,s)$ such that $\|\delta_{\mathcal{M}^{c}}\|_{1}\leq 3\|\delta_{\mathcal{M}}\|_{1}$ for any $|\mathcal{M}|\leq s$, we have $\|\widetilde{\delta}_{\mathcal{M}^{c}}\|_{1}\leq\widehat{\varsigma}^{(x)}\|\widetilde{\delta}_{\mathcal{M}}\|_{1}$ and $\widetilde{\delta}\in\mathcal{R}\left(3\widehat{\varsigma}^{(x)},s\right)$. Then
$\kappa_{D^{(x)}}(\widehat{\Sigma}^{(x)},3,s)$ $=\inf_{\delta\in\mathcal{R}(3,s)}\dfrac{\delta^{\top}(D^{(x)})^{-1}\widehat{\Sigma}^{(x)}(D^{(x)})^{-1}\delta}{\delta^{\top}\delta}=\inf_{\delta\in\mathcal{R}\left(3,s\right)}\dfrac{\widetilde{\delta}^{\top}\widehat{\Sigma}^{(x)}\widetilde{\delta}}{\widetilde{\delta}^{\top}(D^{(x)})^{2}\widetilde{\delta}}$
$\geq\inf_{\widetilde{\delta}\in\mathcal{R}\left(3\widehat{\varsigma}^{(x)},s\right)}\dfrac{\widetilde{\delta}^{\top}\widehat{\Sigma}^{(x)}\widetilde{\delta}}{\widetilde{\delta}^{\top}(D^{(x)})^{2}\widetilde{\delta}}\geq(\widehat{\sigma}_{{\rm\max}}^{(x)})^{-2}\inf_{\widetilde{\delta}\in\mathcal{R}\left(3\widehat{\varsigma}^{(x)},s\right)}\dfrac{\widetilde{\delta}^{\top}\widehat{\Sigma}^{(x)}\widetilde{\delta}}{\widetilde{\delta}^{\top}\widetilde{\delta}}=\dfrac{\kappa(\widehat{\Sigma}^{(x)},3\widehat{\varsigma}^{(x)},s)}{(\widehat{\sigma}_{{\rm\max}}^{(x)})^{2}}.$
Taking $L=3\widehat{\varsigma}^{(x)}$. By Lemma A.4 we have $\kappa(\widehat{\Sigma}^{(x)},3\widehat{\varsigma}^{(x)},s)\geq\dfrac{cn}{9s\log p(\widehat{\sigma}_{{\rm\max}}^{(x)})^{2}\cdot(\widehat{\varsigma}^{(x)})^{2}}$ w.p.a.1 for some absolute constant $c$. By (A.36), there exists some absolute constant $c^{\prime}$ such that
[…]
Therefore,
$\kappa(\widehat{\Sigma}^{(x)},3\widehat{\varsigma}^{(x)},s)\geq\dfrac{cn}{9s(\log p)^{4}(c^{\prime})^{2}}$
w.p.a.1. Then Lemma A.7 follows with $c_{\kappa}^{(x)}=c/(3c^{\prime})^{2}.$
The following lemma delivers the bounds of the standard deviations used for scaling in Slasso.
Lemma A.8.
Under Assumptions 1-4,
  • (a)
    For stationary regressors, there exists some absolute constants $\sigma_{\min}<\sigma_{\max}$, such that with probability approaching one,
    $\sigma_{\min}\leq\min_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}\leq\max_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}\leq\sigma_{\max}.$ (A.35)
  • (b)
    For nonstationary regressors,
    $\sqrt{\dfrac{n}{\log p}}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}\leq\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}.$ (A.36)
Proof of Lemma A.8.
For Part (a), the proof follows (B.60) and (B.61) in the appendix of MS24.
For Part (b), the lower bound follows by
$\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\geq\kappa_{{\bf I}}(\ddot{\Sigma}^{(1)},3,1)\geq\dfrac{c_{\kappa}^{(1)}}{\log p}$
where the second inequality applies Lemma A.4. For the upper bound, the LUR regressor $x_{j,t}=\sum_{s=0}^{t}\rho_{j}^{*t-s}e_{j,s}$ is a partial sum of a sub-exponential and $\alpha$-mixing sequence. Therefore,
$\max_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}^{2}\leq\max_{j\in\mathcal{M}_{x}}n^{-1}\sum_{t=1}^{n}x_{j,t-1}^{2}\leq\max_{j\in\mathcal{M}_{x},t\in[n]}|x_{j,t}|^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\log p,$
where the last inequality applies (A.25). We complete the proof of Lemma A.8. ∎
The following lemma gives the DB condition. Compared to RE, the DB condition for mixed roots is more straightforward and not substantially distinguished from that in MS.
Lemma A.9.
Under Assumptions 1-4, we have
$\|n^{-1}\sum_{t=1}^{n}Z_{t-1}u_{t}\|_{\infty}+\dfrac{1}{\sqrt{n}}\|n^{-1}\sum_{t=1}^{n}X_{t-1}u_{t}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{1+\frac{1}{2r}}}{\sqrt{n}},$ (A.37)
$\dfrac{1}{\sqrt{n}}\|n^{-1}\sum_{t=1}^{n}X_{t-1}e_{t-1}^{\top}\|_{\infty}+\dfrac{1}{\sqrt{n}}\|n^{-1}\sum_{t=1}^{n}X_{t-1}Z_{t-1}^{\top}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{1+\frac{1}{2r}}}{\sqrt{n}}$ (A.38)
as $n\to\infty.$
Proof of Lemma A.9.
$\|n^{-1}\sum_{t=1}^{n}Z_{t-1}u_{t}\|_{\infty}$ can be bounded following the proofs in (B.29) of MS24. $\|n^{-1}\sum_{t=1}^{n}X_{t-1}u_{t}\|_{\infty}$, $\|n^{-1}\sum_{t=1}^{n}X_{t-1}Z_{t-1}^{\top}\|_{\infty}$, and $\|n^{-1}\sum_{t=1}^{n}X_{t-1}e_{t-1}^{\top}\|_{\infty}$ can be bounded following exactly the same procedures in the proof of MS24’s Proposition B.2 about deviation bound (DB) for unit root. Take $\|n^{-1}\sum_{t=1}^{n}X_{t-1}e_{t-1}^{\top}\|_{\infty}$ as an example. The essential modification is the expression of $T_{2}$ above MS24’s Equation (B.12). It should be changed into
$T_{2}$ $=\sup_{k\in\mathcal{M}_{x},j\in[p],t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{j}^{{}^{*}t-1-r}e_{j,r}\right|,$
which can be bounded following the same procedures in MS24. ∎
A.1.3 Proofs of Lemma 2
Proof of Lemma 1.
Proof of (3.7). We have for any $\delta$
$\delta^{\top}D^{-1}\widehat{\Sigma}D{}^{-1}\delta$ $\geq\delta^{\top}D^{-1}\widehat{\Delta}D^{-1}\delta-\|D^{-1}(\widehat{\Sigma}-\widehat{\Delta})D^{-1}\|_{\infty}\|\delta\|_{1}^{2}.$
Lemmas A.7 and (A.33) suggest that for any $\delta\in\mathcal{R}(3,s)$
$\dfrac{\delta^{\top}D^{-1}\widehat{\Delta}D^{-1}\delta}{\|\delta\|_{2}^{2}}\geq\dfrac{c}{s(\log p)^{4}}$
for some absolute constant $c$. By (A.24) and Lemma A.3,
$\dfrac{\|D^{-1}(\widehat{\Sigma}-\widehat{\Delta})D^{-1}\|_{\infty}\|\delta\|_{1}^{2}}{\|\delta\|_{2}^{2}}$ $=O_{p}\left(\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\right)\cdot\dfrac{\|\delta\|_{1}^{2}}{\|\delta\|_{2}^{2}}$
$=O_{p}\left(\dfrac{s(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\right).$
Therefore,
$\dfrac{\delta^{\top}D^{-1}\widehat{\Sigma}D{}^{-1}\delta}{\|\delta\|_{2}^{2}}\geq\dfrac{c}{s(\log p)^{4}}+O_{p}\left(\dfrac{s(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\right)\geq\dfrac{0.5c}{s(\log p)^{4}}$
when $n$ is sufficiently large, where the second inequality applies the fact that $s(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}\gg 1/[s(\log p)^{4}]$, which is implied by Assumption 4. Then (3.7) follows with $c_{\kappa}=0.5c.$
Proof of (3.8). The DB condition follows by
$n^{-1}\|\sum_{t=1}^{n}D^{-1}W_{t-1}u_{t}\|_{\infty}$ $=\max_{j\in[p]}n^{-1}\left|\sum_{t=1}^{n}\frac{w_{j,t-1}}{\widehat{\sigma}_{j}}u_{t}\right|$
$\leq\frac{1}{\min_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\|n^{-1}\sum_{t=1}^{n}Z_{t-1}u_{t}\|_{\infty}+\dfrac{\sqrt{n}}{\min_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\dfrac{1}{\sqrt{n}}\|n^{-1}\sum_{t=1}^{n}X_{t-1}u_{t}\|_{\infty}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(1+\sqrt{\log p})\dfrac{(\log p)^{1+\frac{1}{2r}}}{\sqrt{n}}$
$=\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}},$
where the third row applies Lemmas A.9 and A.8. ∎
Proof of Lemma 2.
By Lemma 1 of MS, our Lemma 1 implies that
$\|D(\widehat{\beta}^{{\rm S}}-\beta^{*})\|_{1}=O_{p}\left(\dfrac{s(\log p)^{\frac{3}{2}+\frac{1}{2r}}/\sqrt{n}}{1/s(\log p)^{4}}\right)=O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n}}\right).$ (A.39)
as $n\to\infty$, which verifies Lemma 2. ∎
A.2 Proofs for Section 3.2
A.2.1 Local Unit Root Regressors
We first introduce and prove several technical lemmas.
Lemma A.10.
Under Assumptions 14, for any fixed $j\in\mathcal{M}_{x}$
$\sup_{t\in[n]}|\zeta_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.40)
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=0}^{n-1}e_{k,t}\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}},$ (A.41)
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$ (A.42)
Furthermore, recall that $\widehat{\varsigma}_{j}=\sqrt{n^{-1}\sum_{t=1}^{n}(\zeta_{j,t-1}-\bar{\zeta}_{j})^{2}}.$ Then for any $j\in\mathcal{M}_{x}$,
$\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}.$ (A.43)
Proof of Lemma A.10.
We work on these inequalities one by one.
Proof of (A.40). Recall from (3.5) that $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$. By (13) in phillips2009econometric, when $j\in\mathcal{M}_{x}$ we have
$\zeta{}_{j,t}=\zeta_{j,t}^{0}+\frac{C}{n}\psi_{j,t}^{0}\text{ for }t\geq 1,$ (A.44)
where
$\zeta_{j,t}^{0}=\sum_{s=1}^{t}\rho_{\zeta}^{s}e_{j,t-s}\text{ and }\psi_{j,t}^{0}=\sum_{s=0}^{t-1}\rho_{\zeta}^{s}x_{j,t-s-1}$
is a partial sum of $\alpha$-mixing sup-exponential components $e_{j,t-s}$ weighted by $\rho_{\zeta}^{s}.$
We first bound $\zeta_{j,t}^{0}$. Define
$a_{n}:=\lfloor n^{\tau}(\log p)^{2}\rfloor.$
Note that $\rho_{\zeta}^{s}e_{j,t-s}$ is sub-exponential with an exponentially decaying $\alpha$-mixing coefficient, and thus $\zeta^{0}{}_{j,t}$ is the partial sum of $t$ observations from a sub-exponential and $\alpha$-mixing time series. MS24’s Lemma B.2 yields
$\sup_{t\leq a_{n}}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{a_{n}\cdot\log p}=O\left[n^{\tau/2}(\log p)^{3/2}\right].$ (A.45)
In addition, when $t>a_{n}$,
$|\zeta_{j,t}^{0}|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\left|\sum_{a_{n}<s\leq t}\rho_{\zeta}^{s}e_{j,t-s}\right|\leq\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|+\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|.$ (A.46)
By the same arguments for (A.45), we bound the two sums on the right-hand side of (A.46) by
$\sup_{a_{n}<t\leq n}\left|\sum_{s\leq a_{n}}\rho_{\zeta}^{s}e_{j,t-s}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.47)
and
$\sup_{a_{n}<t\leq n}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{(n-a_{n})\cdot\log p}\text{.}$
Besides, under the assumption $p\geq n^{\nu_{1}},$ the sequence
$\rho_{\zeta}^{a_{n}}=\left(1-C_{\zeta}/n^{\tau}\right)^{\lfloor n^{\tau}(\log p)^{2}\rfloor}=O\left(\exp\left(-C_{\zeta}(\log p)^{2}\right)\right)=O\left(p^{-C_{\zeta}\log p}\right)$
converges to zero faster than the reciprocal of any polynomial function of $n.$ Thus,
$\sup_{a_{n}<t\leq n}\rho_{\zeta}^{a_{n}}\left|\sum_{0<s\leq t-a_{n}}\rho_{\zeta}^{s-a_{n}}e_{j,t-s+a_{n}}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\rho_{\zeta}^{a_{n}}\sqrt{(n-a_{n})\cdot\log p}=o\left(n^{\tau/2}(\log p)^{3/2}\right).$ (A.48)
By (A.46), (A.47), and (A.48), it follows that
$\sup_{a_{n}<t\leq n}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}.$ (A.49)
By (A.45) and (A.49), we have
$\sup_{1<t\leq n}|\zeta_{j,t}^{0}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}.$ (A.50)
We then bound $\psi_{j,t}^{0}$. By (A.25),
$\sup_{t\in[n]}\left|\psi_{j,t}^{0}\right|$ $\leq\sum_{s=0}^{n}\rho_{\zeta}^{s}\sup_{t\in[n]}\left|x_{j,t-s-1}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}\cdot\sum_{s=0}^{n}\rho_{\zeta}^{s}$
$\leq\sqrt{n\log p}\cdot\frac{1}{1-\rho_{\zeta}}=O(n^{\frac{1}{2}+\tau}\sqrt{\log p}).$ (A.51)
Then by (A.44), (A.50) and (A.51),
$\sup_{t\in[n]}\left|\zeta_{j,t}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2}+n^{\tau-\frac{1}{2}}\sqrt{\log p}=O(n^{\tau/2}(\log p)^{3/2}),$
where the second step applies the fact that $\tau\in(0,1)$, which implies $\tau-\frac{1}{2}<\frac{\tau}{2}$. Then (A.40) follows.
Proof of (A.41). Using the result (A.40), the proof of (A.41) follows exactly the same procedures in the proof of MS24’s Proposition B.2 about deviation bound (DB) for unit root. Essential modifications include:
  1. 1.
    Change the bound $\sup_{t}|x_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n\log p}$ in MS24’s Equation (B.12) into
    $\sup_{t}|\zeta{}_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$
    for a fixed $j$, which has been established in (A.40). With this result, we can deduce the same upper bound of $T_{1}$ in MS24’s Equation (B.13).
  2. 2.
    Change the expression of $T_{2}$ above MS24’s Equation (B.12) into
    $T_{2}$ $=\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}\Delta x_{j,r}\right|$
    $\leq\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|+\dfrac{|c_{j}|}{n}\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}x_{j,r-1}\right|,$
    where the second step follows the DGP of LURs $x_{j,t}=(1-c_{j}^{*}/n)x_{j,t-1}+e_{j,t}$. Following the same way as MS24 to bound $T_{2}$, we can bound the first term by
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}e_{j,r}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\log p.$ (A.52)
    In addition, it is easy to show that
    $\sup_{k\in\mathcal{M}_{x},t\in[n]}|e_{k,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p$ (A.53)
    given that $e_{k,t}$ is sub-exponential. Therefore, the second term is bounded by
    $n^{-1}\sup_{k\in\mathcal{M}_{x},t\in[n]}\left|\sum_{t=G+1}^{n}e_{k,t}\sum_{r=t-G+1}^{t-1}\rho_{\zeta}^{t-1-r}x_{j,r-1}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}n^{-1}\sum_{t=G+1}^{n}\left|e_{k,t}\right|\sup_{k\in\mathcal{M}_{x},t\in[n]}\sum_{r=t-G+1}^{t-1}|x_{j,r-1}|$
    $\leq G\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}|e_{k,t}|\cdot\sup_{t\in[n]}|x_{j,t}|$ (A.54)
    $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n}(\log p)^{3/2},$ (A.55)
    where the last step follows (A.25) and (A.53). By (A.52) and (A.55), for our case we also have
    $T_{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\log p.$
  3. 3.
    Change the definition of the event $\mathcal{X}_{t}$ below MS24’s Equation (B.16) into
    $\mathcal{X}_{t}=\{|\zeta{}_{j,t}|\leq C_{X}\sqrt{n\log p}\}$
    for some absolute constant $C_{X}$. Equation (A.40) in the current paper implies that $\Pr\{\bigcup_{t=1}^{n}\mathcal{X}_{t}^{c}\}\to 0$ as $n\to\infty$. Therefore, the arguments below MS24’s Equation (B.24) are still valid. We can thus show that the upper bound of $T_{3}$ defined above into MS24’s Equation (B.12) still holds.
Proof of (A.42). By the decomposition (A.44), we have
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\leq\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|+\dfrac{C_{\zeta}}{n}\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\psi_{j,t-1}^{0}\right|.$ (A.56)
We first bound $\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|.$ Note that when $j\in\mathcal{M}_{x}$, we have the following AR(1) representation
$\zeta_{j,t}^{0}=\rho_{\zeta}\zeta_{j,t-1}^{0}+e_{j,t}.$ (A.57)
Thus, for any $k\in\mathcal{M}_{x},$
$x_{k,t-1}\zeta_{j,t}^{0}=\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}^{0}+x_{k,t-1}e_{j,t}$
where $\rho_{\zeta}$ is defined in (3.5). By $x_{k,t-1}\zeta_{j,t}^{0}=(\rho_{k}^{*})^{-1}(x_{k,t}-e_{k,t})\zeta_{j,t}^{0},$ we obtain
$x_{k,t}\zeta_{j,t}^{0}=\rho_{k}^{*}\rho_{\zeta}x_{k,t-1}\zeta_{j,t-1}^{0}+\rho_{k}^{*}x_{k,t-1}e_{j,t}+e_{k,t}\zeta_{j,t}^{0}.$ (A.58)
Summing up both sides of (A.58) and using the fact that $\sum_{t=1}^{n}x_{k,t}\zeta_{j,t}^{0}=\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}-x_{k,0}\zeta_{j,0}+x_{k,n}\zeta_{j,n}^{0}$, we deduce
$\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}-x_{k,0}\zeta_{j,0}^{0}+x_{k,n}\zeta_{k,n}^{0}=\rho_{k}^{*}\rho_{\zeta}\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}+\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}.$
It can be further arranged into
$(1-\rho_{k}^{*}\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}=\left(x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right)+\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}+\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}.$ (A.59)
By (A.25) and (A.40),
$\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}.$ (A.60)
By (A.38) and (A.41),
$\sup_{k\in\mathcal{M}_{x}}\left|\rho_{k}^{*}\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n(\log p)^{1+\frac{1}{2r}}.$ (A.61)
By (A.59), (A.60), and (A.61),
$\sup_{k\in\mathcal{M}_{x}}\left|(1-\rho_{k}^{*}\rho_{\zeta})\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|$ $\leq\sup_{k\in\mathcal{M}_{x}}\left|x_{k,0}\zeta_{j,0}^{0}-x_{k,n}\zeta_{j,n}^{0}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}e_{j,t}\right|+\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}e_{k,t}\zeta_{j,t}^{0}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\frac{1+\tau}{2}}(\log p)^{2}+n(\log p)^{1+\frac{1}{2r}}$
$\leq 2n(\log p)^{1+\frac{1}{2r}}.$
Since $\sup_{k\in\mathcal{M}_{x}}\left|\frac{1}{1-\rho_{k}^{*}\rho_{\zeta}}\right|=O(n^{\tau})$, we have
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}^{0}\right|\preccurlyeq n^{1+\tau}(\log p)^{1+\frac{1}{2r}}.$ (A.62)
In addition,
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\psi_{j,t-1}^{0}\right|$ $\leq n\sup_{k\in\mathcal{M}_{x},t\in[n]}|x_{k,t-1}|\cdot\sup_{k\in\mathcal{M}_{x},t\in[n]}|\psi_{j,t-1}^{0}|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n\cdot\sqrt{n\log p}\cdot n^{\frac{1}{2}+\tau}\sqrt{\log p}=n^{2+\tau}\log p,$ (A.63)
where the second step applies (A.25) and (A.51). By (A.56), (A.62) and (A.63),
$\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}x_{k,t-1}\zeta_{j,t-1}\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{1+\tau}(\log p)^{1+\frac{1}{2r}}+\frac{n^{2+\tau}\log p}{n}=O(n^{2+\tau}\log p).$
We complete the proof of (A.42).
Proof of (A.43). We have the following decomposition
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}.$
By the law of large number (LLN) in phillips2009econometric, we have
$\dfrac{\sum_{t=1}^{n}(\zeta_{j,t-1})^{2}}{n^{1+\tau}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}},$ (A.64)
where ${\rm lvar}(e_{j,t})$ is the long-run variance of $e_{j,t}$.
We then bound $\sum_{t=1}^{n}\zeta_{j,t-1}.$ By (A.44), we have
$\sum_{t=1}^{n}\zeta_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}^{0}+\frac{C_{\zeta}}{n}\sum_{t=1}^{n}\psi_{j,t-1}^{0}.$ (A.65)
We first bound $\sum_{t=1}^{n}\zeta_{j,t-1}^{0}$. Without loss of generality, assume $\zeta_{j,0}^{0}=0$. Summing up both sides of (A.57), we have
$\sum_{t=1}^{n}e_{j,t}=\sum_{t=1}^{n}\zeta_{j,t}^{0}-\rho_{\zeta}\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=\zeta_{j,n}^{0}+(1-\rho_{\zeta})\sum_{t=1}^{n}\zeta_{j,t-1}^{0}.$
Since $1-\rho_{\zeta}=C_{\zeta}n^{-\tau},$we have $\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=n^{\tau}C_{\zeta}^{-1}(\sum_{t=1}^{n}e_{j,t}-\zeta_{j,n}^{0}).$ Note that $\sum_{t=1}^{n}e_{j,t}$ is a unit root, and thus $\sum_{t=1}^{n}e_{j,t}=O_{p}(\sqrt{n\log p})$ by (A.25). Also, $|\zeta_{j,n}^{0}|=O_{p}(n^{\tau/2}(\log p)^{3/2})$ by (A.40). Therefore,
$\sum_{t=1}^{n}\zeta_{j,t-1}^{0}=O_{p}\left[n^{\tau}\cdot\left(\sqrt{n\log p}+n^{\tau/2}(\log p)^{3/2}\right)\right]=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right).$ (A.66)
In addition, by (A.51)
$\sum_{t=1}^{n}\psi_{j,t-1}^{0}\leq n\cdot\sup_{t\in[n]}|\psi_{j,t-1}^{0}|=O_{p}(n^{3/2+\tau}\sqrt{\log p}).$ (A.67)
By (A.65), (A.66), and (A.67) we have
$\sum_{t=1}^{n}\zeta_{j,t-1}=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}+\frac{n^{3/2+\tau}\sqrt{\log p}}{n}\right)=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right),$ (A.68)
and thus
$\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}=O_{p}\left(\dfrac{\log p}{n^{1-\tau}}\right)\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.69)
By (A.64) and (A.69),
$\dfrac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}=\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{n^{1+\tau}}-\dfrac{1}{n^{\tau}}\left(\dfrac{\sum_{t=1}^{n}\zeta_{j,t-1}}{n}\right)^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}}.$ (A.70)
This completes the proof of Lemma A.10. ∎
With these preparatory lemmas, we will prove Proposition 1 for $j\in\mathcal{M}_{x}$.
Proof of Proposition 1 for $j\in\mathcal{M}_{x}$..
For simplicity of exposition, define $\widetilde{\zeta}_{j}=(\widetilde{\zeta}_{j,1},\dots,\widetilde{\zeta}_{j,n-1})^{\top}$ and recall that $W_{-j,\cdot}=(W_{-j,1},\dots,W_{-j,n-1})^{\top}.$ By the definition of Slasso, we have
$\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j,\cdot}\widehat{\varphi}^{(j)}\|_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{1}{n}\|\widetilde{\zeta}_{j}-W_{-j}\varphi\|_{2}^{2}+\mu\|D_{-j}\varphi\|_{1}.$
for an arbitrary (p1)(p-1)-dimensional vector $\varphi$. We can write the above inequality into
$n^{-1}\lVert W_{-j,\cdot}\left(\widehat{\varphi}^{(j)}-\varphi\right)\rVert_{2}^{2}+\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq$ $\dfrac{2}{n}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\mu\|D_{-j}\varphi\|_{1}.$ (A.71)
Write the coefficient vector as $\varphi=(\varphi_{k})_{k\in[p],k\neq j}$, where $\varphi_{k}$ is the coefficient of $w_{k,t}$. Define $\varphi_{\mathcal{M}_{x}}=(\varphi_{k})_{k\in\mathcal{M}_{x},k\neq j}$ and $\varphi_{\mathcal{M}_{z}}=(\varphi_{k})_{k\in\mathcal{M}_{z}}.$ Take a vector $\varphi$ such that
$\|\varphi_{\mathcal{M}_{x}}\|_{1}$ $=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+\tau\wedge(1-\tau)}}}\text{ and }\|\varphi_{\mathcal{M}_{z}}\|_{1}=\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.72)
for a positive number $C_{\varphi}=O(1)$. Without loss of generality, suppose that $s\geq 1$. We will show that
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\leq\frac{\mu}{2}\left(1-\frac{1}{2s^{2}}\right),$ (A.73)
Under (A.73), we can deduce by (A.71) that
$\mu\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi)\|_{1}+\|D_{-j}\varphi\|_{1}\right]$
$\leq\mu\left[\left(1-\frac{1}{2s^{2}}\right)\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}+2\|D_{-j}\varphi\|_{1}\right],$
which implies $\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq 4s^{2}\|D_{-j}\varphi\|_{1}.$ In addition, by (A.35) and (A.36) in Lemma A.8, there exists a absolute constant $C_{\sigma}$ such that
$\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\leq C_{\sigma}\sqrt{n\log p}\text{ and }\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\leq C_{\sigma}$ (A.74)
w.p.a.1. Therefore, by (A.72) and (A.74),
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$ $\leq 4s^{2}\|D_{-j}\varphi\|_{1}$
$\leq 4s^{2}\left(\max_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}\right)$
$\leq\dfrac{4C_{\varphi}\cdot C_{\sigma}s^{2}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}$ (A.75)
w.p.a.1, which implies (3.15).
It is thus sufficient to prove (A.73). Note that
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}\leq(n\widehat{\varsigma}_{j})^{-1}\left(\sup_{k\in\mathcal{M}_{x}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|+\sup_{k\in\mathcal{M}_{z}}\left|\sum_{t=1}^{n}\widehat{\sigma}_{k}^{-1}w_{k,t-1}\zeta_{j,t-1}\right|\right).$
Repeatedly using the bounds in (A.35) and (A.36) in Lemma A.8, and (A.41), (A.42), and (A.43) in Lemma A.10,
$\|n^{-1}D_{-j}^{-1}W_{-j,\cdot}^{\top}\widetilde{\zeta}_{j}\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(n^{1+\tau/2})^{-1}\left(\dfrac{n^{1+\tau}(\log p)^{1+\frac{1}{2r}}}{\sqrt{n/\log p}}+n(\log p)^{1+\frac{1}{2r}}\right)$
$\leq\ 2(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}.$ (A.76)
In addition,
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$
$\leq$ $\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{x}}\|_{1}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\|\varphi_{\mathcal{M}_{z}}\|_{1}$
$\leq$ $\ \max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\cdot\dfrac{C_{\varphi}(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}},$ (A.77)
where the second inequality applies (A.72). By the sup-exponential distributions of the stationary components, we can deduce that
$\sup_{k\in\mathcal{M}_{z},t\in[n]}|w_{k,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p.$ (A.78)
Therefore,
$\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$ $\leq\max_{j\in[p],k\in\mathcal{M}_{x}}\max_{t\in[n]}\left|\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|z_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|x_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{x},t\in[n]}\left|w_{k,t}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p\cdot\sqrt{n\log p}=\sqrt{n(\log p)^{3}},$ (A.79)
where the last row applies (A.25), (A.35), (A.36), and (A.78). Following similar arguments we can deduce
$\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$ $\leq\max_{j\in[p],k\in\mathcal{M}_{z}}\max_{t\in[n]}\left|\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|$
$\leq\left(\dfrac{\max_{j\in\mathcal{M}_{z},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}\vee\dfrac{\max_{j\in\mathcal{M}_{x},t\in[n]}\left|w_{j,t}\right|}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}\right)\cdot\max_{k\in\mathcal{M}_{z},t\in[n]}\left|w_{k,t}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\log p\cdot\log p=(\log p)^{2}.$ (A.80)
Thus, by (A.77), (A.79), and (A.80),
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}W_{-j,\cdot}\varphi\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n(\log p)^{3}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{1+(1-\tau)\wedge\tau}}}+(\log p)^{2}\cdot\dfrac{(\log p)^{\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{2(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$ (A.81)
(A.76) and (A.81) yield
$n^{-1}\|D_{-j}^{-1}W_{-j,\cdot}^{\top}(\widetilde{\zeta}_{j}-W_{-j,\cdot}\varphi)\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}.$
Therefore, (A.73) holds as $\mu=C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}/\sqrt{n^{(1-\tau)\wedge\tau}}$ with a sufficiently large $C_{{\rm a}}$. This completes the proof of Proposition 1 for $j\in\mathcal{M}_{x}$. ∎
A.2.2 Stationary Regressor
Recall $\rho_{\zeta}=1-C_{\zeta}/n^{\tau}$ as defined in (3.5), and by (23) in phillips2009econometric we have the following decomposition
$\zeta_{j,t}=w_{j,t}-\dfrac{C_{\zeta}}{n^{\tau}}\phi_{j,t},\ \phi_{j,t}:=\sum_{s=1}^{t}\rho_{\zeta}^{t-s}w_{j,s-1}.$ (A.82)
In addition, we define
$\eta_{j,t}=\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$ (A.83)
where $\varphi_{0z}^{(j)*}$ is defined in (3.12). Compare $\eta_{j,t}$ to its standardized version $\widetilde{\eta}_{j,t}$ defined in (3.14), we have
$\eta_{j,t}=\widehat{\varsigma}_{j}\widetilde{\eta}_{j,t}.$ (A.84)
Finally, define the standardized regressors
$\widetilde{W}_{-j,t}=D_{-j}^{-1}W_{-j,t},$ (A.85)
and $\widetilde{W}_{-j,\cdot}=(W_{-j,0},W_{-j,1},\dots,W_{-j,n-1})^{\top}.$
Lemma A.11.
Under Assumptions 15, for $j\in\mathcal{M}_{z}$
$\sup_{t}|\phi_{j,t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau/2}(\log p)^{3/2},$ (A.86)
$\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1,$ (A.87)
and
$\lVert n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\widetilde{\eta}_{j,t}\rVert_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{-\tau/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}},$ (A.88)
with $\widetilde{\eta}_{j,t}$ defined in (3.14), and $\widetilde{W}_{-j,t}$ defined in (A.85).
Proof of Lemma A.11.
By definition of $\phi_{j,t}$ in (A.82), we can easily deduce the following recursive formula for $\phi_{j,t}$ that
$\phi_{j,t}=\rho_{\zeta}\phi_{j,t-1}+w_{j,t-1}$
Note that $\phi_{j,t-1}\in\sigma(w_{j,0},\dots,w_{j,t-2}).$ Then $\phi_{j,t}$ is an AR(1) process with coefficient $1-C_{\zeta}/n^{\tau}$ and innovation $w_{j,t-1}$. Recall that $w_{j,t}$ is $\alpha$-mixing and sub-exponential; we obtain (A.86) following the same arguments for (A.40).
In addition, following the same arguments for (A.43), we have
$n^{-1}\sum_{t=1}^{n}\phi_{j,t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}.$
By (A.41), the cross-product between a mildly integrated $\phi_{j,t}$ and a stationary $w_{j,t}$ is bounded by
$\sum_{t=1}^{n}w_{j,t}\phi_{j,t}=o_{p}(n^{1+\tau}).$
Thus by (A.82),
$n^{-1}\sum_{t=1}^{n}\zeta_{j,t}^{2}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}+n^{-(1+2\tau)}\sum_{t=1}^{n}\phi_{j,t}^{2}+n^{-(1+\tau)}\sum_{t=1}^{n}w_{j,t}\phi_{j,t}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}+o_{p}(1).$ (A.89)
Furthermore, (A.68) implies that the mildly integrated time series $\phi_{j,t}$ satisfies
$\sum_{t=1}^{n}\phi_{j,t}=O_{p}\left(n^{\tau+1/2}\sqrt{\log p}\right)=o_{p}(n^{1+\tau}).$
We thus have by (A.82) that
$n^{-1}\sum_{t=1}^{n}\zeta_{j,t}=n^{-1}\sum_{t=1}^{n}w_{j,t}+o_{p}(1).$ (A.90)
(A.89), (A.90), and a standard law of large number imply
$\widehat{\varsigma}_{j}^{2}=n^{-1}\sum_{t=1}^{n}\zeta_{j,t}^{2}-\left(n^{-1}\sum_{t=1}^{n}\zeta_{j,t}\right)^{2}=n^{-1}\sum_{t=1}^{n}w_{j,t}^{2}-\left(n^{-1}\sum_{t=1}^{n}w_{j,t}\right)^{2}+o_{p}(1)\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}\left(w_{j,t}\right),$ (A.91)
which verifies (A.87).
We then show (A.88). Note that
$\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-(1+\tau)}\sum_{t=1}^{n}Z_{-j,t-1}\phi_{j,t}\|_{\infty}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{-(1+\tau)}\cdot n\sqrt{\log p}\cdot n^{\tau/2}(\log p)^{3/2}$
$=n^{-\tau/2}(\log p)^{2},$
where the first row applies the decomposition (A.82), and the second row applies the bounds (A.78) and (A.86). Also,
$\|n^{-1}\sum_{t=1}^{n}X_{t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-(1+\tau)}\sum_{t=1}^{n}X_{t-1}\phi_{j,t}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}},$
where the second inequality applies the rate in (A.42) of the cross product between a local unit root and a mildly integrated series. Thus,
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\|n^{-1}\sum_{t=1}^{n}X_{t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}(\zeta_{j,t}-w_{j,t})\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}+\dfrac{(\log p)^{2}}{n^{\tau/2}}\leq\dfrac{2(\log p)^{2}}{n^{\tau/2}}$
with $\tau\in(0,1)$ and $n$ large enough, where the second inequality applies by (A.35) and (A.36). Further by (A.43),
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2}}{n^{\tau/2}}.$ (A.92)
In addition, define
$\eta_{j,t}^{(1)}=w_{j,t}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*}.$ (A.93)
By the definition of $\eta_{j,t}$, we have
$\eta_{j,t}^{(1)}=(\zeta_{j,t}-w_{j,t})+\eta_{j,t}.$ (A.94)
note that the time series $\eta_{j,t}^{(1)}$ is stationary and $\mathbb{E}\left[Z_{-j,t-1}\eta_{j,t}^{(1)}\right]=0$. We then have
$\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{\dfrac{\log p}{n}}$
by standard concentration inequalities; e.g. MS24’s Eq. (B.31). Also, following the same way to prove (A.38), we have
$\|n^{-1}\sum_{t=1}^{n}X_{t-1}\eta_{j,t}^{(1)}\|_{\infty}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{1+\frac{1}{2r}}.$
Thus,
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\|n^{-1}\sum_{t=1}^{n}X_{t-1}\eta_{j,t}^{(1)}\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|n^{-1}\sum_{t=1}^{n}Z_{-j,t-1}\eta_{j,t}^{(1)}\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}+\sqrt{\dfrac{\log p}{n}}\leq\dfrac{2(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}.$ (A.95)
Furthermore, by (A.87) and (A.95) we deduce that
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n}}\leq\dfrac{(\log p)^{2}}{n^{\tau/2}}.$ (A.96)
By (A.94) and (A.84), we have the following decomposition
$\widetilde{\eta}_{j,t}=\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}+\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}.$
Then by the triangular inequality, we have
$\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\widetilde{\eta}_{j,t}\|_{\infty}$ $\leq\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\zeta_{j,t}-w_{j,t}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}+\|n^{-1}\sum_{t=1}^{n}\widetilde{W}_{-j,t-1}\left(\frac{\eta_{j,t}^{(1)}}{\widehat{\varsigma}_{j}}\right)\|_{\infty}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{(\log p)^{2}}{n^{\tau/2}}.$
where the second row applies (A.92) and (A.96). We complete the proof of Lemma A.11. ∎
Proof of Proposition 1 for $j\in\mathcal{M}_{z}$..
According to MS24’s Proposition 3(c), the RE is bounded from below by $\frac{c}{s(\log p)^{4}}$ for some absolute constant $c$ w.p.a.1. In addition, by (A.88), the DB is $O_{p}\left(\dfrac{\log p}{n^{\tau/2}}\right)$ when $\mu$ follows the order in Proposition 1. Then by MS24’s Lemma 1
$\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{(j)*})\|_{1}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}s\cdot\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}/(s(\log p)^{4})=\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}.$
This completes the proof of Proposition 1 for $j\in\mathcal{M}_{z}$. ∎
A.3 Proofs for Section 3.3
A.3.1 Technical Lemmas
Lemma A.12.
Suppose Assumptions 1-5 hold. Then for any $j\in[p]$ ,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm sgn}(G_{j}^{*})$ (A.97)
where
$G_{j}^{*}=\begin{cases}\dfrac{1}{C_{\zeta}}\left({\rm lvar}\left(e_{j,t}\right)+\int_{0}^{1}\mathcal{U}_{j}(r)d\mathcal{U}_{j}(r)\right),&j\in\mathcal{M}_{x},\\ {\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right),&j\in\mathcal{M}_{z},\end{cases}$ (A.98)
with ${\rm lvar}\left(e_{j,t}\right)$ being the long-run variance of $e_{j,t}$, $\mathcal{U}_{j}(r)=\int_{0}^{1}{\rm e}^{c_{j}^{*}(r-s)}{\rm d}\mathcal{B}_{j}(s)$ being an OU process, $\mathcal{B}_{j}$ being the Brownian motion of variance ${\rm lvar}\left(e_{j,t}\right)$, and $\eta_{j,t}^{(1)}$ defined in (A.93). In addition,
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}H_{j}$ (A.99)
where
$H_{j}=\begin{cases}1,&j\in\mathcal{M}_{x},\\ \frac{{\rm var}(\eta_{j,t}^{(1)})}{{\rm var}\left(w_{j,t}\right)}&j\in\mathcal{M}_{z}.\end{cases}$
Proof of Lemma A.12.
We first prove (A.97). The first step is to show
$\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.100)
By definition of the Slasso residual,
$\widehat{u}_{t}=y_{t}-W_{t-1}^{\top}\widehat{\theta}^{{\rm S}}=u_{t}+W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}).$
Thus,
$\widehat{\sigma}_{u}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}\widehat{u}_{t}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}+\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}+\dfrac{2}{n}\sum_{t=1}^{n}u_{t}W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}).$ (A.101)
By MS24’s Theorem 3, we have
$\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}=\|n^{-1}W(\theta^{*}-\widehat{\theta}^{{\rm S}})\|_{2}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.102)
By the Cauchy-Schwartz inequality,
$\dfrac{2}{n}\sum_{t=1}^{n}u_{t}W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}})\leq 2\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}}\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}(W_{t-1}^{\top}(\theta^{*}-\widehat{\theta}^{{\rm S}}))^{2}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.103)
Combining (A.101), (A.102), and (A.103), we have
$\widehat{\sigma}_{u}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}+o_{p}(1).$ (A.104)
Using a standard law of large number, we deduce
$\dfrac{1}{n}\sum_{t=1}^{n}u_{t}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}.$ (A.105)
(A.104) and (A.105) imply that $\widehat{\sigma}_{u}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}$. Then for (A.97) it suffices to show
$\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|/\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm sgn}(G_{j}^{*}).$
CASE I: $j\in\mathcal{M}_{x}$. Define
$\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j},$ (A.106)
and by definition of $\widehat{r}_{j,t}$ in (2.11) we have
$\check{r}_{j,t}=\zeta_{j,t}-\widehat{\varsigma}_{j}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}.$ (A.107)
Then
$\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}-\widehat{\varsigma}_{j}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}.$ (A.108)
Note that $w_{j,t}$ is unit root and $\zeta_{j,t}$ is the IV. By the functional CLT for the case of local unit roots in phillips2016robust, we have
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}G_{j}^{*}:=\dfrac{1}{C_{\zeta}}\left({\rm lvar}(e_{j,t})+\int_{0}^{1}\mathcal{U}_{j}d\mathcal{U}_{j}\right).$ (A.109)
By the bound of $\min_{k\in[p]}\widehat{\sigma}_{k}$ by (A.35) and (A.36), the bound of LUR processes (A.25), and the bound of a stationary component (A.78), we have
$\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}\leq\dfrac{\sup_{k\in\mathcal{M}_{x}}|w_{k,t}|}{\min_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}}+\dfrac{\sup_{k\in\mathcal{M}_{z}}|w_{k,t}|}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}=O_{p}(\log p).$ (A.110)
Note that in (A.75), we allow for a small $C_{\varphi}$ that shrinks to zero as $n\to\infty$. Let $C_{\varphi}=1/(4C_{\sigma}s^{2}n^{1-\tau/2})$ with the absolute constant $C_{\sigma}$ in (A.75). Then
$\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\leq\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\leq\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}},$ (A.111)
where the first inequality applies (A.75). Therefore,
$\left|n^{-(1+\tau)}\widehat{\varsigma}_{j}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\leq\widehat{\varsigma}_{j}\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq\widehat{\varsigma}_{j}\lVert n^{-(1+\tau)}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\rVert_{\infty}\cdot\dfrac{4C_{\sigma}C_{\varphi}s^{2}(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\leq\dfrac{\widehat{\varsigma}_{j}\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}|w_{j,t-1}|}{n^{\tau}}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{n^{\tau/2}\cdot\sqrt{n}(\log p)^{\frac{3}{2}}}{n^{\tau}\cdot}\cdot\dfrac{(\log p)^{\frac{1}{2}+\frac{1}{2r}}}{n^{1-\tau/2}\cdot\sqrt{n^{(1-\tau)\wedge\tau}}}=\dfrac{(\log p)^{\frac{3}{2}+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\to 0,$
where the fourth inequality applies $\widehat{\varsigma}_{j}=O_{p}(n^{\tau/2})$ in (A.43), and the bound of $\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t-1}\|_{\infty}$ in (A.110). Thus,
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}G_{j}^{*}.$ (A.112)
By $\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j}$ and the continuous mapping theorem,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}=\dfrac{\left|\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}{\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\dfrac{\left|G_{j}^{*}\right|}{G_{j}^{*}}={\rm sgn}(G_{j}^{*}).$ (A.113)
Then (A.97) is implied by (A.100) and (A.113).
For (A.99), we have $\check{r}_{j,t}^{2}=\zeta_{j,t}^{2}+\widehat{\varsigma}_{j}^{2}(W_{-j,t}^{\top}\widehat{\varphi}^{(j)})^{2}-2\widehat{\varsigma}_{j}\zeta_{j,t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ . When $j\in\mathcal{M}_{x}$,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $\leq\widehat{\varsigma}_{j}^{2}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}^{2}\cdot\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$
$\ \ \ \ +2\widehat{\varsigma}_{j}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}\cdot\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}.$ (A.114)
By (A.35), (A.36), (A.79), and (A.80),
$\|n^{-1}\widetilde{W}_{-j,\cdot}^{\top}\widetilde{W}_{-j,\cdot}\|_{\infty}$ $\leq\max_{j\in[p],k\in\mathcal{M}_{x}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\dfrac{1}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}+\max_{j\in[p],k\in\mathcal{M}_{z}}\left|\frac{1}{n}\sum_{t=1}^{n}\dfrac{w_{j,t}w_{k,t}}{\widehat{\sigma}_{j}}\right|\dfrac{1}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}$
$=O_{p}\left((\log p)^{2}\right).$ (A.115)
Furthermore,
$\|\dfrac{1}{n}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\zeta_{j,t}\|_{\infty}$ $\leq\dfrac{\|\dfrac{1}{n}\sum_{t=1}^{n}X_{-j,t-1}\zeta_{j,t}\|_{\infty}}{\inf_{j\in\mathcal{M}_{x}}\widehat{\sigma}_{j}}+\dfrac{\|\dfrac{1}{n}\sum_{t=1}^{n}Z_{-j,t-1}\zeta_{j,t}\|_{\infty}}{\inf_{j\in\mathcal{M}_{z}}\widehat{\sigma}_{j}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}+(\log p)^{1+\frac{1}{2r}}$
$=O\left(n^{\tau-1/2}(\log p)^{\frac{3}{2}+\frac{1}{2r}}\right)=o_{p}(n^{\tau/2}).$ (A.116)
where the second inequality applies (A.41), (A.42), (A.35), and (A.36). Combining (A.114), (A.115), (A.116), the rate of $\widehat{\varsigma}_{j}$ in (A.43), and Proposition 1, we have
$\left|\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}-\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\right|$ $=O_{p}\left(n^{\tau-\tau\wedge(1-\tau)}(\log p)^{4+1/r}\right)+O_{p}(n^{\tau/2})\cdot o_{p}(n^{\tau/2})$
$=o_{p}\left(n^{\tau}\right).$ (A.117)
In addition, (A.64) implies $\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}n^{\tau}$. Then by (A.117),
$\left|\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}-1\right|\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0$
as $n\to\infty$, or equivalently
$\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}\bigg/\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.118)
Recall from (A.64) and (A.70) that $\frac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}$ and $\frac{\widehat{\varsigma}_{j}^{2}}{n^{\tau}}$ have the same probability limit, and thus
$\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\bigg/\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$ (A.119)
Recall that $\widehat{r}_{j,t}=\check{r}_{j,t-1}/\widehat{\varsigma}_{j}$ as shown in (A.106). Thus, (A.118) and (A.119) imply
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}=\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}/\widehat{\varsigma}_{j}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1.$
Then (A.99) is verified for CASE I.
CASE II: $j\in\mathcal{M}_{z}$. The definition of $\widehat{r}_{j,t}$ gives
$\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}=\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}+\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\left(\widehat{\varphi}^{(j)}-\varphi^{*(j)}\right),$ (A.120)
where $\widetilde{\eta}_{j,t}$ is defined below (3.14). Note that
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}$ $=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}(w_{j,t-1}-Z_{-j,t-1}^{\top}\varphi_{0z}^{(j)*})$
$=\frac{1}{n\widehat{\varsigma}_{j}}\sum_{t=1}^{n}w_{j,t-1}\eta_{j,t-1}^{(1)}.$
By (A.91), we have
$\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}$ (A.121)
where $\eta_{j,t}^{(1)}=w_{j,t}-Z_{-j,t}^{\top}\varphi_{0z}^{(j)*}$ was defined in (A.93). In addition, we deduce that
$\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\|_{\infty}$ $\leq\dfrac{\max_{k\in\mathcal{M}_{x}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|}{\min_{k\in\mathcal{M}_{x}}\widehat{\sigma}_{k}}+\dfrac{\max_{k\in\mathcal{M}_{z}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|}{\min_{k\in\mathcal{M}_{z}}\widehat{\sigma}_{k}}$
$=O_{p}\left(\sqrt{\dfrac{\log p}{n}}\cdot(\log p)^{1+\frac{1}{2r}}\right)+O_{p}(1)=O_{p}(1),$
where the second step follows (A.38) bounding the cross product between LURs and a stationary component, (A.35) and (A.35) bounding the standard deviations, and the fact that
$\max_{k\in\mathcal{M}_{z}}\left|n^{-1}\sum_{t=1}^{n}w_{k,t-1}w_{j,t-1}\right|=O_{p}(1)$
following MS24’s Eq.(B.28). Thus, by Proposition 1 in our main text,
$\left|\dfrac{1}{n}\sum_{t=1}^{n}w_{j,t-1}W_{-j,t-1}^{\top}\left(\widehat{\varphi}^{(j)}-\varphi^{*(j)}\right)\right|$ $\leq\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}w_{j,t-1}\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(j)}-\varphi^{*(j)})\|_{1}$
$=o_{p}(1).$ (A.122)
Combining (A.120), (A.121), and (A.122),
$\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}{\sqrt{{\rm var}\left(w_{j,t}\right)}}.$ (A.123)
Thus,
$\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{\left|{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)\right|}{{\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)}={\rm sgn}\left({\rm cov}\left(w_{j,t},\eta_{j,t}^{(1)}\right)\right),$
which together with (A.100) implies (A.97) for $j\in\mathcal{M}_{z}$.
For (A.99), note that $\widehat{r}_{j,t}$ is the LASSO residual of regression (2.11). Recall that $\widetilde{\eta}_{j,t}$ is the error term of the pseudo-true model (3.14). Following the ideas in the proof of (A.104), we can show that $n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ and $n^{-1}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}$ share the same probability limit, which is
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}$ $={\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widetilde{\eta}_{j,t}^{2}.$
$={\rm plim}_{n\to\infty}\dfrac{\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{\widehat{\varsigma}_{j}^{2}}=\dfrac{{\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{{\rm var}(w_{j,t})},$ (A.124)
where the last step applies (A.91). In addition, note that
$\eta_{j,t}$ $=\eta_{j,t-1}^{(1)}+\zeta_{j,t}-w_{j,t}=\eta_{j,t-1}^{(1)}-n^{-\tau}C_{\zeta}\phi_{j,t-1},$ (A.125)
where the first equality applies (A.94), and the second equality applies (A.82). Therefore,
$\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}$ $=\dfrac{1}{n}\sum_{t=1}^{n}(\eta_{j,t-1}^{(1)})^{2}+\dfrac{C_{\zeta}^{2}}{n^{2\tau+1}}\sum_{t=1}^{n}\phi_{j,t-1}^{2}-\dfrac{2C_{\zeta}}{n^{1+\tau}}\sum_{t=1}^{n}\phi_{j,t-1}\eta_{j,t-1}^{(1)}$
$=\dfrac{1}{n}\sum_{t=1}^{n}(\eta_{j,t-1}^{(1)})^{2}+o_{p}(1)$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}(\eta_{j,t}^{(1)}),$
where the second row applies the same arguments for (A.89). Then we have
${\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t}^{2}=\dfrac{{\rm plim}_{n\to\infty}\dfrac{1}{n}\sum_{t=1}^{n}\eta_{j,t}^{2}}{{\rm var}(w_{j,t})}=\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{{\rm var}(w_{j,t})}.$
This completes the proof of Lemma A.12. ∎
A.3.2 Proofs of main results in Section 3.3
Proof of Theorem 1.
By the definition of the XDlasso estimator,
$\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}$ $=\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}+\frac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}.$
Then the $t$-statistic can be decomposed as
$\dfrac{\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}}{\widehat{\omega}_{j}^{{\rm XD}}}$ $=\mathcal{Z}_{j}+\varDelta_{j},\text{ where }$
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}},$ (A.126)
$\varDelta_{j}$ $=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})}{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}.$ (A.127)
We first bound $\varDelta_{j}$. By the Karush-Kuhn-Tucker condition, we can establish
$\|\sum_{t=1}^{n}D_{-j}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\leq\dfrac{C_{{\rm a}}(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}$ (A.128)
w.p.a.1 as in (3.16). Thus,
$\left|\varDelta_{j}\right|$ $\leq\sqrt{n}\dfrac{\left|n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}W_{-j,t-1}^{\top}D_{-j}^{-1}D_{-j}(\theta_{-j}^{*}-\widehat{\theta}_{-j})\right|}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}$
$\leq\dfrac{\sqrt{n}\cdot\|n^{-1}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}\widehat{r}_{j,t-1}\|_{\infty}\cdot\|D_{-j}(\theta{}_{-j}^{*}-\widehat{\theta}_{-j})\|_{1}}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{\sqrt{n}}{\sqrt{n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\left(\dfrac{(\log p)^{2+\frac{1}{2r}}}{\sqrt{n^{(1-\tau)\wedge\tau}}}\right)\cdot\left(\dfrac{s^{2}}{\sqrt{n}}(\log p)^{6+\frac{1}{2r}}\right)$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\dfrac{s^{2}(\log p)^{8+\frac{1}{r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0$ (A.129)
where the third inequality applies (A.128) and Lemma 2, and the fourth inequality applies
$n^{-1}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ (A.130)
implied by (A.99).
It then suffices to show that $\mathcal{Z}_{j}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ Given the limit in (A.97), when $j\in\mathcal{M}_{z}$ it suffices to show
$\mathcal{Z}_{j}^{(1)}:=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}u_{t}}{\text{$\sigma_{u}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.131)
When $j\in\mathcal{M}_{x}$, we need to additionally show the asymptotic distribution in (A.131) is independent of the Brownian motion in the $G_{j}^{*}$ of limit (A.97), so that
$\mathcal{Z}_{j}$ $=\dfrac{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}\right|}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t}}\cdot\dfrac{\sigma_{u}}{\widehat{\sigma}_{u}}\cdot\mathcal{Z}_{j}^{(1)}$ (A.132)
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}{\rm sgn}(G_{j}^{*})\cdot\mathcal{N}(0,1)=\mathcal{MN}\left(0,{\rm sgn}(G_{j}^{*})^{2}\right)=\mathcal{N}(0,1),$
where $\mathcal{MN}$ denotes a mixed normal distribution.
CASE I. When $j\in\mathcal{M}_{x}$,
$\mathcal{Z}_{j}^{(1)}=-\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\cdot$
We bound the first term by
$\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-j,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}},$ (A.133)
where the second step applies (3.8) and Proposition 1. Therefore, by (A.99)
$\dfrac{\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\widehat{\varphi}^{(j)}\right|}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}/\sqrt{n^{(1-\tau)\wedge\tau}}\to 0,$
which implies
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+o_{p}(1)\cdot$ (A.134)
In addition, by the central limit theorem in Lemma B4(ii) of kostakis2015robust and Eq. (28) in their appendix, the law of large numbers Eq. (13) and (21) in the appendix of the same reference, and the Slutsky’s Theorem, we have
$\frac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.135)
Besides, phillips2009econometric’s Lemma 3.2 shows that the asymptotic distribution in (A.135) is independent of the Brownian motion in the expression (A.98) of $G_{j}^{*}$. Also, recall that $\widetilde{\zeta}_{j,t-1}=\zeta_{j,t-1}\widehat{\varsigma}_{j}$ and $\check{r}_{j,t-1}=\widehat{r}_{j,t-1}\widehat{\varsigma}_{j}.$ By (A.118) and the Slutsky’s Theorem, we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\dfrac{n^{-1/2}\sum_{t=1}^{n}\widetilde{\zeta}_{j,t-1}u_{t}}{\sigma_{u}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\cdot\dfrac{\widehat{\varsigma}_{j}}{\widehat{\varsigma}_{j}}=\sqrt{\dfrac{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}{\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}}\cdot\dfrac{n^{-1/2}\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}}{\sigma_{u}\sqrt{\frac{1}{n}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$
This completes the proof of (A.131) when $j\in\mathcal{M}_{x}.$
CASE II. When $j\in\mathcal{M}_{z}$, recall that we have defined $\eta_{j,t}=\zeta_{j,t}-Z_{-j,t}^{\top}\varphi_{0,z}^{(j)*}$ in (A.83), with $\varphi_{0,z}^{(j)*}$ defined as (3.12). Then
$\widehat{r}_{j,t}$ $=\widetilde{\zeta}_{j,t}-W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t}}{\widehat{\varsigma}_{j}}$
$=W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\dfrac{\eta_{j,t-1}^{(1)}-n^{-\tau}C_{\zeta}\phi_{j,t-1}}{\widehat{\varsigma}_{j}},$ (A.136)
where the first equality is by the definition of $\widehat{r}_{j,t}$ in (2.11), the second row applies the pseudo-true regression model (3.14) and the equality (A.84), and the third row applies (A.125). Then by the definition of $\mathcal{Z}_{0,j}$ in (A.131), we have the following decomposition
$\mathcal{Z}_{j}^{(1)}=\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+\dfrac{n^{-1/2}\sum_{t=1}^{n}\left(\eta_{j,t-1}^{(1)}+n^{-\tau}\phi_{j,t-1}\right)u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}.$ (A.137)
We first bound the first term by
$\left|n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\right|$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\|n^{-1/2}\sum_{t=1}^{n}D_{-j}^{-1}W_{-h,t-1}u_{t}\|_{\infty}\cdot\|D_{-j}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)\|_{1}$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}(\log p)^{3/2+1/(2r)}\cdot\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\to 0,$ (A.138)
where the second step applies MS24’s (B.63), and Proposition 1 in this current paper. By (A.99), we have $\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\left|\dfrac{n^{-1/2}\sum_{t=1}^{n}u_{t}W_{-j,t-1}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)}{\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\right|=o_{p}(1).$ (A.139)
We then show the central limit theorem for the second term. Recall that $u_{t}$ is m.d.s. by Assumption 1, and $\eta_{j,t}^{(1)}$ is stationary and strong mixing. By a standard martingale central limit theorem we have $\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ By (A.99) and (A.91), we have $\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}{\rm var}(\eta_{j,t}^{(1)})$. Thus, by the Slutsky’s Theorem we have
$\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}=\sqrt{\dfrac{{\rm var}(\eta_{j,t}^{(1)})}{\widehat{\varsigma}_{j}^{2}\cdot\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}}\cdot\frac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\sigma_{u}\sqrt{{\rm var}(\eta_{j,t}^{(1)})}}\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}(0,1).$ (A.140)
Finally, note that $\phi_{j,t-1}$ is mildly integrated. Again by kostakis2015robust we have $\sum_{t=1}^{n}\phi_{j,t-1}u_{t}=O_{p}\left(n^{(\tau+1)/2}\right),$ and thus
$n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.141)
By (A.99) and (A.91), we further have $\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}\stackrel{{\scriptstyle\mathrm{p}}}{{\asymp}}1$ and thus
$\dfrac{n^{-1/2}\sum_{t=1}^{n}n^{-\tau}\phi_{j,t-1}u_{t}}{\widehat{\varsigma}_{j}\text{$\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}0.$ (A.142)
By (A.137), (A.138), (A.140), and (A.142), we have
$\mathcal{Z}_{j}^{(1)}$ $=o_{p}(1)+\dfrac{n^{-1/2}\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}}{\text{$\sigma_{u}\widehat{\varsigma}_{j}\cdot\sqrt{\dfrac{1}{n}\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}+o_{p}(1)$ (A.143)
$\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}0,$ (A.144)
which verifies (A.131) when $j\in\mathcal{M}_{z}.$ This completes the proof of Theorem 1. ∎
Proof of Theorem 2.
Recall that
$\widehat{\omega}_{j}^{{\rm XD}}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|}=\frac{\text{$\sqrt{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}$}}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}$ (A.145)
where $\check{r}_{j,t}=\widehat{r}_{j,t}\cdot\widehat{\varsigma}_{j}$ as defined in (A.106).
CASE I. When $j\in\mathcal{M}_{x}$, (A.118) and (A.64) implies
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}=\dfrac{\sum_{t=1}^{n}\check{r}_{j,t-1}^{2}}{\sum_{t=1}^{n}\zeta_{j,t-1}^{2}}\cdot\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j,t-1}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{{\rm lvar}(e_{j,t})}{2C_{\zeta}},$ (A.146)
which implies $\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}=O_{p}(n^{1+\tau}).$ In addition, the weak convergence (A.112) implies that $\frac{1}{\left|\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\right|}=O_{p}\left(\frac{1}{n^{1+\tau}}\right)$. Then
$\widehat{\omega}_{j}^{{\rm XD}}=O_{p}\left(\dfrac{\sqrt{n^{1+\tau}}}{n^{1+\tau}}\right)=O_{p}\left(\dfrac{1}{n^{(1+\tau)/2}}\right).$
CASE II. When $j\in\mathcal{M}_{z}$, by (A.123) and (A.99),
$\widehat{\omega}_{j}^{{\rm XD}}=O_{p}\left(\dfrac{\sqrt{n}}{n}\right)=O_{p}\left(\dfrac{1}{\sqrt{n}}\right).$
We complete the proof of Theorem 2. ∎
Proof of Theorem 3.
Define $\widehat{\Pi}_{j}=\sum_{t=1}^{n}\check{r}_{j,t-1}\left[u_{t}+W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})\right]$, and $\widehat{\Pi}_{\mathcal{A}}=(\widehat{\Pi}_{j})_{j\in\mathcal{A}}$for any subset $\mathcal{A}\in[p]$. Note that
$\widehat{\theta}_{j}^{{\rm XD}}-\theta_{j}^{*}$ $=\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\left[u_{t}+W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})\right]}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}}=\dfrac{\widehat{\Pi}_{j}}{\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}},$ (A.147)
where the second row applies the fact the equality $\check{r}_{j,t}=\widehat{r}_{j,t}\widehat{\varsigma}_{j}$ in (A.106). Furthermore, define the matrix $\widehat{\Theta}_{\mathcal{J}}=(\sum_{t=1}^{n}\check{r}_{j,t-1}\check{r}_{k,t-1})_{j\in\mathcal{J},k\in\mathcal{J}}.$ Also, note that
$\widehat{\Omega}_{\mathcal{J}}^{{\rm XD}}$ $=\widehat{\sigma}_{u}^{2}\left(\dfrac{\sum_{t=1}^{n}\widehat{r}_{j,t-1}\widehat{r}_{k,t-1}}{\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\sum_{t=1}^{n}\widehat{r}_{k,t-1}w_{k,t-1}}\right)_{j,k\in\mathcal{J}}$
$=\widehat{\sigma}_{u}^{2}\left(\dfrac{\sum_{t=1}^{n}\check{r}_{j,t-1}\check{r}_{k,t-1}}{\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1}\sum_{t=1}^{n}\check{r}_{k,t-1}w_{k,t-1}}\right)_{j,k\in\mathcal{J}}$
$=\widehat{\sigma}_{u}^{2}\left[{\rm diag}(\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1})_{j\in\mathcal{J}}\right]^{-1}\widehat{\Theta}_{\mathcal{J}}\left[{\rm diag}(\sum_{t=1}^{n}\check{r}_{j,t-1}w_{j,t-1})_{j\in\mathcal{J}}\right]^{-1}.$ (A.148)
By (A.147) and (A.148), some fundamental calculation yields that under $\mathbb{H}_{0}:\theta_{\mathcal{J}}^{*}=\theta_{0,\mathcal{J}},$
${\rm Wald}_{\mathcal{J}}^{{\rm XD}}=\dfrac{1}{\widehat{\sigma}_{u}^{2}}\widehat{\Pi}_{\mathcal{J}}^{\top}\widehat{\Theta}_{\mathcal{J}}^{-1}\widehat{\Pi}_{\mathcal{J}}.$
The proof will consist of the following essential steps: ∎
  1. 1.
    Show that
    $\dfrac{1}{\sqrt{n^{1+\tau}}}\left|\widehat{\Pi}_{j}-\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}\right|=o_{p}(1)\text{ for any fixed }j\in\mathcal{M}_{x}.$ (A.149)
  2. 2.
    Show that
    $\dfrac{1}{\sqrt{n}}\left|\widehat{\Pi}_{j}-\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}\right|=o_{p}(1)\text{ for any fixed }j\in\mathcal{M}_{z}.$ (A.150)
  3. 3.
    Show that
    $\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}u_{t}\\ \dfrac{1}{\sqrt{n}}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}u_{t}\end{array}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}\left(0,\sigma_{u}^{2}\Theta_{\mathcal{J}}\right),$ (A.151)
    where $\zeta_{\mathcal{J}_{x},t-1}=(\zeta_{j,t-1})_{j\in\mathcal{J}_{x}}$, $\eta_{\mathcal{J}_{z},t-1}^{(1)}=(\eta_{j,t-1}^{(1)})_{j\in\mathcal{M}_{z}}$, and $\Theta_{\mathcal{J}}$ is a nonrandom positive definite matrix.
  4. 4.
    Show that
    $\widetilde{\Theta}_{\mathcal{J}}:=\left(\begin{array}[]{cc}\sqrt{n^{1+\tau}}\mathbf{I}_{|\mathcal{J}_{x}|}\\ &\sqrt{n}\mathbf{I}_{|\mathcal{J}_{z}|}\end{array}\right)^{-1}\widehat{\Theta}_{\mathcal{J}}\left(\begin{array}[]{cc}\sqrt{n^{1+\tau}}\mathbf{I}_{|\mathcal{J}_{x}|}\\ &\sqrt{n}\mathbf{I}_{|\mathcal{J}_{z}|}\end{array}\right)^{-1}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\Theta_{\mathcal{J}}.$ (A.152)
    Equations (A.149), (A.150), and (A.151) imply that
    $\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\widehat{\Pi}_{\mathcal{J}_{x}}\\ \dfrac{1}{\sqrt{n}}\widehat{\Pi}_{\mathcal{J}_{z}}\end{array}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}\left(0,\sigma_{u}^{2}\Theta_{\mathcal{J}}\right).$ (A.153)
    Recall that we have shown $\widehat{\sigma}_{u}^{2}/\sigma_{u}^{2}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}1$ in (A.100). By (A.152), (A.153), and the Slutsky’s Theorem, we have
    $\widehat{\sigma}_{u}^{-1}\widetilde{\Theta}_{\mathcal{J}}^{-1/2}\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\widehat{\Pi}_{\mathcal{J}_{x}}\\ \dfrac{1}{\sqrt{n}}\widehat{\Pi}_{\mathcal{J}_{z}}\end{array}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\mathcal{N}\left(0,\mathbf{I}_{|\mathcal{J}|}\right),$
    and thus
    ${\rm Wald}_{\mathcal{J}}^{{\rm XD}}=\dfrac{1}{\widehat{\sigma}_{u}^{2}}\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\widehat{\Pi}_{\mathcal{J}_{x}}\\ \dfrac{1}{\sqrt{n}}\widehat{\Pi}_{\mathcal{J}_{z}}\end{array}\right)^{\top}\widetilde{\Theta}_{\mathcal{J}}^{-1}\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\widehat{\Pi}_{\mathcal{J}_{x}}\\ \dfrac{1}{\sqrt{n}}\widehat{\Pi}_{\mathcal{J}_{z}}\end{array}\right)\stackrel{{\scriptstyle\mathrm{d}}}{{\to}}\chi_{|\mathcal{J}|}^{2},$
    which verifies Theorem 3.
Proof of (A.149). By (A.107), we have $\check{r}_{j,t}=\zeta_{j,t}-\widehat{\varsigma}_{j}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$. Therefore,
$\widehat{\Pi}_{j}$ $=\sum_{t=1}^{n}\check{r}_{j,t-1}u_{t}+\sum_{t=1}^{n}\check{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})$
$=\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}-\widehat{\varsigma}_{j}\sum_{t=1}^{n}u_{t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}+\widehat{\varsigma}_{j}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}\varDelta_{j}$
$=\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}+O_{p}(\sqrt{n^{\tau}})o_{p}(\sqrt{n})+O_{p}(\sqrt{n^{\tau}})O_{p}(\sqrt{n})o_{p}(1)$
$=\sum_{t=1}^{n}\zeta_{j,t-1}u_{t}+o_{p}(\sqrt{n^{1+\tau}}),$
where the second row applies the definition of $\varDelta_{j}$ in (A.127), and third row applies $\widehat{\varsigma}_{j}=O_{p}(\sqrt{n^{\tau}})$ by (A.43), $\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}=O_{p}(n)$ by (A.124), the rate of $\varDelta_{j}$ by (A.129), and the rate of $\sum_{t=1}^{n}u_{t}W_{-j,t}^{\top}\widehat{\varphi}^{(j)}$ by (A.133).
Proof of (A.150). By (A.136) and the definition $\check{r}_{j,t}=\widehat{\varsigma}_{j}\widehat{r}_{j,t}$, we have
$\check{r}_{j,t}=\widehat{\varsigma}_{j}W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)+\eta_{j,t-1}^{(1)}-n^{-\tau}C_{\zeta}\phi_{j,t-1}.$ (A.154)
Similar to the proof of (A.149), we have
$\widehat{\Pi}_{j}$ $=\sum_{t=1}^{n}\check{r}_{j,t-1}u_{t}+\sum_{t=1}^{n}\check{r}_{j,t-1}W_{-j,t-1}^{\top}(\theta_{-j}^{*}-\widehat{\theta}_{-j})$
$=\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}+\widehat{\varsigma}_{j}\sum_{t=1}^{n}u_{t}W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)-\dfrac{C_{\zeta}}{n^{\tau}}\sum_{t=1}^{n}\phi_{j,t-1}u_{t}+\widehat{\varsigma}_{j}\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}}\varDelta_{j}$
$=\sum_{t=1}^{n}\eta_{j,t-1}^{(1)}u_{t}+o_{p}(\sqrt{n}),$
where the third row applies $\widehat{\varsigma}_{j}=O_{p}(1)$ by (A.87), and the rate of $\varDelta_{j}$ by (A.129), $\sum_{t=1}^{n}u_{t}W_{-j,t}^{\top}\left(\varphi^{*(j)}-\widehat{\varphi}^{(j)}\right)=o_{p}(\sqrt{n})$ by (A.138), and the rate of $n^{-\tau}\sum_{t=1}^{n}\phi_{j,t-1}u_{t}$ by (A.141).
Proof of (A.151). Following the proof of (31) in phillips2009econometric, we can show the following Lindeberg condition for the IVs of the LURs:
$\lim_{n\to\infty}\mathbb{E}\left(\|n^{-\frac{1+\tau}{2}}\zeta_{\mathcal{J}_{x},t}\|_{2}\cdot\textbf{1}\{\|n^{-\frac{1+\tau}{2}}\zeta_{\mathcal{J}_{x},t}\|_{2}>\epsilon\}\right)=0$
for any fixed $\epsilon>0$. In addition, by standard argument it can be shown that parallel Lindeberg condition holds for $n^{-1/2}\eta_{\mathcal{J}_{z},t}^{(1)}$, since $\eta_{\mathcal{J}_{z},t}^{(1)}$ is a vector of stationary and weakly dependent components.
Let ${\rm var}_{t-1}(\cdot)$ denote the conditional covariance matrix given the information up to time t1t-1. According to the martingale central limit theorem hall1980martingale, it suffices to show that
$\sum_{t=1}^{n}{\rm var}_{t-1}\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\zeta_{\mathcal{J}_{x},t-1}u_{t}\\ \dfrac{1}{\sqrt{n}}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}u_{t}\end{array}\right)$ $\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}\Theta_{\mathcal{J}},$ (A.157)
$\text{ where }\Theta_{\mathcal{J}}$ $=\left(\begin{array}[]{cc}\dfrac{1}{2C_{\zeta}}{\rm lvar}(e_{\mathcal{J}_{x},t})\\ &{\rm var}(\eta_{\mathcal{J}_{z},t}^{(1)})\end{array}\right).$ (A.160)
By Lemma 3.1 (iii) and Equation (14) of phillips2009econometric, we have $\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\zeta_{\mathcal{J}_{x},t-1}^{\top}\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\dfrac{1}{2C_{\zeta}}{\rm lvar}(e_{\mathcal{J}_{x},t}).$ By standard CLT, we can show that $\dfrac{1}{n}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}$. Since $\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}$ is the cross-product of the mildly integrated IVs $\zeta_{\mathcal{J}_{x},t-1}$ and the stationary components $\eta_{\mathcal{J}_{z},t-1}^{(1)},$ we have $\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}=O_{p}(n)$ by Lemma B2 (i) of kostakis2015robust. Therefore,
$\sum_{t=1}^{n}{\rm var}_{t-1}\left(\begin{array}[]{c}\dfrac{1}{\sqrt{n^{1+\tau}}}\zeta_{\mathcal{J}_{x},t-1}u_{t}\\ \dfrac{1}{\sqrt{n}}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}u_{t}\end{array}\right)$ $=\sigma_{u}^{2}\left(\begin{array}[]{cc}\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\zeta_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}\\ \dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}\zeta_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}\end{array}\right)$ (A.165)
$\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\sigma_{u}^{2}\Theta_{\mathcal{J}},$ (A.166)
where $\Theta_{\mathcal{J}}$is defined in (A.160). We complete the proof of (A.151).
Proof of (A.152). By (A.107), we have for any $j\in\mathcal{M}_{x}$
$\sup_{t\in[n]}\left|\check{r}_{j,t}-\zeta_{j,t}\right|$ $=\widehat{\varsigma}_{j}\sup_{t\in[n]}\left|W_{-j,t}^{\top}\widehat{\varphi}^{(j)}\right|$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n^{\tau}}\cdot\sup_{t\in[n]}\|D_{-j}^{-1}W_{-j,t}\|_{\infty}\|D_{-j}\widehat{\varphi}^{(j)}\|_{1}$
$\leq\sqrt{n^{\tau}}\cdot O_{p}(\log p)\cdot o_{p}(n^{-1+\tau/2})=o_{p}(n^{\tau-1/2}).$
where second row applies (A.43) and (A.75). Therefore, for any $j_{1},j_{2}\in\mathcal{M}_{x}$,
$\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{j_{1},t-1}\check{r}_{j_{2},t-1}-\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j_{1},t-1}\zeta_{j_{2},t-1}$
$=$ $\dfrac{o_{p}(n^{\tau-1/2})}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j_{1},t-1}+\dfrac{o_{p}(n^{\tau-1/2})}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{j_{2},t-1}+\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}o_{p}(n^{2\tau-1})$
$=$ $o_{p}(\sqrt{n^{(\tau-1)}(\log p)^{3}})+o_{p}(n^{\tau-1})=o_{p}(1),$ (A.167)
where the second row applies $\sup_{t\in[n]}|\zeta_{j_{1},t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n^{\tau}(\log p)^{3}}$ by (A.40). In addition, by (A.154), we have for any $k\in\mathcal{M}_{z}$,
$\sup_{t\in[n]}\left|\check{r}_{k,t}-\eta_{k,t-1}^{(1)}\right|$ $\leq\widehat{\varsigma}_{k}\cdot\sup_{t\in[n]}\left|W_{-k,t}^{\top}\left(\varphi^{*(k)}-\widehat{\varphi}^{(k)}\right)+n^{-\tau}C_{\zeta}\phi_{k,t-1}\right|$
$\leq O_{p}(1)\cdot\sup_{t\in[n]}\|D_{-k}^{-1}W_{-k,t}\|_{\infty}\|D_{-j}(\widehat{\varphi}^{(k)}-\varphi^{*(k)})\|_{1}$
$=O_{p}(1)\cdot O_{p}(\log p)\cdot O_{p}\left(\dfrac{s^{2}(\log p)^{6+\frac{1}{2r}}}{\sqrt{n^{\tau\wedge(1-\tau)}}}\right)=o_{p}(1/(\log p)^{3}),$
[…]
$\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{k_{1},t-1}\check{r}_{k_{2},t-1}-\dfrac{1}{n}\sum_{t=1}^{n}\eta_{k_{1},t-1}^{(1)}\eta_{k_{2},t-1}^{(1)}$
$=$ $o_{p}\left(\frac{1}{n(\log p)^{3}}\right)\sum_{t=1}^{n}\eta_{k_{1},t-1}^{(1)}+o_{p}\left(\frac{1}{n(\log p)^{3}}\right)\sum_{t=1}^{n}\eta_{k_{2},t-1}^{(1)}+\dfrac{1}{n}\sum_{t=1}^{n}o_{p}\left(\frac{1}{(\log p)^{3}}\right)$ (A.168)
$=$ $o_{p}(1).$
In addition, for any $j\in\mathcal{M}_{x}$ and $k\in\mathcal{M}_{z}$,
$\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\check{r}_{j,t-1}\check{r}_{k,t-1}-\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\zeta_{j,t-1}\eta_{k,t-1}^{(1)}$
$=$ $\dfrac{o_{p}(n^{\tau-1/2})}{n^{1+\tau/2}}\sum_{t=1}^{n}\eta_{k,t-1}^{(1)}+o_{p}\left(\frac{1}{n^{1+\tau/2}(\log p)^{3}}\right)\sum_{t=1}^{n}\zeta_{j,t-1}+\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}o_{p}\left(\frac{n^{\tau-1/2}}{(\log p)^{3}}\right)$ (A.169)
$=$ $o_{p}(n^{(\tau-1)/2})+o_{p}\left(\frac{n^{1+\tau/2}(\log p)^{3/2}}{n^{1+\tau/2}(\log p)^{3}}\right)+o_{p}(1),$ (A.170)
where the second row applies $\sup_{t\in[n]}|\zeta_{j_{1},t}|\stackrel{{\scriptstyle\mathrm{p}}}{{\preccurlyeq}}\sqrt{n^{\tau}(\log p)^{3}}$ by (A.40). Therefore,
$\widetilde{\Theta}_{\mathcal{J}}$ $=\left(\begin{array}[]{cc}\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\check{r}_{\mathcal{J}_{x},t-1}\check{r}_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\check{r}_{\mathcal{J}_{x},t-1}\check{r}_{\mathcal{J}_{z},t-1}^{\top}\\ \dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\check{r}_{\mathcal{J}_{z},t-1}\check{r}_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n}\sum_{t=1}^{n}\check{r}_{\mathcal{J}_{z},t-1}\check{r}_{\mathcal{J}_{z},t-1}^{\top}\end{array}\right)$
$=\left(\begin{array}[]{cc}\dfrac{1}{n^{1+\tau}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\zeta_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\zeta_{\mathcal{J}_{x},t-1}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}\\ \dfrac{1}{n^{1+\tau/2}}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}\zeta_{\mathcal{J}_{x},t-1}^{\top}&\dfrac{1}{n}\sum_{t=1}^{n}\eta_{\mathcal{J}_{z},t-1}^{(1)}\eta_{\mathcal{J}_{z},t-1}^{(1)\top}\end{array}\right)+o_{p}(1)$
$\stackrel{{\scriptstyle\mathrm{p}}}{{\to}}\Theta_{\mathcal{J}},$
where the $o_{p}(1)$ in the second row applies (A.167), (A.168), and (A.169), and the limit follows (A.166). With the essential equations (A.149), (A.150), (A.151) , and (A.152) verified, we complete the proof of Theorem 3.
Appendix B Additional Simulation Results
B.1 Simulation Results with More Nonzero Coefficients
We follow the same setup in Section 4.1, with a modification in (4.4) to have
$\gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},\frac{0.25}{6^{2}},\ldots,\frac{0.25}{10^{2}},0_{p_{z}-10}^{\top})^{\top}.$ (B.1)
The empirical sizes are reported in Table B.1 and B.2, and the empirical power is depicted in Figure B.1. The results mirror the results of the benchmark setup in Section 4, which demonstrates the robust performance of XDlasso in finite sample with more control variables associated with nonzero coefficients.
Table B.1: Empirical size and length of confidence interval: IID innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.037 0.217 0.143 0.099 0.047 0.223 0.377 0.104 0.060 0.230 0.436 0.156
300 0.047 0.155 0.142 0.066 0.047 0.164 0.430 0.078 0.064 0.169 0.524 0.120
400 0.046 0.122 0.140 0.050 0.052 0.133 0.479 0.064 0.067 0.135 0.547 0.096
500 0.047 0.101 0.143 0.040 0.054 0.112 0.498 0.054 0.072 0.115 0.577 0.077
600 0.044 0.087 0.135 0.033 0.045 0.097 0.509 0.047 0.057 0.099 0.579 0.065
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.046 0.215 0.147 0.099 0.046 0.220 0.371 0.101 0.057 0.228 0.498 0.159
300 0.033 0.154 0.145 0.066 0.044 0.162 0.452 0.076 0.052 0.169 0.620 0.130
400 0.039 0.122 0.142 0.050 0.048 0.129 0.517 0.062 0.064 0.134 0.689 0.110
500 0.046 0.101 0.141 0.040 0.048 0.111 0.557 0.053 0.069 0.114 0.704 0.089
600 0.039 0.088 0.148 0.033 0.049 0.095 0.605 0.046 0.070 0.100 0.738 0.077
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.042 0.218 0.141 0.100 0.041 0.215 0.361 0.096 0.047 0.222 0.495 0.140
300 0.051 0.155 0.134 0.066 0.045 0.157 0.435 0.072 0.055 0.166 0.594 0.114
400 0.045 0.122 0.146 0.049 0.047 0.127 0.485 0.059 0.062 0.135 0.649 0.095
500 0.040 0.101 0.146 0.039 0.048 0.108 0.532 0.050 0.060 0.114 0.690 0.084
600 0.037 0.087 0.153 0.033 0.050 0.092 0.581 0.044 0.056 0.099 0.736 0.073
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.044 0.379 0.054 0.327 0.066 0.325 0.078 0.288 0.069 0.323 0.080 0.287
300 0.047 0.298 0.055 0.265 0.067 0.265 0.062 0.240 0.070 0.264 0.065 0.240
400 0.048 0.253 0.054 0.229 0.059 0.229 0.064 0.210 0.062 0.229 0.062 0.210
500 0.044 0.223 0.050 0.204 0.054 0.204 0.057 0.189 0.056 0.205 0.063 0.190
600 0.045 0.201 0.050 0.186 0.054 0.186 0.056 0.174 0.053 0.187 0.057 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.050 0.377 0.062 0.326 0.069 0.325 0.065 0.289 0.074 0.320 0.068 0.284
300 0.047 0.297 0.060 0.265 0.063 0.264 0.064 0.239 0.067 0.261 0.063 0.237
400 0.044 0.252 0.057 0.229 0.053 0.228 0.056 0.209 0.054 0.227 0.060 0.209
500 0.052 0.223 0.051 0.204 0.059 0.203 0.057 0.188 0.060 0.203 0.061 0.188
600 0.045 0.202 0.051 0.186 0.062 0.185 0.060 0.173 0.059 0.186 0.060 0.173
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.042 0.379 0.053 0.328 0.062 0.328 0.055 0.291 0.069 0.315 0.057 0.281
300 0.041 0.299 0.050 0.266 0.062 0.265 0.060 0.241 0.059 0.260 0.060 0.237
400 0.040 0.253 0.052 0.229 0.058 0.228 0.066 0.210 0.056 0.226 0.065 0.208
500 0.048 0.223 0.062 0.205 0.064 0.203 0.069 0.189 0.067 0.202 0.071 0.188
600 0.050 0.202 0.061 0.187 0.064 0.185 0.064 0.173 0.061 0.185 0.067 0.173
Notes: The data generating process corresponds to (4.1). The coefficients are specified in (B.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.2: Empirical size and length of confidence: AR(1) innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.046 0.164 0.151 0.074 0.048 0.168 0.419 0.079 0.073 0.173 0.484 0.134
300 0.046 0.112 0.140 0.048 0.051 0.122 0.460 0.059 0.078 0.125 0.563 0.096
400 0.047 0.088 0.150 0.036 0.051 0.098 0.521 0.048 0.081 0.100 0.590 0.072
500 0.040 0.073 0.151 0.028 0.048 0.083 0.553 0.041 0.071 0.083 0.609 0.056
600 0.049 0.062 0.141 0.024 0.049 0.071 0.561 0.035 0.071 0.072 0.606 0.046
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.044 0.159 0.144 0.073 0.055 0.165 0.393 0.076 0.090 0.174 0.544 0.131
300 0.039 0.113 0.147 0.048 0.052 0.120 0.494 0.057 0.084 0.125 0.650 0.106
400 0.035 0.088 0.140 0.036 0.055 0.096 0.556 0.047 0.085 0.098 0.698 0.087
500 0.043 0.073 0.153 0.029 0.060 0.081 0.605 0.040 0.089 0.084 0.739 0.069
600 0.036 0.063 0.147 0.023 0.055 0.070 0.631 0.035 0.079 0.072 0.753 0.057
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.045 0.162 0.144 0.073 0.053 0.162 0.388 0.072 0.088 0.169 0.543 0.113
300 0.051 0.112 0.149 0.048 0.049 0.117 0.477 0.054 0.079 0.126 0.605 0.087
400 0.034 0.088 0.150 0.035 0.044 0.093 0.547 0.044 0.066 0.100 0.676 0.074
500 0.044 0.074 0.149 0.028 0.052 0.077 0.577 0.037 0.072 0.083 0.717 0.065
600 0.044 0.063 0.155 0.023 0.049 0.067 0.631 0.033 0.067 0.072 0.746 0.057
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for AR(1) regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.040 0.385 0.062 0.316 0.067 0.334 0.075 0.275 0.072 0.331 0.081 0.273
300 0.048 0.301 0.054 0.255 0.065 0.269 0.068 0.229 0.069 0.268 0.074 0.227
400 0.052 0.253 0.050 0.220 0.066 0.231 0.066 0.200 0.068 0.231 0.071 0.200
500 0.048 0.223 0.050 0.196 0.054 0.205 0.060 0.181 0.058 0.205 0.063 0.180
600 0.046 0.200 0.053 0.179 0.054 0.186 0.064 0.166 0.055 0.187 0.068 0.165
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.041 0.386 0.056 0.318 0.067 0.334 0.069 0.275 0.071 0.328 0.076 0.271
300 0.044 0.301 0.049 0.256 0.060 0.268 0.067 0.228 0.065 0.265 0.067 0.225
400 0.039 0.254 0.047 0.220 0.058 0.230 0.057 0.199 0.057 0.228 0.060 0.198
500 0.041 0.223 0.045 0.196 0.051 0.204 0.060 0.179 0.052 0.204 0.060 0.179
600 0.045 0.201 0.049 0.178 0.053 0.185 0.056 0.164 0.056 0.185 0.057 0.164
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.035 0.387 0.044 0.319 0.064 0.337 0.057 0.278 0.061 0.326 0.067 0.269
300 0.042 0.302 0.059 0.257 0.068 0.269 0.067 0.229 0.071 0.265 0.067 0.225
400 0.047 0.254 0.057 0.220 0.065 0.230 0.072 0.199 0.064 0.228 0.070 0.197
500 0.047 0.223 0.061 0.196 0.060 0.204 0.066 0.179 0.057 0.203 0.065 0.178
600 0.044 0.201 0.058 0.178 0.061 0.184 0.065 0.164 0.058 0.184 0.066 0.164
Notes: The data generating process corresponds to (4.1). The coefficients are specified in (B.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Figure B.1: Power curves of XDlasso inference
(a) IID Innovations
(b) AR(1) Innovations
Notes: The left and right panels correspond to DGPs (4.1) and (4.2), respectively. The coefficients are specified in (B.1). In each subplot, the first row depicts the empirical power function for $\beta_{1}^{*}$, associated with a nonstationary regressor, across various $\left(p_{x},p_{z}\right)$ configurations, while the second row pertains to $\gamma_{1}^{*}$, associated with a stationary regressor. The empirical power is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution.
B.2 Simulation Results with Cointegrated Regressors
In this section, we follow the data generating process in Section 4.1 with the same innovation processes. The LUR regressors are generated by $X_{j,t}=\rho_{j}^{*}X_{j,t-1}+e_{j,t},$ for $j=1,2,\cdots,p_{x}-3,p_{x}-1$ with $\rho^{*}=(1,1-1/n,1+1/n,1,1-1/n,1+1/n,\cdots)^{\top}\in\mathbb{R}^{p_{x}-2}$, and $X_{j,t}=X_{j-1,t}-e_{j,t},$ for $j=p_{x}-2$ and $p_{x}$, so that the last four LURs are cointegrated. The true coefficient vectors are:
$\beta^{*}=\left(\beta_{1}^{*},\dfrac{0.5}{\sqrt{n}}\times 1_{4}^{\top},0_{p_{x}-7}^{\top},0.5,-0.5\right)^{\top},\ \ \gamma^{*}=(\gamma_{1}^{*},0.5\times 1_{2}^{\top},0.25\times 1_{2}^{\top},0_{p_{z}-5}^{\top})^{\top},$ (B.2)
so that we include one cointegration residual with nonzero coefficients, while the other is treated as redundant control. The empirical sizes are reported in Table B.3 and B.4. In this setting, XDlasso keeps demonstrating good size control in finite sample with similar performance as in the benchmark setup.
Table B.3: Empirical size and length of confidence interval with cointegrated regressors: IID innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for I(1) regressor (raw data)
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.051 0.217 0.146 0.099 0.040 0.224 0.366 0.106 0.050 0.230 0.408 0.154
300 0.053 0.154 0.150 0.066 0.046 0.167 0.449 0.080 0.060 0.170 0.528 0.121
400 0.056 0.121 0.140 0.049 0.050 0.134 0.490 0.065 0.060 0.137 0.575 0.095
500 0.054 0.101 0.142 0.039 0.043 0.113 0.523 0.055 0.059 0.114 0.593 0.078
600 0.046 0.087 0.137 0.033 0.049 0.098 0.540 0.048 0.065 0.099 0.594 0.064
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.051 0.217 0.149 0.100 0.062 0.224 0.379 0.103 0.067 0.225 0.497 0.160
300 0.043 0.155 0.139 0.066 0.054 0.165 0.484 0.077 0.073 0.170 0.596 0.130
400 0.041 0.121 0.140 0.050 0.057 0.132 0.535 0.063 0.076 0.138 0.675 0.107
500 0.044 0.101 0.130 0.040 0.053 0.111 0.560 0.054 0.074 0.114 0.702 0.089
600 0.044 0.086 0.125 0.033 0.046 0.096 0.609 0.047 0.066 0.099 0.747 0.078
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.041 0.215 0.150 0.099 0.057 0.219 0.346 0.097 0.067 0.223 0.491 0.142
300 0.036 0.154 0.132 0.067 0.054 0.158 0.441 0.073 0.059 0.166 0.585 0.113
400 0.043 0.121 0.136 0.050 0.056 0.129 0.505 0.060 0.062 0.134 0.647 0.096
500 0.048 0.101 0.130 0.040 0.054 0.107 0.545 0.051 0.067 0.112 0.679 0.085
600 0.044 0.087 0.129 0.033 0.059 0.092 0.600 0.044 0.072 0.098 0.729 0.074
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.045 0.371 0.052 0.322 0.066 0.325 0.066 0.288 0.072 0.322 0.072 0.287
300 0.045 0.295 0.047 0.263 0.066 0.265 0.061 0.240 0.069 0.264 0.063 0.239
400 0.037 0.251 0.054 0.228 0.054 0.229 0.061 0.210 0.056 0.229 0.061 0.210
500 0.039 0.222 0.045 0.203 0.054 0.204 0.058 0.189 0.056 0.205 0.060 0.190
600 0.046 0.201 0.051 0.185 0.058 0.187 0.060 0.174 0.059 0.187 0.061 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.037 0.372 0.044 0.323 0.059 0.326 0.062 0.289 0.065 0.321 0.067 0.285
300 0.047 0.294 0.054 0.263 0.064 0.264 0.058 0.240 0.064 0.261 0.064 0.237
400 0.039 0.251 0.047 0.228 0.056 0.228 0.056 0.210 0.057 0.227 0.058 0.209
500 0.045 0.222 0.051 0.204 0.060 0.204 0.058 0.189 0.059 0.204 0.058 0.189
600 0.040 0.201 0.044 0.186 0.057 0.186 0.057 0.173 0.056 0.186 0.057 0.173
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.035 0.372 0.042 0.323 0.056 0.329 0.054 0.292 0.058 0.317 0.057 0.282
300 0.038 0.296 0.056 0.263 0.058 0.265 0.057 0.241 0.061 0.260 0.061 0.236
400 0.043 0.251 0.053 0.228 0.065 0.228 0.060 0.210 0.061 0.226 0.059 0.208
500 0.039 0.221 0.044 0.203 0.057 0.203 0.052 0.188 0.056 0.202 0.053 0.188
600 0.039 0.201 0.043 0.185 0.051 0.185 0.050 0.173 0.053 0.185 0.051 0.172
Notes: The data generating process corresponds to (4.1) with cointegrated regressors described in Section B.2. The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.4: Empirical size and length of confidence interval with cointegrated regressors: AR(1) innovations
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for I(1) regressor (raw data)
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.048 0.161 0.157 0.072 0.065 0.169 0.411 0.079 0.092 0.175 0.500 0.138
300 0.047 0.113 0.154 0.048 0.061 0.123 0.461 0.059 0.079 0.127 0.549 0.095
400 0.044 0.087 0.153 0.035 0.052 0.096 0.518 0.048 0.076 0.100 0.598 0.073
500 0.045 0.073 0.144 0.028 0.053 0.082 0.551 0.040 0.079 0.083 0.610 0.058
600 0.044 0.062 0.139 0.023 0.053 0.070 0.565 0.035 0.071 0.072 0.619 0.047
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.046 0.158 0.150 0.072 0.058 0.166 0.419 0.076 0.100 0.173 0.550 0.134
300 0.046 0.112 0.144 0.047 0.054 0.120 0.512 0.057 0.089 0.126 0.643 0.105
400 0.051 0.088 0.142 0.035 0.052 0.095 0.562 0.047 0.086 0.100 0.681 0.083
500 0.043 0.073 0.134 0.028 0.042 0.080 0.606 0.040 0.076 0.083 0.737 0.068
600 0.043 0.063 0.124 0.023 0.056 0.069 0.639 0.035 0.084 0.072 0.736 0.056
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.050 0.162 0.153 0.072 0.057 0.160 0.385 0.071 0.090 0.170 0.530 0.112
300 0.048 0.113 0.144 0.047 0.061 0.116 0.454 0.054 0.088 0.126 0.615 0.087
400 0.039 0.088 0.142 0.035 0.057 0.093 0.522 0.044 0.079 0.099 0.667 0.071
500 0.049 0.073 0.133 0.029 0.058 0.077 0.580 0.038 0.082 0.084 0.723 0.065
600 0.044 0.063 0.143 0.023 0.050 0.066 0.622 0.033 0.071 0.071 0.753 0.057
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor (raw data)
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.041 0.380 0.051 0.311 0.069 0.334 0.073 0.275 0.076 0.330 0.079 0.272
300 0.045 0.297 0.049 0.253 0.058 0.269 0.061 0.229 0.059 0.268 0.065 0.227
400 0.037 0.251 0.044 0.219 0.052 0.231 0.053 0.200 0.055 0.231 0.056 0.200
500 0.047 0.221 0.048 0.195 0.058 0.206 0.061 0.181 0.060 0.206 0.061 0.180
600 0.051 0.200 0.053 0.177 0.057 0.187 0.060 0.166 0.057 0.187 0.062 0.166
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.038 0.379 0.046 0.314 0.058 0.334 0.052 0.276 0.068 0.329 0.059 0.272
300 0.041 0.297 0.051 0.253 0.057 0.268 0.062 0.228 0.066 0.266 0.068 0.225
400 0.040 0.252 0.050 0.219 0.064 0.230 0.058 0.200 0.067 0.230 0.067 0.198
500 0.049 0.222 0.056 0.195 0.066 0.204 0.066 0.180 0.066 0.205 0.070 0.180
600 0.041 0.200 0.055 0.178 0.067 0.186 0.063 0.165 0.067 0.186 0.068 0.165
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.035 0.380 0.062 0.312 0.067 0.335 0.071 0.277 0.069 0.325 0.077 0.267
300 0.042 0.297 0.059 0.253 0.061 0.268 0.067 0.228 0.067 0.263 0.069 0.224
400 0.039 0.251 0.054 0.219 0.067 0.230 0.061 0.199 0.065 0.227 0.062 0.197
500 0.038 0.221 0.043 0.195 0.056 0.203 0.058 0.179 0.060 0.202 0.053 0.178
600 0.043 0.200 0.047 0.178 0.063 0.184 0.054 0.164 0.057 0.184 0.055 0.163
Notes: The data generating process corresponds to (4.2) with cointegrated regressors described in Section B.2. The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
B.3 Simulation Results on Conditional Heteroskedasticity
In this section, we conduct simulation experiments to investigate the finite sample properties of XDlasso with conditional heteroskedasticity and heteroskedastic-robust standard error. In the experiment, we adapt the data generating process in Section 4.1 to incorporate possibly heteroskedastic error terms. The innovations $v_{t}=\left(u_{0,t},e_{t}^{\top},Z_{t}^{\top}\right)^{\top}$ are generated following (4.1) and (4.2). We examine two cases for the error term $u_{t}$:
$\text{IID Error Term:\,}u_{t}$ $=u_{0,t},$ (B.3)
$\text{GARCH(1,1) Error Term:}\>u_{t}$ $=\sqrt{h_{t}}u_{0,t},\,h_{t}=\alpha_{0}+\alpha_{u}u_{t-1}^{2}+\alpha_{h}h_{t-1},$ (B.4)
where we specify $\alpha_{0}=0.6,\alpha_{u}=\alpha_{h}=0.2$, and initialize $h_{1}=1$.
We consider both the homoskedasticity-only standard error as in (2.14) and the heteroskedasticity-robust standard error given as
$\widehat{\omega}_{j}^{{\rm XD,Robust}}$ $=\dfrac{\sqrt{\sum_{t=1}^{n}\widehat{r}_{j,t-1}^{2}\widehat{u}_{t}^{2}}}{\left|\sum_{t=1}^{n}\widehat{r}_{j,t-1}w_{j,t-1}\right|},$ (B.5)
for the construction of the test statistic in (2.15).
The empirical sizes based on homoskedastic standard errors in DGPs with GARCH(1,1) error term $u_{t}$ are reported in Table B.5 and B.6. The results echo our conjecture in Remark 7 that the homoskedastic standard error (2.14) is robust to conditional heteroskedasticity as in kostakis2015robust.
The empirical sizes based on the heteroskedastic-robust standard error (B.5) are reported in Table B.7 and B.8 when the error term $u_{t}$ is IID, and in Table B.9 and B.10 when $u_{t}$ follows a GARCH model. The finite sample performance of XDlasso with robust standard error demonstrates good size control, and motivates our practice in the empirical analysis carried out in Section C.2.
Table B.5: Empirical size and length of confidence interval with homoskedastic S.E.: IID innovations and GARCH error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor, $\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.043 0.216 0.141 0.097 0.057 0.226 0.367 0.107 0.064 0.226 0.422 0.154
300 0.048 0.153 0.148 0.066 0.054 0.167 0.447 0.080 0.062 0.168 0.524 0.117
400 0.040 0.122 0.146 0.050 0.056 0.135 0.478 0.065 0.070 0.134 0.541 0.093
500 0.040 0.102 0.153 0.040 0.054 0.114 0.522 0.055 0.067 0.114 0.597 0.078
600 0.051 0.088 0.143 0.033 0.060 0.099 0.541 0.048 0.072 0.099 0.604 0.065
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.052 0.216 0.148 0.099 0.050 0.222 0.384 0.104 0.055 0.225 0.497 0.162
300 0.039 0.153 0.137 0.065 0.051 0.164 0.483 0.078 0.061 0.169 0.598 0.128
400 0.048 0.121 0.148 0.049 0.042 0.133 0.529 0.064 0.060 0.137 0.672 0.108
500 0.046 0.101 0.147 0.039 0.048 0.112 0.581 0.054 0.064 0.115 0.714 0.092
600 0.049 0.086 0.138 0.033 0.047 0.097 0.616 0.047 0.059 0.099 0.720 0.076
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.041 0.214 0.162 0.098 0.057 0.221 0.351 0.099 0.063 0.225 0.474 0.141
300 0.046 0.153 0.148 0.066 0.053 0.159 0.455 0.074 0.065 0.165 0.582 0.111
400 0.049 0.121 0.147 0.049 0.056 0.128 0.516 0.060 0.069 0.133 0.638 0.098
500 0.048 0.101 0.153 0.039 0.058 0.108 0.567 0.051 0.068 0.113 0.703 0.083
600 0.054 0.087 0.151 0.033 0.066 0.094 0.600 0.045 0.074 0.099 0.736 0.078
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor, $\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.036 0.372 0.056 0.323 0.064 0.325 0.061 0.289 0.065 0.323 0.066 0.287
300 0.042 0.294 0.057 0.262 0.055 0.265 0.062 0.240 0.056 0.264 0.063 0.240
400 0.047 0.251 0.052 0.227 0.061 0.229 0.059 0.210 0.063 0.228 0.064 0.210
500 0.043 0.221 0.042 0.203 0.056 0.204 0.058 0.189 0.060 0.204 0.057 0.189
600 0.041 0.200 0.049 0.185 0.054 0.186 0.056 0.174 0.056 0.186 0.056 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.045 0.371 0.052 0.323 0.062 0.326 0.059 0.289 0.071 0.319 0.066 0.284
300 0.048 0.294 0.059 0.263 0.065 0.264 0.067 0.240 0.066 0.262 0.071 0.238
400 0.051 0.251 0.058 0.228 0.065 0.228 0.064 0.209 0.066 0.227 0.068 0.209
500 0.043 0.222 0.056 0.203 0.059 0.203 0.067 0.189 0.062 0.203 0.067 0.189
600 0.050 0.200 0.055 0.185 0.060 0.185 0.060 0.173 0.061 0.186 0.059 0.173
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.033 0.372 0.050 0.323 0.062 0.329 0.059 0.292 0.066 0.316 0.066 0.281
300 0.039 0.294 0.043 0.264 0.063 0.266 0.057 0.241 0.060 0.260 0.056 0.236
400 0.041 0.251 0.041 0.228 0.060 0.228 0.050 0.210 0.057 0.226 0.051 0.208
500 0.036 0.222 0.037 0.204 0.055 0.203 0.049 0.189 0.053 0.202 0.048 0.187
600 0.042 0.201 0.050 0.186 0.058 0.185 0.051 0.173 0.054 0.185 0.050 0.173
Notes: The data generating process corresponds to (4.1) and (B.4). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.6: Empirical size and length of confidence with homoskedastic S.E.: AR(1) innovations and GARCH error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor, $\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.055 0.160 0.167 0.072 0.067 0.168 0.403 0.078 0.084 0.172 0.493 0.134
300 0.045 0.113 0.162 0.047 0.048 0.121 0.472 0.058 0.076 0.124 0.576 0.092
400 0.050 0.088 0.163 0.036 0.056 0.096 0.517 0.048 0.085 0.098 0.593 0.071
500 0.047 0.073 0.151 0.028 0.054 0.081 0.550 0.040 0.084 0.083 0.625 0.055
600 0.048 0.063 0.151 0.024 0.058 0.071 0.568 0.035 0.080 0.072 0.607 0.046
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.044 0.159 0.162 0.072 0.060 0.166 0.433 0.075 0.093 0.174 0.542 0.132
300 0.044 0.111 0.153 0.047 0.048 0.121 0.510 0.057 0.080 0.127 0.646 0.104
400 0.048 0.087 0.149 0.035 0.048 0.096 0.543 0.046 0.082 0.100 0.702 0.084
500 0.048 0.073 0.149 0.028 0.053 0.080 0.600 0.039 0.074 0.083 0.737 0.067
600 0.049 0.063 0.141 0.023 0.049 0.069 0.639 0.034 0.072 0.072 0.747 0.056
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.052 0.160 0.167 0.071 0.060 0.158 0.361 0.071 0.097 0.171 0.518 0.110
300 0.057 0.113 0.157 0.047 0.055 0.115 0.460 0.054 0.072 0.124 0.622 0.087
400 0.048 0.088 0.142 0.035 0.054 0.091 0.528 0.044 0.085 0.099 0.689 0.077
500 0.055 0.074 0.151 0.028 0.061 0.076 0.584 0.037 0.089 0.082 0.729 0.065
600 0.049 0.063 0.153 0.023 0.061 0.065 0.623 0.032 0.073 0.071 0.750 0.054
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor, $\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.040 0.379 0.054 0.311 0.062 0.331 0.073 0.273 0.066 0.330 0.075 0.272
300 0.043 0.296 0.050 0.252 0.058 0.267 0.062 0.228 0.059 0.268 0.062 0.227
400 0.043 0.251 0.048 0.218 0.051 0.230 0.067 0.200 0.051 0.230 0.068 0.199
500 0.042 0.221 0.053 0.194 0.059 0.204 0.058 0.180 0.065 0.205 0.061 0.179
600 0.040 0.199 0.047 0.177 0.055 0.186 0.057 0.165 0.056 0.187 0.057 0.165
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.046 0.381 0.057 0.311 0.069 0.332 0.071 0.273 0.072 0.328 0.074 0.270
300 0.043 0.297 0.049 0.254 0.062 0.267 0.059 0.227 0.065 0.266 0.061 0.225
400 0.045 0.252 0.059 0.218 0.062 0.229 0.065 0.198 0.065 0.229 0.070 0.198
500 0.049 0.221 0.058 0.195 0.062 0.204 0.069 0.179 0.062 0.204 0.069 0.179
600 0.045 0.199 0.055 0.178 0.053 0.185 0.063 0.164 0.056 0.186 0.064 0.164
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.038 0.380 0.046 0.314 0.064 0.332 0.059 0.275 0.063 0.325 0.062 0.268
300 0.037 0.298 0.040 0.253 0.065 0.267 0.062 0.227 0.065 0.264 0.057 0.224
400 0.040 0.253 0.040 0.219 0.056 0.228 0.054 0.198 0.054 0.228 0.052 0.197
500 0.039 0.222 0.042 0.195 0.058 0.203 0.055 0.178 0.057 0.203 0.053 0.178
600 0.046 0.200 0.052 0.178 0.063 0.184 0.056 0.164 0.060 0.184 0.057 0.164
Notes: The data generating process corresponds to (4.2) and (B.4). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (2.15) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.7: Empirical size and length of confidence interval with robust S.E.: IID innovations and IID error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.040 0.216 0.145 0.099 0.056 0.220 0.373 0.105 0.057 0.227 0.434 0.159
300 0.041 0.155 0.140 0.066 0.046 0.166 0.450 0.079 0.060 0.171 0.533 0.120
400 0.043 0.121 0.141 0.050 0.052 0.132 0.472 0.064 0.063 0.136 0.551 0.093
500 0.051 0.102 0.159 0.040 0.049 0.113 0.511 0.054 0.057 0.115 0.597 0.077
600 0.043 0.088 0.154 0.033 0.049 0.098 0.547 0.047 0.066 0.100 0.608 0.064
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.048 0.219 0.138 0.098 0.050 0.219 0.381 0.101 0.058 0.227 0.504 0.158
300 0.042 0.154 0.155 0.066 0.040 0.160 0.460 0.076 0.059 0.168 0.596 0.126
400 0.039 0.121 0.150 0.050 0.046 0.129 0.519 0.062 0.063 0.135 0.653 0.108
500 0.051 0.101 0.151 0.039 0.046 0.110 0.564 0.053 0.063 0.114 0.711 0.090
600 0.047 0.088 0.149 0.033 0.044 0.096 0.607 0.046 0.065 0.100 0.746 0.077
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.044 0.219 0.158 0.100 0.051 0.213 0.339 0.097 0.056 0.223 0.475 0.140
300 0.048 0.156 0.159 0.067 0.055 0.158 0.430 0.073 0.063 0.167 0.571 0.113
400 0.055 0.123 0.158 0.050 0.057 0.125 0.494 0.059 0.067 0.133 0.651 0.096
500 0.045 0.102 0.155 0.040 0.057 0.105 0.545 0.050 0.068 0.113 0.702 0.086
600 0.045 0.087 0.138 0.033 0.060 0.090 0.573 0.044 0.070 0.098 0.737 0.077
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.037 0.374 0.046 0.324 0.055 0.325 0.068 0.288 0.059 0.324 0.064 0.287
300 0.038 0.295 0.043 0.263 0.053 0.265 0.056 0.240 0.055 0.265 0.058 0.240
400 0.038 0.251 0.045 0.227 0.057 0.228 0.062 0.210 0.060 0.229 0.065 0.210
500 0.038 0.222 0.051 0.203 0.052 0.204 0.055 0.189 0.055 0.204 0.059 0.189
600 0.040 0.201 0.046 0.185 0.057 0.186 0.057 0.174 0.059 0.187 0.058 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.054 0.372 0.062 0.324 0.074 0.326 0.078 0.289 0.079 0.320 0.085 0.285
300 0.052 0.294 0.059 0.264 0.070 0.264 0.071 0.239 0.074 0.262 0.070 0.238
400 0.048 0.250 0.056 0.228 0.071 0.228 0.066 0.210 0.074 0.228 0.069 0.209
500 0.045 0.222 0.058 0.204 0.061 0.204 0.066 0.189 0.064 0.204 0.068 0.189
600 0.045 0.200 0.054 0.185 0.055 0.185 0.056 0.173 0.056 0.186 0.056 0.173
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.039 0.375 0.054 0.324 0.062 0.328 0.061 0.292 0.071 0.316 0.074 0.281
300 0.036 0.295 0.044 0.263 0.053 0.265 0.056 0.240 0.055 0.260 0.059 0.237
400 0.046 0.251 0.051 0.227 0.060 0.228 0.064 0.210 0.055 0.226 0.059 0.208
500 0.043 0.222 0.049 0.203 0.050 0.203 0.054 0.188 0.049 0.202 0.052 0.188
600 0.038 0.200 0.049 0.185 0.054 0.185 0.061 0.173 0.050 0.185 0.059 0.173
Notes: The data generating process corresponds to (4.1). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (B.5) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.8: Empirical size and length of confidence with robust S.E.: AR(1) innovations and IID error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.044 0.161 0.146 0.071 0.057 0.169 0.412 0.078 0.090 0.173 0.487 0.135
300 0.044 0.112 0.156 0.047 0.058 0.123 0.459 0.059 0.075 0.126 0.570 0.096
400 0.052 0.089 0.149 0.036 0.054 0.098 0.521 0.048 0.076 0.100 0.604 0.072
500 0.046 0.074 0.156 0.029 0.050 0.083 0.549 0.041 0.071 0.084 0.605 0.057
600 0.053 0.063 0.150 0.024 0.052 0.072 0.584 0.035 0.073 0.073 0.615 0.047
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.051 0.163 0.164 0.072 0.053 0.169 0.431 0.076 0.094 0.175 0.548 0.136
300 0.045 0.112 0.150 0.048 0.049 0.119 0.501 0.057 0.084 0.124 0.636 0.102
400 0.047 0.088 0.156 0.036 0.051 0.095 0.555 0.047 0.086 0.100 0.709 0.084
500 0.057 0.073 0.153 0.028 0.049 0.080 0.617 0.040 0.080 0.084 0.746 0.067
600 0.057 0.063 0.145 0.024 0.052 0.069 0.655 0.035 0.075 0.072 0.746 0.057
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.049 0.162 0.168 0.073 0.059 0.160 0.372 0.072 0.079 0.169 0.512 0.108
300 0.050 0.113 0.156 0.048 0.062 0.117 0.458 0.054 0.080 0.124 0.603 0.089
400 0.052 0.088 0.145 0.036 0.058 0.091 0.525 0.044 0.078 0.097 0.672 0.075
500 0.048 0.073 0.152 0.028 0.066 0.077 0.592 0.037 0.084 0.082 0.716 0.065
600 0.051 0.062 0.155 0.024 0.057 0.066 0.622 0.033 0.071 0.071 0.765 0.057
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.037 0.380 0.053 0.313 0.066 0.332 0.067 0.274 0.072 0.332 0.071 0.273
300 0.047 0.297 0.049 0.253 0.063 0.268 0.064 0.228 0.066 0.268 0.068 0.228
400 0.041 0.251 0.052 0.218 0.061 0.230 0.071 0.200 0.065 0.230 0.074 0.199
500 0.046 0.221 0.052 0.195 0.062 0.205 0.066 0.180 0.062 0.205 0.065 0.179
600 0.048 0.199 0.051 0.178 0.056 0.186 0.056 0.165 0.056 0.186 0.056 0.165
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.045 0.380 0.057 0.312 0.078 0.332 0.063 0.274 0.089 0.328 0.076 0.271
300 0.047 0.297 0.049 0.253 0.068 0.267 0.064 0.227 0.073 0.266 0.066 0.226
400 0.050 0.251 0.050 0.218 0.066 0.229 0.057 0.199 0.064 0.229 0.059 0.198
500 0.050 0.221 0.052 0.195 0.060 0.203 0.059 0.179 0.057 0.204 0.059 0.179
600 0.049 0.199 0.053 0.178 0.060 0.185 0.063 0.164 0.060 0.186 0.063 0.164
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.034 0.380 0.053 0.311 0.060 0.334 0.065 0.275 0.061 0.323 0.073 0.267
300 0.039 0.297 0.044 0.252 0.057 0.267 0.059 0.227 0.054 0.263 0.058 0.224
400 0.035 0.251 0.049 0.218 0.059 0.228 0.056 0.198 0.059 0.227 0.055 0.196
500 0.036 0.221 0.044 0.194 0.052 0.202 0.059 0.178 0.054 0.203 0.060 0.178
600 0.038 0.199 0.045 0.177 0.051 0.184 0.054 0.164 0.051 0.185 0.056 0.164
Notes: The data generating process corresponds to (4.2). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (B.5) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.9: Empirical size and length of confidence interval with robust S.E.: IID innovations and GARCH error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{1}:\gamma_{1}^{*}=0.1$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.042 0.218 0.135 0.100 0.045 0.222 0.340 0.105 0.052 0.227 0.415 0.153
300 0.039 0.154 0.155 0.066 0.051 0.164 0.434 0.079 0.059 0.167 0.517 0.118
400 0.045 0.122 0.157 0.050 0.050 0.132 0.462 0.064 0.067 0.135 0.555 0.095
500 0.045 0.102 0.147 0.040 0.054 0.112 0.512 0.054 0.060 0.115 0.591 0.077
600 0.049 0.088 0.142 0.033 0.053 0.098 0.531 0.047 0.064 0.099 0.610 0.065
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.042 0.220 0.162 0.101 0.043 0.224 0.390 0.104 0.055 0.229 0.516 0.160
300 0.048 0.154 0.148 0.067 0.045 0.164 0.471 0.077 0.060 0.170 0.602 0.131
400 0.045 0.123 0.159 0.051 0.046 0.130 0.531 0.063 0.064 0.135 0.675 0.106
500 0.050 0.103 0.154 0.040 0.055 0.110 0.589 0.053 0.077 0.114 0.720 0.091
600 0.051 0.089 0.151 0.034 0.056 0.097 0.598 0.047 0.072 0.100 0.734 0.077
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.044 0.214 0.138 0.098 0.054 0.219 0.335 0.097 0.064 0.226 0.498 0.139
300 0.040 0.154 0.157 0.066 0.056 0.158 0.423 0.073 0.055 0.166 0.594 0.112
400 0.044 0.123 0.153 0.050 0.055 0.127 0.500 0.060 0.061 0.135 0.672 0.095
500 0.049 0.101 0.150 0.040 0.052 0.106 0.546 0.051 0.056 0.113 0.703 0.084
600 0.050 0.088 0.152 0.033 0.059 0.093 0.589 0.044 0.061 0.099 0.726 0.076
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.044 0.370 0.057 0.323 0.067 0.325 0.065 0.288 0.065 0.323 0.070 0.287
300 0.038 0.295 0.054 0.263 0.055 0.265 0.059 0.241 0.055 0.264 0.058 0.239
400 0.050 0.251 0.057 0.228 0.055 0.229 0.056 0.211 0.059 0.229 0.058 0.211
500 0.044 0.222 0.056 0.203 0.054 0.205 0.064 0.190 0.058 0.205 0.064 0.190
600 0.045 0.201 0.053 0.185 0.054 0.187 0.058 0.174 0.054 0.187 0.061 0.174
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.043 0.374 0.061 0.324 0.061 0.327 0.064 0.290 0.065 0.322 0.066 0.286
300 0.043 0.295 0.056 0.263 0.059 0.265 0.057 0.240 0.067 0.263 0.062 0.238
400 0.044 0.251 0.053 0.227 0.053 0.229 0.057 0.210 0.053 0.228 0.058 0.209
500 0.048 0.221 0.055 0.203 0.057 0.204 0.058 0.189 0.058 0.204 0.059 0.189
600 0.047 0.201 0.047 0.185 0.058 0.186 0.058 0.173 0.061 0.186 0.057 0.173
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.040 0.371 0.053 0.325 0.063 0.329 0.064 0.292 0.066 0.316 0.076 0.282
300 0.042 0.295 0.052 0.263 0.057 0.266 0.062 0.241 0.054 0.260 0.064 0.236
400 0.038 0.251 0.049 0.227 0.058 0.228 0.062 0.210 0.055 0.226 0.062 0.208
500 0.042 0.222 0.049 0.203 0.056 0.203 0.063 0.189 0.050 0.202 0.060 0.188
600 0.042 0.200 0.048 0.185 0.058 0.185 0.060 0.173 0.055 0.185 0.059 0.173
Notes: The data generating process corresponds to (4.1) and (B.4). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$, respectively, at a 5% nominal significance level. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (B.5) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
Table B.10: Empirical size and length of confidence with robust S.E.: AR(1) innovations and GARCH error terms
$n$ Oracle Calibrated CV
IVX Oracle OLS Oracle XDlasso Dlasso XDlasso Dlasso
Size Len. Size Len. Size Len. Size Len. Size Len. Size Len.
$\mathbb{H}_{0}:\beta_{1}^{*}=0$ for nonstationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.055 0.161 0.157 0.072 0.063 0.170 0.403 0.079 0.090 0.174 0.497 0.137
300 0.050 0.113 0.157 0.047 0.056 0.121 0.478 0.059 0.085 0.123 0.573 0.093
400 0.061 0.088 0.155 0.035 0.059 0.097 0.526 0.048 0.085 0.099 0.601 0.072
500 0.055 0.074 0.153 0.028 0.052 0.083 0.552 0.041 0.062 0.083 0.613 0.057
600 0.045 0.063 0.141 0.024 0.051 0.072 0.559 0.036 0.072 0.072 0.604 0.046
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.039 0.162 0.158 0.073 0.057 0.172 0.411 0.077 0.095 0.178 0.545 0.136
300 0.050 0.114 0.161 0.048 0.055 0.122 0.512 0.058 0.090 0.127 0.677 0.115
400 0.048 0.090 0.165 0.036 0.064 0.096 0.581 0.047 0.096 0.099 0.706 0.083
500 0.052 0.074 0.157 0.028 0.059 0.080 0.616 0.040 0.089 0.084 0.737 0.072
600 0.054 0.063 0.154 0.024 0.053 0.070 0.648 0.035 0.080 0.072 0.751 0.054
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.041 0.161 0.147 0.072 0.057 0.165 0.363 0.073 0.087 0.172 0.556 0.109
300 0.044 0.115 0.161 0.048 0.060 0.118 0.458 0.055 0.083 0.124 0.628 0.087
400 0.047 0.089 0.159 0.036 0.054 0.092 0.544 0.045 0.079 0.100 0.666 0.072
500 0.041 0.074 0.155 0.028 0.051 0.077 0.587 0.038 0.075 0.083 0.715 0.065
600 0.050 0.064 0.162 0.024 0.054 0.067 0.628 0.033 0.071 0.072 0.744 0.058
$\mathbb{H}_{0}:\gamma_{1}^{*}=0$ for stationary regressor
$\left(p_{x},p_{z}\right)=\left(50,100\right)$
200 0.041 0.379 0.055 0.312 0.060 0.334 0.064 0.274 0.062 0.330 0.068 0.271
300 0.043 0.298 0.054 0.252 0.056 0.269 0.056 0.228 0.057 0.268 0.060 0.227
400 0.045 0.251 0.053 0.218 0.053 0.231 0.061 0.200 0.058 0.231 0.062 0.200
500 0.044 0.221 0.053 0.195 0.057 0.205 0.065 0.180 0.060 0.205 0.067 0.180
600 0.046 0.199 0.052 0.178 0.059 0.187 0.059 0.165 0.060 0.187 0.064 0.165
$\left(p_{x},p_{z}\right)=\left(100,150\right)$
200 0.046 0.382 0.059 0.313 0.061 0.334 0.071 0.275 0.072 0.330 0.078 0.271
300 0.046 0.299 0.053 0.254 0.064 0.269 0.063 0.228 0.064 0.267 0.063 0.226
400 0.039 0.252 0.054 0.219 0.051 0.230 0.055 0.200 0.054 0.230 0.056 0.198
500 0.041 0.221 0.049 0.195 0.052 0.204 0.054 0.179 0.051 0.204 0.058 0.178
600 0.040 0.200 0.045 0.178 0.048 0.186 0.052 0.165 0.049 0.186 0.058 0.165
$\left(p_{x},p_{z}\right)=\left(150,300\right)$
200 0.036 0.380 0.054 0.314 0.060 0.335 0.064 0.276 0.059 0.323 0.067 0.267
300 0.039 0.298 0.054 0.253 0.060 0.268 0.063 0.228 0.061 0.263 0.064 0.224
400 0.040 0.251 0.053 0.218 0.060 0.229 0.072 0.198 0.056 0.227 0.065 0.197
500 0.038 0.221 0.055 0.195 0.062 0.203 0.065 0.179 0.058 0.203 0.063 0.178
600 0.047 0.199 0.049 0.178 0.059 0.184 0.058 0.164 0.053 0.184 0.057 0.164
Notes: The data generating process corresponds to (4.2) and (B.4). The upper and lower panels report the empirical size of testing the null hypotheses $\mathbb{H}_{0}:\beta_{1}^{*}=0$ and $\mathbb{H}_{0}:\gamma_{1}^{*}=0$ at a 5% nominal significance level, respectively. “Size” is calculated as $R^{-1}\sum_{r=1}^{R}\mathbf{1}\left[|t^{\text{XD}(r)}|>\mathrm{\Phi}_{0.975}\right]$ across $R=2,000$ replications, where $t^{\text{XD}(r)}$ is computed based on (B.5) for the $r$-th replication, and the critical value $\mathrm{\Phi}_{0.975}\left(\approx 1.96\right)$ is the 97.5-th percentile of the standard normal distribution. “Len.” refers to the median length of the 95% confidence intervals across replications. The IVX oracle and OLS oracle are infeasible estimators. The “Calibrated” and “CV” columns refer to the methods used for choosing the tuning parameters through calibration and cross-validation, respectively.
B.4 Variance Ratio Test on Slasso Residual
In this section, we examine the finite sample rejection rate of performing the automatic variance ratio test based on wild bootstrap (kim2009automatic) on the first step Slasso residuals. We follow the data generating process in Section 4.1 with AR(1) innovations, where $u_{t}\sim i.i.d.N\left(0.1\right)$, and focus on $\left(p_{x},p_{z}\right)=\left(50,100\right)$ and $n\in\left\{200,400,600\right\}$. Table B.11 reports the proportion of rejection at nominal significance levels 1%, 5% and 10% across $2,000$ replications. Even though Slasso is consistent, we still observe severe over-rejection of performing the variance ratio test on $\widehat{u}_{t}$ in finite sample.
Table B.11: Rej. Rate of the Wild Bootstrapped Automatic Variance Ratio Test on Slasso Residual
$n$ 1% 5% 10%
$200$ 4.4% 14.0% 22.0%
$400$ 3.4% 10.7% 17.2%
$600$ 2.2% 10.0% 15.9%
Notes: We follow the data generating process in Section 4.1 with AR(1) innovations, where $u_{t}\sim i.i.d.N\left(0.1\right)$, and focus on $\left(p_{x},p_{z}\right)=\left(50,100\right)$ and $n\in\left\{200,400,600\right\}$. This table reports the rejection rates of the automatic variance ratio test based on the wild bootstrap (kim2009automatic). The test is applied to the first-step Slasso residuals. Rejection rates are shown at nominal significance levels of 1%, 5%, and 10%, based on 2,000 replications.
Appendix C Additional Empirical Results
C.1 Sensitivity to the Classification of I(2) Time Series and Logarithmic Transformation
To further evaluate the robustness of our findings, we consider two alternative specifications in addition to the main analysis in Section 5. First, we exclude nonstationary variables based on their integration orders as determined by the bootstrap sequential testing procedure of smeekes2015bootstrap, following the summary in smeekes2020unit. Second, we apply only logarithmic transformations as indicated by TCODE, without differencing, and remove highly nonstationary time series according to both TCODE and the classifications in smeekes2020unit. Table C.1 shows that, across all specifications, XDlasso consistently finds no evidence that the log earnings-price ratio log earnings-price ratio has predictive power for stock returns stock returns, which confirms our main findings in Section 5.1. Similarly, the inflation prediction results in Table C.3 are largely in line with those in Section 5.2. The only exception occurs in the full-sample estimation using untransformed data and the classification of smeekes2020unit when labor variables are excluded. In these exceptions, diagnostic tests suggest a violation of the m.d.s. condition, which undermines the validity of the inference.
Table C.1: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in stock return prediction: Alternative set of I(2) variables and transformation
(a) Untransformed Data: Excluding I(2) Variables Based on smeekes2020unit
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.020 -0.008 0.000 0.025 0.012 0.831
(Jan. 1960 - Apr. 2025) (0.016) (0.011) (0.015) (0.010)
Pre-1994 0.035 -0.208 0.049 0.044 0.168 0.494
(Jan. 1960 - Dec. 1993) (0.055) (0.215) (0.042) (0.154)
Post-1994 -0.003 -0.022 0.011 -0.000 -0.009 0.163
(Jan. 1994 - Apr. 2025) (0.009) (0.017) (0.009) (0.017)
(b) Log Transformation: Excluding I(2) Variables Based on TCODE
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.011 -0.005 0.003 0.021 0.012 0.576
(Jan. 1960 - Apr. 2025) (0.014) (0.011) (0.014) (0.010)
Pre-1994 0.039 -0.360 0.033 0.033 0.087 0.489
(Jan. 1960 - Dec. 1993) (0.049) (0.291) (0.048) (0.213)
Post-1994 0.041∗∗ 0.020 0.010 0.045∗∗ 0.033 0.258
(Jan. 1994 - Apr. 2025) (0.020) (0.030) (0.020) (0.029)
(c) Log Transformation: Excluding I(2) Variables Based on smeekes2020unit
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.016 -0.009 0.001 0.022 0.013 0.559
(Jan. 1960 - Apr. 2025) (0.012) (0.016) (0.012) (0.010)
Pre-1994 0.106 -0.394 0.032 0.108∗∗ 0.101 0.476
(Jan. 1960 - Dec. 1993) (0.055) (0.317) (0.054) (0.292)
Post-1994 0.018 -0.001 0.010 0.021 0.016 0.216
(Jan. 1994 - Apr. 2025) (0.017) (0.037) (0.017) (0.027)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table C.3: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in inflation prediction: Alternative set of I(2) variables and transformation
(a) Untransformed Data: Excluding I(2) Variables Based on smeekes2020unit
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.172** 0.118 0.104 0.069** 0.116*** 0.015
(Jan. 1960 - Apr. 2025) (0.070) (0.231) (0.029) (0.041)
Pre-Volcker -0.169 -0.022 0.516 0.018 0.106 0.555
(Jan. 1960 - Jul. 1979) (0.150) (0.285) (0.053) (0.108)
Volcker-Greenspan -0.041 -0.200 0.669 -0.009 -0.189 0.641
(Aug. 1979 - Jan. 2006) (0.207) (0.270) (0.058) (0.123)
Bernanke-Yellen-Powell 0.550* 0.294 0.220 0.081 0.162** 0.094
(Feb. 2006 - Apr. 2025) (0.309) (0.510) (0.050) (0.076)
(b) Log Transformation: Excluding I(2) Variables Based on TCODE
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.014 -0.103 0.118 -0.063*** 0.044 0.005
(Jan. 1960 - Apr. 2025) (0.048) (0.103) (0.016) (0.056)
Pre-Volcker -0.155 -0.056 0.516 -0.029 0.131 0.827
(Jan. 1960 - Jul. 1979) (0.164) (0.266) (0.048) (0.117)
Volcker-Greenspan 0.064 -0.068 0.448 -0.082 -0.057 0.427
(Aug. 1979 - Jan. 2006) (0.141) (0.198) (0.060) (0.332)
Bernanke-Yellen-Powell 0.121 1.231 0.398 -0.040 0.142 0.119
(Feb. 2006 - Apr. 2025) (0.182) (0.754) (0.057) (0.102)
(c) Log Transformation: Excluding I(2) Variables Based on smeekes2020unit
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.016 -0.101 0.110 -0.030** 0.016 0.020
(Jan. 1960 - Apr. 2025) (0.049) (0.102) (0.013) (0.028)
Pre-Volcker -0.137 -0.068 0.430 -0.017 0.052 0.556
(Jan. 1960 - Jul. 1979) (0.166) (0.250) (0.052) (0.110)
Volcker-Greenspan -0.054 -0.136 0.871 -0.014 -0.172 0.702
(Aug. 1979 - Jan. 2006) (0.135) (0.304) (0.064) (0.124)
Bernanke-Yellen-Powell 0.018 0.555 0.541 0.087* 0.098* 0.180
(Feb. 2006 - Apr. 2025) (0.133) (0.439) (0.047) (0.055)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
C.2 Heteroskedasticity-Robust Standard Errors
To assess the robustness of our results to conditional heteroskedasticity, we recompute the heteroskedasticity-robust standard errors in (B.5). Tables C.5 and C.7 present the robust results for stock return and inflation prediction parallel to those in Section 5 but with heteroskedasticity-robust standard errors. Tables C.9 and C.11 report the sensitivity analyses like Section C.1 with robust standard errors. Across all cases, our results remain consistent with those in the baseline analysis.
Table C.5: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in stock return prediction with heteroskedasticity-robust standard errors
(a) TCODE Transformed Data
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.009 0.003 0.074 0.009 0.005 0.216
(Jan. 1960 - Apr. 2025) (0.006) (0.014) (0.006) (0.014)
Pre-1994 0.025*** 0.059** 0.227 0.024** 0.062** 0.296
(Jan. 1960 - Dec. 1993) (0.010) (0.029) (0.009) (0.029)
Post-1994 0.002 -0.001 0.053 0.002 -0.001 0.049
(Jan. 1994 - Apr. 2025) (0.006) (0.015) (0.006) (0.015)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.013 -0.008 0.001 0.019 0.012 0.811
(Jan. 1960 - Apr. 2025) (0.014) (0.010) (0.014) (0.010)
Pre-1994 0.064** -0.312 0.046 0.055* 0.096 0.467
(Jan. 1960 - Dec. 1993) (0.030) (0.296) (0.031) (0.070)
Post-1994 -0.003 -0.022 0.016 -0.000 -0.004 0.280
(Jan. 1994 - Apr. 2025) (0.008) (0.015) (0.007) (0.014)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table C.7: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in inflation prediction with heteroskedasticity-robust standard errors
(a) TCODE Transformed Data
Sample Period Dlasso XDlasso VR Test
Full Sample 0.018*** -0.024 0.000
(Jan. 1960 - Apr. 2025) (0.006) (0.074)
Pre-Volcker 0.074*** 0.013 0.125
(Jan. 1960 - Jul. 1979) (0.018) (0.211)
Volcker-Greenspan -0.020 0.161 0.025
(Aug. 1979 - Jan. 2006) (0.020) (0.137)
Bernanke/Yellen/Powell -0.002 -0.054 0.002
(Feb. 2006 - Apr. 2025) (0.010) (0.098)
(b) Untransformed Data: Excluding I(2) Variables Based on TCODE
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.077 0.068 0.080 -0.050** 0.033 0.065
(Jan. 1960 - Apr. 2025) (0.068) (0.200) (0.022) (0.031)
Pre-Volcker -0.129 -0.014 0.522 0.007 0.113 0.548
(Jan. 1960 - Jul. 1979) (0.102) (0.274) (0.051) (0.089)
Volcker-Greenspan 0.094 -0.272 0.741 -0.092* -0.259 0.339
(Aug. 1979 - Jan. 2006) (0.185) (0.279) (0.054) (0.224)
Bernanke-Yellen-Powell 0.550 0.585 0.283 0.001 0.040 0.230
(Feb. 2006 - Apr. 2025) (0.522) (0.796) (0.076) (0.080)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table C.9: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in stock return prediction: Alternative set of I(2) variables and transformation with heteroskedasticity-robust standard errors
(a) Untransformed Data: Excluding I(2) Variables Based on smeekes2020unit
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.020 -0.008 0.000 0.025 0.012 0.831
(Jan. 1960 - Apr. 2025) (0.018) (0.010) (0.017) (0.010)
Pre-1994 0.035 -0.208 0.049 0.044 0.168 0.494
(Jan. 1960 - Dec. 1993) (0.051) (0.229) (0.042) (0.143)
Post-1994 -0.003 -0.022 0.011 -0.000 -0.009 0.163
(Jan. 1994 - Apr. 2025) (0.008) (0.015) (0.007) (0.015)
(b) Log Transformation: Excluding I(2) Variables Based on TCODE
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.011 -0.005 0.003 0.021 0.012 0.576
(Jan. 1960 - Apr. 2025) (0.014) (0.010) (0.014) (0.010)
Pre-1994 0.039 -0.360 0.033 0.033 0.087 0.489
(Jan. 1960 - Dec. 1993) (0.043) (0.294) (0.041) (0.206)
Post-1994 0.041* 0.020 0.010 0.045** 0.033 0.258
(Jan. 1994 - Apr. 2025) (0.021) (0.028) (0.020) (0.028)
(c) Log Transformation: Excluding I(2) Variables Based on smeekes2020unit
Without $\text{Return}_{t-1}$ Include $\text{Return}_{t-1}$
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.016 -0.009 0.001 0.022* 0.013 0.559
(Jan. 1960 - Apr. 2025) (0.012) (0.013) (0.012) (0.010)
Pre-1994 0.106* -0.394 0.032 0.108** 0.101 0.476
(Jan. 1960 - Dec. 1993) (0.056) (0.329) (0.055) (0.275)
Post-1994 0.018 -0.001 0.010 0.021 0.016 0.216
(Jan. 1994 - Apr. 2025) (0.016) (0.035) (0.016) (0.025)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.
Table C.11: Test $\mathbb{H}_{0}:\theta_{1}^{*}=0$ in inflation prediction: Alternative set of I(2) variables and transformation with heteroskedasticity-robust standard errors
(a) Untransformed Data: Excluding I(2) Variables Based on smeekes2020unit
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.172** 0.118 0.104 0.069* 0.116*** 0.015
(Jan. 1960 - Apr. 2025) (0.072) (0.206) (0.036) (0.042)
Pre-Volcker -0.169 -0.022 0.516 0.018 0.106 0.555
(Jan. 1960 - Jul. 1979) (0.134) (0.281) (0.054) (0.097)
Volcker-Greenspan -0.041 -0.200 0.669 -0.009 -0.189 0.641
(Aug. 1979 - Jan. 2006) (0.188) (0.265) (0.053) (0.120)
Bernanke-Yellen-Powell 0.550 0.294 0.220 0.081 0.162* 0.094
(Feb. 2006 - Apr. 2025) (0.395) (0.394) (0.057) (0.083)
(b) Log Transformation: Excluding I(2) Variables Based on TCODE
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample 0.014 -0.103 0.118 -0.063*** 0.044 0.005
(Jan. 1960 - Apr. 2025) (0.056) (0.102) (0.021) (0.057)
Pre-Volcker -0.155 -0.056 0.516 -0.029 0.131 0.827
(Jan. 1960 - Jul. 1979) (0.153) (0.256) (0.048) (0.103)
Volcker-Greenspan 0.064 -0.068 0.448 -0.082 -0.057 0.427
(Aug. 1979 - Jan. 2006) (0.131) (0.197) (0.065) (0.306)
Bernanke-Yellen-Powell 0.121 1.231* 0.398 -0.040 0.142 0.119
(Feb. 2006 - Apr. 2025) (0.183) (0.729) (0.084) (0.100)
(c) Log Transformation: Excluding I(2) Variables Based on smeekes2020unit
Include Labor Variables Exclude Labor Variables
Sample Period Dlasso XDlasso VR Test Dlasso XDlasso VR Test
Full Sample -0.016 -0.101 0.110 -0.030* 0.016 0.020
(Jan. 1960 - Apr. 2025) (0.057) (0.102) (0.016) (0.023)
Pre-Volcker -0.137 -0.068 0.430 -0.017 0.052 0.556
(Jan. 1960 - Jul. 1979) (0.154) (0.246) (0.054) (0.097)
Volcker-Greenspan -0.054 -0.136 0.871 -0.014 -0.172 0.702
(Aug. 1979 - Jan. 2006) (0.131) (0.314) (0.060) (0.118)
Bernanke-Yellen-Powell 0.018 0.555 0.541 0.087* 0.098* 0.180
(Feb. 2006 - Apr. 2025) (0.134) (0.438) (0.049) (0.055)
Notes: We report estimates and the standard error (in parentheses below the estimates) across methods and setups. The symbols *, **, and *** indicate significance levels at 10%, 5%, and 1%, respectively. “VR Test” represents the $p$-value of the variance ratio test (kim2009automatic) on the LASSO residual. The tuning parameter for LASSO estimation is selected through 10-fold block cross-validation. In XDlasso, instruments are generated based on (2.9) and (3.5) with $C_{\zeta}=5$ and $\tau=0.5$.